### v1

# Computing Wasserstein Barycenters through Gradient Flows

Abstract
Wasserstein barycenters provide a powerful tool for aggregating probability measures, while leveraging the geometry of their ambient space. Existing discrete methods suffer from poor scalability, as they require access to the complete set of samples from input measures. We address this issue by recasting the original barycenter problem as a gradient flow in the Wasserstein space. Our approach offers two advantages. First, we achieve scalability by sampling mini-batches from the input measures. Second, we incorporate functionals over probability measures, which regularize the barycenter problem through internal, potential, and interaction energies. We present two algorithms for empirical and Gaussian mixture measures, providing convergence guarantees under the Polyak-Łojasiewicz inequality. Experimental validation on toy datasets and domain adaptation benchmarks show that our methods outperform previous discrete and neural net-based methods for computing Wasserstein barycenters.
1 Introduction
(a) Static barycenter.
(b) Barycenter flow.
Figure 1: In (a), we show the usual static notion of the Wasserstein barycenter (in red), which minimizes the sum of distances to the input measures (in blue). In (b), we show our notion of barycenter as a gradient flow, flowing an initial measure $P_{0}$ (purple) to the barycenter $P^{\star}$ (yellow) of the input measures.
Wasserstein barycenters [1] constitute a building block in the analysis of probability measures under Optimal Transport (OT) [2]. This mathematical object extends the notion of the barycenter of points to Wasserstein spaces, thus defining a notion of the average of probability measures. Since it relies on the Wasserstein metric, these barycenters capture the geometry of the underlying space over which the probability measures are defined [3]. As such, they contributed to several areas of machine learning, including model averaging [4], ensembling [5], data augmentation [6], distillation [7, 8], dictionary learning [9], domain adaptation [10, 11, 12], and Bayesian learning [13].
Existing methods for computing Wasserstein barycenters can be divided into three categories. The first category relies on discretization of the input measures and the barycentric measure through empirical measures. These algorithms may be subdivided into fixed-support [14, Algorithm 1][15, 16], and free-support [14, Algorithm 2][11].
The second category relies on parametric models for the input measures, such as Gaussian, or Gaussian mixture models. In the Gaussian case, these barycenters are known as Bures-Wasserstein barycenters [17, 18]. For Gaussian mixtures, [19] shows connections with multi-marginal OT [20] between the components of the input mixtures, while [12] shows that, for components with diagonal covariances, there is an iterative algorithm similar to [14, Algorithm 2].
The third category of methods for computing Wasserstein barycenters relies on neural nets and comes from the effort to scale OT to measures in high-dimensional spaces with a large number of samples. In this sense, [21] proposes parameterizing the barycentric measure through a specific kind of neural net architecture, known as Input Convex Neural Networks (ICNNs) [22]. Furthermore, [23] proposes a bi-level adversarial learning approach that can be used for general ground costs, which is made more robust by semi-unbalanced OT [24]. More recently, [25] proposes computing barycenters with normalizing flows.
Current Wasserstein barycenter algorithms mostly focus on scalability with respect to the number of samples $n$, and dimensions $d$. In this sense, neural networks have a clear advantage over discrete methods, since they can operate at the level of mini-batches. There is another level of scalability, oftentimes overlooked, which is with respect the number of input measures $K$. So far, only [25] has considered experiments in this context. In this regard, methods such as [21, 26] are difficult to scale, as they require $\mathcal{O}(K)$ neural networks to compute the Wasserstein barycenter.
In this paper we propose a new perspective on discrete Wasserstein barycenters, through the lens of Wasserstein Gradient Flows (WGFs) [27, 28], which is an appealing framework for optimization in the space of probability measures. This new understanding of Wasserstein barycenters has several advantages. First, we can scale the discrete barycenter problem, by drawing mini-batches from the input measures. Second, we can scale with respect to the number of input measures, as the representation of the barycentric measure is independent of $K$. Third, based on the established literature on gradient flows, we can define regularized Wasserstein barycenters with suitable functionals on the space of probability measures, which were not possible with previous methods. Lastly, we can modify the underlying metric to integrate label information between samples, leading to barycenters that respect class structure and provide more accurate and semantically meaningful estimates. We show an overview of our approach in Figure 1. Our contributions are:
  • We introduce a unified framework for computing Wasserstein barycenters based on gradient flows, which provides a principled and flexible approach to optimizing in the space of probability measures.
  • We generalize algorithms from the state-of-the-art [14, 11, 12] under a common framework.
  • We offer theoretical guarantees for the convergence of our algorithms, which also apply to previous works.
  • We offer a new theoretical result for the mixture-Wasserstein distance between labeled Gaussian mixtures [19, 12].
  • We conduct extensive experiments benchmarking existing methods, including neural network-based Wasserstein barycenter algorithms [21, 23, 24, 25], on both toy datasets and domain adaptation  [^1] In contrast with discrete methods, most experiments in neural network solvers involve generative modeling. We thus establish a bridge between these methods and domain adaptation benchmarks, and show that our approach establishes a new state-of-the-art.
Paper organization. Section 2 presents the notation and preliminaries of our paper. Section 3 covers our proposed algorithms, including the theoretical results. Section 4 describes our experiments. Finally, section 5 concludes this paper.
2 Background
Throughout this paper, $(\Omega,d)$ is a metric space. We denote by $P$ and $Q$, probability measures on $\mathcal{P}(\Omega)$, which is the space of measures over $\Omega$. For $p\in[1,+\infty)$ we further denote by $\mathcal{P}_{p,ac}(\Omega)$, the sub-set of absolutely continuous measures w.r.t. the Lebesgue measure such that $\int_{\Omega}d(z,z_{0})^{p}dP(z)<+\infty$. In our case, we have access to these measures through their samples, denoted $z^{(P)}\sim P$. The set $\Delta_{K}$ denotes the $K$-simplex, i.e., $y\in\Delta_{n}=\{y\in\mathbb{R}^{n}:\sum_{i=1}^{n}y_{i}=1\text{ and }y_{i}\geq 0\text{ }\forall i\}$. We denote functionals over $\mathcal{P}_{2,\text{ac}}$ with blackboard bold letters (e.g., $\mathbb{F}$), and their (Wasserstein) gradients by $\nabla\mkern-10.0mu\nabla$. We offer a short introduction in Section 3 in the Appendix.
2.1 Empirical Optimal Transport
In this section, given $n$ i.i.d. samples from $P$, we can approximate a probability $P$ empirically through,
$\hat{P}(z)=\dfrac{1}{n}\sum_{i=1}^{n}\delta(z-z_{i}^{(P)}),$ (1)
which gives a finite parametrization for $P$. Henceforth we use a hat to indicate empirical measures (e.g., $\hat{P}$).
OT [2] is a field of mathematics concerned with the transportation of mass at least effort. We refer readers to [29] for a computational treatment of the subject, and [30] for applications in machine learning. Let $P,Q\in\mathcal{P}(\Omega)$, and $c:\Omega^{2}\rightarrow\mathbb{R}$ be a ground-cost. OT was originally founded by Monge [31]. In this formulation, we seek for a mapping $T:\Omega\rightarrow\Omega$ such that,
$T_{P\rightarrow Q}^{\star}=\underset{T_{\sharp}P=Q}{\text{arg inf}}\,\int_{\Omega}c(z,T(z))dP(z).$ (2)
Alternatively, Kantorovich [32] proposed a formulation in terms of a transport plan $\gamma\in\Gamma(P,Q)$, where $\Gamma(P,Q)$ is the set of all joint measures with marginals $P$ and $Q$. In this case,
$\gamma^{\star}=\underset{\gamma\in\Gamma(P,Q)}{\text{arg inf}}\,\int_{\Omega}\int_{\Omega}c(z,z^{\prime})\gamma(z,z)$ (3)
Given $p\in[1,+\infty)$, when $c(z,z^{\prime})=d(z,z^{\prime})^{p}$, the previous problems define the $p$-Wasserstein distance,
$\mathbb{W}_{p}(P,Q)^{p}=\underset{\gamma\in\Gamma(P,Q)}{\text{inf}}\,\int_{\Omega}\int_{\Omega}d(z,z^{\prime})^{p}d\gamma(z,z^{\prime}).$ (4)
Equation 4 defines a metric in $\mathcal{P}_{p,ac}(\Omega)$. Conceptually, the Wasserstein distance lifts the metric $d$ on $\Omega$ to $\mathbb{W}_{p}$ on $\mathcal{P}_{p,ac}(\Omega)$. Based on this idea, we can define barycenters and gradient flows in the space of probability measures.
2.2 Gaussian Mixture Optimal Transport
In comparison with equation 1, we can approximate $P$ through a Gaussian Mixture Model (GMM),
$P(z)=\sum_{i=1}^{n}\pi_{i}^{(P)}\mathcal{N}(z|\mu_{i}^{(P)},\Sigma_{i}^{(P)}),$ (5)
where $\theta=\{(\pi_{i}^{(P)},\mu_{i}^{(P)},\Sigma_{i}^{(P)})\}$ are the parameters of the GMM. The general theory of OT between GMMs was first presented in [19]. The main advantage is that, when the OT plan is restricted to the set of GMMs, i.e, $\gamma\in\Gamma(P,Q)\cap\text{GMM}_{\infty}(d)$, there is a discrete equivalent formulation in terms of the GMM components,
$\omega^{\star}=\underset{\omega\in\Gamma(\pi^{(P)},\pi^{(Q)})}{\text{arg min}}\,\sum_{i=1}^{n}\sum_{j=1}^{m}\omega_{ij}\mathbb{W}_{2}(P_{i},Q_{j})^{2}.$
$\omega^{\star}\in\mathbb{R}^{n\times m}$ is a component-to-component OT plan. The term $\mathbb{W}_{2}(P_{i},Q_{j})$ is the Wasserstein distance between Gaussian components, which has closed-form [33] in terms of their parameters,
$\mathbb{W}_{2}(P,P^{\prime})^{2}=\lVert\mu-\mu^{\prime}\rVert_{2}^{2}+\text{Tr}(\Sigma+\Sigma^{\prime}-2(\Sigma^{\frac{1}{2}}\Sigma^{\prime}\Sigma^{\frac{1}{2}})^{\frac{1}{2}}).$
Based on these ideas, one may define a Wasserstein-type distance between GMMs,
$\mathbb{MW}_{2}(P,Q)^{2}=\underset{\omega\in\Gamma(\pi^{(P)},\pi^{(Q)})}{\text{inf}}\,\sum_{i=1}^{n}\sum_{j=1}^{m}\omega_{ij}\mathbb{W}_{2}(P_{i},Q_{j})^{2}.$ (6)
2.3 Wasserstein Barycenter and its Variants
In the metric setting, the barycenter problem is known as Fréchet [34] or Karcher [35] means. In our case, given a finite family of measures $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}$ and a set of barycentric coordinates $\lambda\in\Delta_{K}$, we define the Wasserstein barycenters over $(\mathcal{P}_{p,ac}(\Omega),\mathbb{W}_{p})$ through the following optimization problem,
$P^{\star}=\underset{P\in\mathcal{P}(\Omega)}{\text{arg min}}\,\biggr\{\mathbb{B}_{Q}(P)=\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{p}(P,Q_{k})^{p}\biggr\}.$ (7)
In general, equation 7 does not have a closed-form solution. However, there are algorithms for computing barycenters for $p=2$, when the measures in $\mathcal{P}$ are either empirical measures [14], Gaussian measures [1, 36] or Gaussian mixtures [19, 12], and a further link can be made to multi-marginal OT [20].
The work of [14] is of particular interest to us, since theirs was the first algorithm, based on gradient descent, to optimize equation 7 on empirical measures. The idea is to initialize $z_{0,1}^{(P)},\cdots,z_{0,n}^{(P)}$ randomly (e.g., from a Gaussian measure), and iterate
$z_{\tau+1,i}^{(P)}=(1-\alpha)z_{\tau,i}^{(P)}+\alpha\sum_{k=1}^{K}\lambda_{k}T_{\gamma_{k}}(z_{\tau,i}^{(P)}),$ (8)
where $T_{\gamma_{k}}(z_{i}^{(P)})=n\sum_{j=1}^{n}\gamma_{k,i,j}z_{j}^{(Q_{k})}$ is the barycentric map between $Q_{k}$ and $P$. This strategy is reminiscent of the fixed-point approach of [36]. Furthermore, [11] shows that, for feature-label joint measures, the labels can be propagated through [37],
$y_{\tau+1,i}^{(P)}=(1-\alpha)y_{\tau,i}^{(P)}+\alpha\sum_{k=1}^{K}\lambda_{k}T_{\gamma_{k}}(y_{\tau,i}^{(P)}),$ (9)
where $y_{\tau,i}^{(P)}$ are one-hot encoded vectors.
Previous work has extended equation 8 to the GMM setting, when the components are axis-aligned Gaussians, i.e., $\Sigma_{i}^{(P)}=\text{diag}(\sigma_{i}^{(P)})$. In this case,
$\mu_{i,\tau+1}^{(P)}=(1-\alpha)\mu_{i,\tau}^{(P)}+\alpha\sum_{k=1}^{K}\lambda_{k}T_{\omega_{k}}(\mu_{i,\tau}^{(P)}),$ (10)
$\sigma_{i,\tau+1}^{(P)}=(1-\alpha)\sigma_{i,\tau}^{(P)}+\alpha\sum_{k=1}^{K}\lambda_{k}T_{\omega_{k}}(\sigma_{i,\tau}^{(P)}),$
There are a few limitations with equations 8 and 10. First, the empirical iterations do not scale well to large datasets. Indeed, they assume the availability of all samples of $Q_{k}$ per iteration. Second, the GMM iterations in equation 10 are restricted to axis-aligned GMM s. This paper addresses these gaps using gradient flows.
2.4 Gradient Flows
In the Euclidean setting, let $F:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a functional. A gradient flow is the solution to the differential equation,
$\begin{cases}\dot{x}(t)=-\nabla F(x(t))&\text{for }t>0,\\ x(0)=x_{0},\end{cases}$ (11)
where $x_{0}$ is the initial condition. In $\mathcal{P}_{p,\text{ac}}(\Omega)$, gradient flows correspond to the continuity equation,
$\partial_{t}P_{t}=-\text{div}(P_{t}\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t})),$ (12)
where $P_{t}$ is a curve in $\mathcal{P}_{p,\text{ac}}(\Omega)$, and $\mathbb{F}:\mathcal{P}_{p,\text{ac}}(\Omega)\rightarrow\mathbb{R}$ is a functional over probability measures. Usually, these functionals take the following form [38],
$\mathbb{F}(P)=$ $\underbrace{\int G(P(z))dz}_{\mathbb{G}(P)}+\underbrace{\int V(z)dP(z)}_{\mathbb{V}(P)}+$ (13)
$\underbrace{\int\int U(z,z^{\prime})dP(z)dP(z^{\prime})}_{\mathbb{U}(P)},$
where $\mathbb{G}$, $\mathbb{V}$, and $\mathbb{U}$ are the internal, potential, and interaction energies, respectively. Here, $G:\mathbb{R}\rightarrow\mathbb{R}$ is convex and superlinear, and $V:\Omega\rightarrow\mathbb{R}$ and $U:\Omega^{2}\rightarrow\mathbb{R}$ are convex and sufficiently smooth. Henceforth, we denote $\mathbb{F}^{\star}=\text{inf}_{P\in\mathcal{P}_{2,\text{ac}}(\Omega)}\mathbb{F}(P)$.
3 Wasserstein Barycenters as Gradient Flows
In this section, we describe a new method for computing empirical and Gaussian mixture Wasserstein barycenters. Our main idea is using the functional:
$\mathbb{F}(P)=\mathbb{B}_{\mathcal{Q}}(P)+\mathbb{G}(P)+\mathbb{V}(P)+\mathbb{U}(P),$ (14)
where $\mathbb{B}_{\mathcal{Q}}(P)$ is the barycenter objective defined in equation 7, and $\mathbb{G},\mathbb{V}$ and $\mathbb{U}$ are the energies defined in equation 13.
3.1 Empirical Flow
In the empirical case, the barycenter is approximated through a finite set of particles (see equation 1) $\{z_{i}^{(P)}\}_{i=1}^{n}$. Hence, we may rewrite equation 14 as,
$\underset{\begin{subarray}{c}z_{1}^{(P)},\cdots,z_{n}^{(P)}\in\Omega,\\ \gamma_{k}\in\Gamma(\hat{P},\hat{Q}_{k})\end{subarray}}{\text{min}}\,\underbrace{\sum_{k=1}^{K}\lambda_{k}\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}d(z_{i}^{(P)},z_{j}^{(Q_{k})})^{p}}_{\mathbb{B}_{\mathcal{Q}}(\hat{P})}$+
$\underbrace{\dfrac{1}{n}\sum_{i=1}^{n}V(z_{i}^{(P)})}_{\mathbb{V}(\hat{P})}+\underbrace{\dfrac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}U(z_{i}^{(P)},z_{j}^{(P)})}_{\mathbb{U}(\hat{P})}$ (15)
Here, we omit the internal energy term, $\mathbb{G}(P)$, as it is not defined for empirical measures. We solve the previous optimization problem through a block-coordinate descent strategy, that is,
$\gamma_{k}$ $=\underset{\gamma\in\Gamma(\hat{P}_{\tau},\hat{Q}_{k})}{\text{arg min}}\,\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{i,j}d(z_{\tau,i}^{(P)},z_{j}^{(Q_{k})})^{p},$
$z_{\tau+1,i}^{(P)}$ $=z_{\tau,i}^{(P)}-\alpha\nabla\mkern-10.0mu\nabla\mathbb{F}(\hat{P}_{\tau}).$ (16)
In equation 16, $\gamma_{k}$ is used to compute $\mathbb{B}_{\mathcal{Q}}$ and is updated for each $\tau$. We provide in Algorithm 1 the pseudo-code for this strategy. Our algorithm assumes access to $Q_{k}$ through sampling, that is, we do not necessarily have access to all samples of $Q_{k}$ at once.
Algorithm 1 Empirical barycenter Wasserstein flow using Gradient Descent.
[…]
2:Barycenter support $\{z_{i}^{(P)}\}_{i=1}^{n}$
3:for $\tau\leftarrow 1$ to $n_{\text{iter}}$ do
4:  $\mathbb{F}(\hat{P}_{\tau})\leftarrow\mathbb{V}(\hat{P}_{\tau})+\mathbb{U}(\hat{P}_{\tau})$
5:  for $k\leftarrow 1$ to $K$ do
6:   Sample $\{z_{i}^{(Q_{k})}\}_{i=1}^{m}$ i.i.d. from $Q_{k}$.
7:   $\mathbb{F}(\hat{P}_{\tau})\leftarrow\mathbb{F}(\hat{P}_{\tau})+\lambda_{k}\mathbb{W}_{p}(\hat{P}_{\tau},\hat{Q}_{k})^{p}$
8:  end for
9:  $z_{\tau+1,i}^{(P)}\leftarrow z_{\tau,i}^{(P)}-\alpha\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{\tau})$
10:end for
Remark 3.1.
In line 5 of Algorithm 1, $\hat{Q}_{k}$ denotes the empirical measure (c.f., equation 1) supported on the mini-batch $\{z_{i}^{(Q_{k})}\}_{i=1}^{m}$ sampled from $Q_{k}$.
3.2 Gaussian Mixture Flow
In the case of GMMs, we flow the parameters $\theta=\{(\pi_{k},\mu_{k},\Sigma_{k})\}_{k=1}^{K}$ of $P_{\theta}=\sum\pi_{i}\mathcal{N}(\cdot|\mu_{i},\Sigma_{i})$. Similarly to equation 16, we have:
$\omega_{k}$ $=\underset{\omega\in\Gamma(\pi,\pi_{k})}{\text{arg min}}\,\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\omega_{i,j}\mathbb{W}_{2}(P_{\theta_{\tau},i},Q_{k,j})^{2},$
$\theta_{\tau+1,i}$ $=\theta_{\tau,i}-\alpha\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{\theta_{\tau}}).$ (17)
Here, $\omega_{k}$ is a transport plan between GMM components. In general, the energy functionals do not have closed form with respect to the GMM parameters. However, we can estimate them using Monte-Carlo and the reparametrization trick, for instance,
$z_{i}$ $=L_{i}\varepsilon+\mu_{i},\quad i\sim\pi,\varepsilon\sim\mathcal{N}(\cdot|0,\text{Id}),$ (18)
This strategy allows us to propagate the gradients to the parameters of $P_{\tau}$. We show the overall approach in Algorithm 2.
Algorithm 2 GMM barycenter flow using Gradient Descent.
[…]
2:Barycenter support $\{z_{i}^{(P)}\}_{i=1}^{n}$
3:for $\tau\leftarrow 1$ to $n_{\text{iter}}$ do
4:  Sample $\{z_{i}=L_{i}\varepsilon+\mu_{i}:i\sim\pi,\varepsilon\sim\mathcal{N}(\cdot|0,\text{Id})\}$,
5:  $\mathbb{F}(\hat{P}_{\tau})\leftarrow\mathbb{G}(\hat{P}_{\tau})+\mathbb{V}(\hat{P}_{\tau})+\mathbb{U}(\hat{P}_{\tau})$
6:  for $k\leftarrow 1$ to $K$ do
7:   $\mathbb{F}(P_{\theta_{\tau}})\leftarrow\mathbb{F}(P_{\theta_{\tau}})+\lambda_{k}\mathbb{MW}_{2}(P_{\theta_{\tau}},Q_{k})^{2}$
8:  end for
9:  $z_{\tau+1,i}^{(P)}\leftarrow z_{\tau,i}^{(P)}-\alpha\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{\theta_{\tau}})$
10:end for
Covariance matrix parametrization. Optimizing $\mathbb{F}$ with respect to $\Sigma_{k}$ is challenging, due to the symmetric positive definite constraint. We enforce this constraint by writing $\Sigma_{k}=L_{k}L_{k}^{T}$, where $L_{k}$ is the lower-triangular Cholesky factor. There are two advantages to this strategy. First, optimizing with respect to $L_{k}$ is empirically more stable than $\Sigma_{k}$. Second, we can use $L_{k}$ directly in the reparametrization trick in equation 18.
3.3 Flows over Joint Measures
In [14], the authors considered barycenters for $p=2$, $\Omega=\mathbb{R}^{d}$, and $d(z,z^{\prime})=\lVert z-z^{\prime}\rVert_{2}$. Therefore, we generalize this setting, since we only need the differentiability of $d$ for our algorithm to work. As we show in our experiments, in machine learning applications one encounters $\Omega=\mathcal{X}\times\mathcal{Y}$, i.e., measures over the joint space of features and labels. Whenever we deal with labeled measures (i.e., $z=(x,y)$), we adopt,
$d(z,z^{\prime})=\sqrt{\lVert x-x^{\prime}\rVert_{2}^{2}+\beta\lVert y-y^{\prime}\rVert_{2}^{2}},$ (19)
where $\beta\geq 0$ is a parameter that balances the feature distance terms and the label distance terms.
For regression applications (i.e., $\mathcal{Y}=\mathbb{R}$), this distance is quite natural. However, for classification, $\mathcal{Y}$ is categorical (e.g., $\mathcal{Y}=\{1,\cdots,n_{\text{classes}}\}$). One possible strategy, used in [10] and [39], is fixing the labels and flowing only the features. In contrast to these works, we embed $\mathcal{Y}$ into the compact continuous space $\Delta_{K}$, through a one-hot encoding operation. For our flow, we parametrize labels through a change of variables,
$y_{i,c}^{(P)}=\text{softmax}(\ell_{i,1}^{(P)},\cdots,\ell_{i,n_{c}}^{(P)})=\dfrac{\exp(\ell_{i,c}^{(P)})}{\sum_{c=1}^{n_{c}}\exp(\ell_{i,c}^{(P)})},$
thus, instead of optimizing over $z=(x,y)$, we optimize over $z=(x,\ell)$. From the soft probabilities, we can retrieve the actual discrete labels with an argmax, $y_{i}^{(P)}=\text{argmax}_{c=1,\cdots,n_{c}}y_{i,c}^{(P)}$.
For GMMs, we equip their components with labels $\nu_{k}\in\Delta_{n_{\text{classes}}}$. Using the ground-cost in equation 19, we are able to write the $\mathbb{MW}_{2}(P,Q)^{2}$ as,
$\sum_{i=1}^{n}\sum_{j=1}^{m}\omega_{ij}^{\star}(\mathbb{W}_{2}(P_{i,x},Q_{j,x})^{2}+\beta\lVert\nu_{i}^{(P)}-\nu_{j}^{(Q)}\rVert_{2}^{2}),$
where $P_{i,x}=\mathcal{N}(\cdot|\mu_{i}^{(P)},\Sigma_{i}^{(P)})$ is the Gaussian feature marginal. This approach was first proposed in [12], but the authors relied on heuristic arguments to justify this modeling choice. We present the following proposition justifying the label term,
Proposition 3.1.
Let $P=\sum\pi_{i}^{(P)}(\mathcal{N}(\mu_{i}^{(P)},\Sigma_{i}^{(P)})\otimes\delta_{\nu_{i}^{(P)}})$ and $Q=\sum\pi_{j}^{(Q)}(\mathcal{N}(\mu_{j}^{(Q)},\Sigma_{j}^{(Q)}\otimes\delta_{\nu_{j}^{(Q)}})$ be two GMMs over $\Omega=\mathcal{X}\times\mathcal{Y}$. Let the ground cost $c$ be,
$c(z,z^{\prime})=\lVert x-x^{\prime}\rVert_{2}^{2}+\rho(y,y^{\prime})^{2},$
where $\rho(y,y^{\prime})$ is a metric over $\mathcal{Y}$. Then,
$\mathbb{MW}_{2}(P,Q)^{2}=\underset{\omega\in\Gamma(\pi^{(P)},\pi^{(Q)})}{\text{min}}\,\sum_{i=1}^{n}\sum_{j=1}^{m}\omega_{ij}C_{ij},$
where $C_{ij}=\mathbb{W}_{2}(P_{i,x},Q_{j,x})^{2}+\rho(\nu_{i}^{(P)},\nu_{j}^{(P)})^{2}$.
Functionals for joint measures. One of the advantages of our proposed method is regularizing the barycenter calculation with internal, interaction, and potential energy functionals. This idea was already used in practice by [39] for transfer learning problems. Here we propose the following functionals,
$V_{E}(z)$ $=-\sum_{c=1}^{n_{\text{classes}}}y_{c}\log y_{c},$ (20)
$U_{R}(z,z^{\prime})$ $=\begin{cases}h(d(x,x^{\prime}))&\text{if }y\neq y^{\prime},\\ 0&\text{otherwise}\end{cases},$ (21)
where $h:\mathbb{R}\rightarrow\mathbb{R}$ is lower semi-continuous and bounded from below. For instance, in our experiments we use the hinge loss, $h(u)=\text{max}(0,\text{margin}-d)$, where $\text{margin}\geq 0$ is a margin parameter. Equations 20 and 21 correspond to entropy and repulsion terms. The first functional penalizes barycenters that have fuzzy labels. The second functional encourages classes to be well separated.
3.4 Convergence
One of the difficulties in analyzing the gradient flow of the functional in equation 14 comes from the fact that $P\mapsto\mathbb{W}_{p}(P,Q)^{p}$ is not geodesically convex in $\mathcal{P}_{p,ac}(\Omega)$ [28, Section 4.4]. As a result, we are minimizing a non-convex functional. Henceforth, we focus on $p=2$, as this is the setting where most results are available. Explicit constants and proofs are available in Section 4 in the Appendix.
Our analysis relies on a measure-theoretic version of the Polyak Łojasiewicz (PL) inequality [40, 41],
$\lVert\nabla\mkern-10.0mu\nabla\mathbb{B}(P)\rVert_{L_{2}(P)}\geq C_{\text{PL}}(\mathbb{B}(P)-\mathbb{B}^{\star}),$ (22)
which is covered in [42]. Under this assumption, we have the following convergence result,
Theorem 3.1.
Let $\Omega=\mathcal{B}(0,R)$ be the closed ball on $\mathbb{R}^{d}$ with radius $R>0$. Let $P^{\star}\in\mathcal{P}_{2,\text{ac}}(\mathcal{B}(0,R))$ be the barycenter of $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K},Q_{k}\in\mathcal{P}_{2,\text{ac}}(\mathcal{B}(0,R))$ $k=1,\cdots,K$, with barycentric coordinates $\lambda=(\lambda_{1},\cdots,\lambda_{K})\in\Delta_{K}$. Assume the inequality 22. Then, the following holds,
$\begin{aligned} \mathbb{E}[\hat{\mathbb{B}}(P_{\tau})-\hat{\mathbb{B}}^{\star}]\leq e^{-C_{\text{PL}}\tau}(\mathbb{B}(P_{0})-\mathbb{B}^{\star})+C_{R}\sqrt{\dfrac{C_{d,m}}{n}}\end{aligned},$ (23)
where the expectation on the l.h.s. is taken with respect to samples from $Q_{k}$. The constants $C_{R}$ and $C_{d,m}$ depend on the radius $R$, number of barycenter samples $m$, and dimensions $d$.
The right hand side of inequality 23 is composed of two terms. The first term comes from the PL inequality, and covers the convergence towards the minimizer of $\mathbb{B}$. The second term covers the error of empirical approximations. In this sense, letting $\tau\rightarrow+\infty$ leads to an error governed by the empirical approximation.
Next, we derive a new result for the convergence of the GMM gradient flow (cf. Algorithm 2). Here, we assume $P$ and $Q_{k}\in\mathcal{Q}$ are GMMs. Furthermore, we denote by $\hat{Q}_{k}$ the GMM fitted on data from $Q_{k}$ with the expectation-maximization algorithm. Our result relies on [43, Theorem 3.1]. Therefore, we assume,
  1. 1.
    bounded GMM parameters, i.e., $\lVert\mu_{i}^{(Q_{k})}\lVert_{2}\leq R_{\mu}$, $\sqrt{\text{Tr}(\Sigma_{i}^{(Q_{k})})}\leq R_{\Sigma}$ $\forall i,\forall k$ (resp. $\hat{\mu}_{i}^{(Q_{k})}$ and $\hat{\Sigma}_{i}^{(Q_{k})}$, the estimated parameters); furthermore,
  2. 2.
    Empirical convergence rates, i.e., $\mathbb{E}[\lVert\pi^{(Q_{k})}-\hat{\pi}^{(Q_{k})}\rVert_{1}]\leq\rho_{\pi}$, $\mathbb{E}[\lVert\mu_{i}^{(Q_{k})}-\hat{\mu}_{i}^{(Q_{k})}\rVert_{2}]\leq\rho_{\mu}$, and $\mathbb{E}[d_{\text{Bures}}(\Sigma_{i}^{(Q_{k})},\hat{\Sigma}_{i}^{(Q_{k})})]\leq\rho_{\Sigma}$, $\forall i$ and $\forall k$.
Here the expectations are taken with respect to samples drawn from $Q_{k}$. Under these conditions we have,
Theorem 3.2.
Let $P$ and $Q$ be two labeled GMMs with bounded parameters and satisfying the empirical convergence rates. Then, the following holds,
$\mathbb{E}[\hat{\mathbb{B}}(P_{\tau})-\hat{\mathbb{B}}^{\star}]\leq e^{-C_{\text{PL}\tau}}(\mathbb{B}(P_{0})-\mathbb{B}^{\star})+C_{\lambda,\mathcal{Q}},$
where $C_{\lambda,\mathcal{Q}}$ is a constant that depends on the coordinates $\lambda$ and GMMs in $\mathcal{Q}$.
Here $C_{\lambda,\mathcal{Q}}$ depends on the $(\pi^{(Q_{k})},\mu^{(Q_{k})},\Sigma^{(Q_{k})})$ of each GMM, but not on $\nu^{(Q_{k})}$. Indeed, since we fit a GMM per class, the labels are estimated exactly (i.e., $P_{i,x}$ and $\hat{P}_{i,x}$ belong to the same class).
4 Experiments
We divide our experiments into a toy example (Section 4.1) to illustrate our methods, and multi-source domain adaptation experiments (Section 4.2). Due to length constraints, we include additional experiments in Sections 5 and 6 in the Appendix. In particular, in Table 1 in the Appendix, we show a running-time comparison demonstrating the scalability of our method.
4.1 Toy Example
In Figure 2, we show the Swiss roll measure $Q_{0}$, alongside four variations obtained via $T_{k,\sharp}Q_{0}$, $k=1,\cdots,4$. This example has been used in several state-of-the-art Wasserstein barycenter works [21, 25]. Originally there are labels associated with the samples of the Swiss-roll measure, corresponding to the position of the points in the underlying manifold. We show a summary of our results in Figure 3.
Figure 2: Location-scatter family generated by a Swiss-roll measure, $Q_{0}$. Each measure $Q_{k}=T_{k,\sharp}Q_{0}$, for $T_{k}(x)=A_{k}x+b_{k}$.
First, we compare unsupervised barycenter solvers in Figures 3 (a) through (h). Quantitatively, empirical methods, notably [14, Algorithm 2] and our WGF algorithm (cf. Algorithm 1) achieve the lowest Wasserstein distance to the ground truth in Figure 2. We note that, overall, while the neural network solvers are usually more scalable than discrete methods in terms of number of samples, their optimization problem is more complicated and very sensitive to hyper-parameters.
(a) CW2B [21].
(b) NOT [23].
(c) Discrete [14].
(d) WGF.
[…]
(g) U-NOT [24].
(h) NormFlow [25].
(i) Discrete [11].
(j) WGF.
[…]
Figure 3: Comparison between Wasserstein barycenter solvers. Colored scatter plots indicate labeled barycenters. For each solver, we compute the Wasserstein distance between its solution $\hat{P}$ and the ground-truth $P^{\star}$, shown in the title of each sub figure (best seen on screen). Overall, using label information leads to barycenters that better approximate the ground-truth barycenter.
Second, we experiment with integrating labels in the ground cost as described in Section 3.3. These methods are shown in Figures 3 (i) through (l). In all cases, integrating labels produces barycenters that are closer to the ground-truth. We conclude that using the labels gives a strong inductive bias in the barycenter computation, which helps explain the gain in performance of labeled barycenters in the next section.
4.2 Multi-Source Domain Adaptation
Problem Formulation. One of the main applications of Wasserstein barycenters is multi-source domain adaptation (MSDA) [44, 45, 46]. In this setting one needs to adapt multiple labeled source measures $Q_{1},\cdots,Q_{K}$ to a single unlabeled target measure $Q_{T}$. The goal is to learn, from samples $\{\{x_{i}^{(Q_{k})},y_{i}^{(Q_{k})}\}_{i=1}^{n_{k}}\}_{k=1}^{K}$ and $\{x_{i}^{(Q_{T})}\}_{i=1}^{n_{T}}$, a classifier $h$ that achieves low risk in the target domain measure,
$\mathcal{R}_{Q_{T}}(h)=\mathbb{E}_{(x,y)\sim Q_{T}}[\mathcal{L}(y,h(x))],$
for a loss function $\mathcal{L}$ (e.g., cross-entropy loss). We isolate the quality of barycenters by doing adaptation at the level of embeddings. This approach allows us to perform domain adaptation in a higher semantic space, where distributions are more meaningful and comparable across domains. Thus, we assume that a meaningful feature extractor $\phi$, called the backbone, has been previously learned. We obtain the feature extractor by fine-tuning a neural network on the labeled source domain data. In addition, similar to [10, 12], we align the barycenter with the target through OT [47]. More details are available in Section 6 in the Appendix.
Benchmark Backbone # Samples # Domains # Dim. # Classes
Office31 ResNet50 [48] 3287 3 2048 31
BCI-CIV-2a CBraMod [49] 5184 10 200 4
TEP CNN [50] 17289 6 128 29
Office Home ResNet101 [48] 15500 4 2048 65
ISRUC CBraMod [49] 89240 100 512 5
Table 1: Overview of benchmarks used for domain adaptation, sorted by number of samples.
Experimental Setting. We run our experiments on five benchmarks: Office 31 [51], Office Home [52], BCI-CIV-2a [53], ISRUC [54], and TEP [50]. The first two, second two, and last benchmarks correspond to computer vision, neuroscience, and chemical engineering benchmarks, respectively. We show in Table 1 an overview of our experimental setting.
Compared Methods. Overall, we compare seven Wasserstein barycenter strategies with ours. These methods are: discrete barycenters [14, Algorithm 2], normalizing flows [25], continuous 2-Wasserstein barycenters [21], neural OT barycenters [23], and unbalanced neural OT barycenters [24]. We also include labeled barycenter strategies [11, 55]. A brief review of these methods is available in Section 2 in the Appendix. For completeness, we include four other state-of-the-art methods in domain adaptation over embedding vectors. Those methods are: WJDOT [56], DaDiL-R and E [11], and GMM-DaDiL [12]. For each benchmark, we use one domain as the target domain (e.g., Amazon vs. $\{\text{dSLR, Webcam}\}$ in the Office 31 benchamrk), and we measure the classification accuracy, i.e., the percentage of correct predictions.
Benchmark $\mathcal{X}\times\mathcal{Y}$ Office31 OfficeHome BCI-CIV-2a ISRUC TEP
Backbone - ResNet-50 ResNet-101 CBraMod CBraMod CNN
Source-Only - 86.40 75.95 50.30 76.63 78.48
WJDOT [56] - 86.80 76.59 N/A 76.95 86.13
DaDiL-R [11] - 89.91 77.86 53.41 74.68 86.14
DaDiL-E [11] - 89.79 78.14 N/A 75.89 85.87
GMM-DaDiL [12] - 90.63 78.81 57.10 75.47 86.85
Discrete [14] 81.94 70.42 57.38 70.67 83.81
NormFlow [25] 85.91 76.73 55.45 78.04 82.89
CW2B [21] 86.37 76.44 57.25 75.84 85.83
NOT [23] 86.22 75.36 57.29 76.39 84.93
U-NOT [24] 86.97 76.85 57.17 77.06 85.43
Discrete [11] 87.93 77.09 57.43 78.20 86.09
GMM [12] 88.54 77.87 57.04 74.58 84.67
WGF (ours) 88.14 77.83 57.50 79.78 86.21
WGF-GMM (ours) 89.31 78.71 57.35 78.72 86.77
Table 2: Average classification accuracy on target domains. In total, we compare eight methods against our WGF strategy. For barycenter solvers, we indicate by ✓methods that compute barycenters on $\Omega=\mathcal{X}\times\mathcal{Y}$. Bold numbers represent the best barycenter method, and underlined numbers represent second best.
Main results. We present our main results in Table 2, which reports the average performance per domain on each benchmark. We provide fine-grained results in Section 6 in the Appendix. Overall, labeled barycenter methods (the four last rows in Table 2) have a clear advantage with respect to the unsupervised methods. We argue that label information is pivotal in domain adaptation success, which is consistent with previous research [47, 10, 11, 12]. Among the labeled barycenter methods, our gradient flow framework achieves either the best performance, or remains competitive, surpassing previous methods in MSDA, such as WJDOT and DaDiL in the ISRUC benchmark.
Figure 4: t-SNE [57] visualization of the barycenter of $\text{Art},\text{Product},\text{Real-World}]$ source domains in the Office home benchmark. Colors represent different classes from 11 to $65$. Overall, the classes tend to be more separated when using the repulsion functional $\mathbb{U}$.
Visualization. We now visualize the effect of adding the repulsion interaction energy functional on the Office-Home benchmark. Adding $\mathbb{U}$ leads to classes that are well separated. For computer vision experiments we use ResNets [48] as the backbone, which produce $2048$-dimensional embeddings. For that reason, we use the cosine distance $d(x,x^{\prime})=1-\text{cossim}(x,x^{\prime})$ in equation 21 and we fix $\text{margin}=1$. This choice encourages class clusters to live in orthogonal sub-spaces of the latent space. We show a comparison of the obtained barycenters in Figure 4. Overall, the classes tend to have less overlap when the interaction energy is added to the barycenter functional.
Ablation. We ablate the effect of $\mathbb{V}$ and $\mathbb{U}$ in the Office Home benchmark. For this benchmark, we use $\mathbb{V}(P)=\mathbb{W}_{2}(\hat{P},\hat{Q}_{T})^{2}$. The intent here is to isolate the contribution of each component to the state-of-the-art results in Table 2. Our results are summarized in Table 3, where we further compare our results with other related barycenter methods. In the empirical setting, the use of Algorithm 1 alone already presents a gain in performance compared to its discrete counterpart. Overall, the combination of $\mathbb{B}+\mathbb{V}+\mathbb{U}$ consistently gives the best performance, showing the importance of adding underlying structure to the Wasserstein barycenter in domain adaptation.
Method $\mathcal{X}\times\mathcal{Y}$ $\mathbb{B}$ $\mathbb{B}+\mathbb{V}$ $\mathbb{B}+\mathbb{U}$ $\mathbb{B}+\mathbb{V}+\mathbb{U}$
Discrete [14] 70.42 N/A N/A N/A
Discrete [11] 77.09 N/A N/A N/A
WGF (ours) 77.48 77.07 77.08 77.83
GMM [12] 77.87 N/A N/A N/A
WGF-GMM (ours) 77.28 78.28 78.12 78.71
Table 3: Ablation of average target domain classification accuracy on the Office Home benchmark.
5 Conclusion
In this paper we have introduced a new framework for computing Wasserstein barycenters, based on Wasserstein gradient flows [27, 28]. Our approach addresses scalability issues in discrete barycenter solvers [14], and is capable of regularizing the underlying barycenter through internal, potential, and interaction functionals. We presented gradient flow Algorithms 1 and 2, for empirical and Gaussian mixture measures, respectively. We further proved convergence guarantees based on the PL inequality.
We empirically tested our method against existing discrete [14, 11] and neural-network [21, 23, 24, 25]-based barycenter solvers on computer vision, neuroscience, and chemical engineering benchmarks, demonstrating that our proposed methods consistently achieve state-of-the-art performance. Our empirical findings (Table 2) further demonstrate that incorporating label information in the OT objective is key for domain adaptation. This finding highlights an interesting gap in the literature, where neural net-based solvers are not capable of exploiting this information, and thus have sub-optimal performance in domain adaptation. We leave an investigation of this issue for future work.

### v2

@@
 In (b), we show our notion of barycenter as a gradient flow, flowing an initial measure $P_{0}$ (purple) to the barycenter $P^{\star}$ (yellow) of the input measures.
-Wasserstein barycenters [1] constitute a building block in the analysis of probability measures under Optimal Transport (OT) [2].
+Wasserstein barycenters [agueh2011barycenters] constitute a building block in the analysis of probability measures under Optimal Transport (OT) [villani2008optimal].
 This mathematical object extends the notion of the barycenter of points to Wasserstein spaces, thus defining a notion of the average of probability measures.
-Since it relies on the Wasserstein metric, these barycenters capture the geometry of the underlying space over which the probability measures are defined [3].
-As such, they contributed to several areas of machine learning, including model averaging [4], ensembling [5], data augmentation [6], distillation [7, 8], dictionary learning [9], domain adaptation [10, 11, 12], and Bayesian learning [13].
+Since it relies on the Wasserstein metric, these barycenters capture the geometry of the underlying space over which the probability measures are defined [kloeckner2010geometric].
+As such, they contributed to several areas of machine learning, including model averaging [singh2020model], ensembling [coz2023barycenter], data augmentation [zhu2023interpolation], distillation [liu2023dataset, montesuma2024multisource], dictionary learning [schmitz2018wasserstein], domain adaptation [montesuma2021wasserstein, montesuma2023multi, montesuma2024lighter], and Bayesian learning [srivastava2018scalable].
 Existing methods for computing Wasserstein barycenters can be divided into three categories.
 The first category relies on discretization of the input measures and the barycentric measure through empirical measures.
-These algorithms may be subdivided into fixed-support [14, Algorithm 1][15, 16], and free-support [14, Algorithm 2][11].
+These algorithms may be subdivided into fixed-support [cuturi2014fast, Algorithm 1][benamou2015iterative, janati2020debiased], and free-support [cuturi2014fast, Algorithm 2][montesuma2023multi].
 The second category relies on parametric models for the input measures, such as Gaussian, or Gaussian mixture models.
-In the Gaussian case, these barycenters are known as Bures-Wasserstein barycenters [17, 18].
-For Gaussian mixtures, [19] shows connections with multi-marginal OT [20] between the components of the input mixtures, while [12] shows that, for components with diagonal covariances, there is an iterative algorithm similar to [14, Algorithm 2].
+In the Gaussian case, these barycenters are known as Bures-Wasserstein barycenters [chewi2020gradient, kroshnin2021statistical].
+For Gaussian mixtures, [delon2020wasserstein] shows connections with multi-marginal OT [pass2015multi] between the components of the input mixtures, while [montesuma2024lighter] shows that, for components with diagonal covariances, there is an iterative algorithm similar to [cuturi2014fast, Algorithm 2].
 The third category of methods for computing Wasserstein barycenters relies on neural nets and comes from the effort to scale OT to measures in high-dimensional spaces with a large number of samples.
-In this sense, [21] proposes parameterizing the barycentric measure through a specific kind of neural net architecture, known as Input Convex Neural Networks (ICNNs) [22].
-Furthermore, [23] proposes a bi-level adversarial learning approach that can be used for general ground costs, which is made more robust by semi-unbalanced OT [24].
-More recently, [25] proposes computing barycenters with normalizing flows.
+In this sense, [korotin2021continuous] proposes parameterizing the barycentric measure through a specific kind of neural net architecture, known as Input Convex Neural Networks (ICNNs) [amos2017input].
+Furthermore, [kolesov2024estimating] proposes a bi-level adversarial learning approach that can be used for general ground costs, which is made more robust by semi-unbalanced OT [gazdieva2024robust].
+More recently, [visentin2025computing] proposes computing barycenters with normalizing flows.
 Current Wasserstein barycenter algorithms mostly focus on scalability with respect to the number of samples $n$, and dimensions $d$.
@@
 There is another level of scalability, oftentimes overlooked, which is with respect the number of input measures $K$.
-So far, only [25] has considered experiments in this context.
-In this regard, methods such as [21, 26] are difficult to scale, as they require $\mathcal{O}(K)$ neural networks to compute the Wasserstein barycenter.
-In this paper we propose a new perspective on discrete Wasserstein barycenters, through the lens of Wasserstein Gradient Flows (WGFs) [27, 28], which is an appealing framework for optimization in the space of probability measures.
+So far, only [visentin2025computing] has considered experiments in this context.
+In this regard, methods such as [korotin2021continuous, korotin2022wasserstein] are difficult to scale, as they require $\mathcal{O}(K)$ neural networks to compute the Wasserstein barycenter.
+In this paper we propose a new perspective on discrete Wasserstein barycenters, through the lens of Wasserstein Gradient Flows (WGFs) [ambrosio2008gradient, santambrogio2017euclidean], which is an appealing framework for optimization in the space of probability measures.
 This new understanding of Wasserstein barycenters has several advantages.
@@
   • We introduce a unified framework for computing Wasserstein barycenters based on gradient flows, which provides a principled and flexible approach to optimizing in the space of probability measures.
-  • We generalize algorithms from the state-of-the-art [14, 11, 12] under a common framework.
+  • We generalize algorithms from the state-of-the-art [cuturi2014fast, montesuma2023multi, montesuma2024lighter] under a common framework.
   • We offer theoretical guarantees for the convergence of our algorithms, which also apply to previous works.
-  • We offer a new theoretical result for the mixture-Wasserstein distance between labeled Gaussian mixtures [19, 12].
-  • We conduct extensive experiments benchmarking existing methods, including neural network-based Wasserstein barycenter algorithms [21, 23, 24, 25], on both toy datasets and domain adaptation  [^1] In contrast with discrete methods, most experiments in neural network solvers involve generative modeling.
+  • We offer a new theoretical result for the mixture-Wasserstein distance between labeled Gaussian mixtures [delon2020wasserstein, montesuma2024lighter].
+  • We conduct extensive experiments benchmarking existing methods, including neural network-based Wasserstein barycenter algorithms [korotin2021continuous, kolesov2024estimating, gazdieva2024robust, visentin2025computing], on both toy datasets and domain adaptation  [^1] In contrast with discrete methods, most experiments in neural network solvers involve generative modeling.
 We thus establish a bridge between these methods and domain adaptation benchmarks, and show that our approach establishes a new state-of-the-art.
@@
 which gives a finite parametrization for $P$. Henceforth we use a hat to indicate empirical measures (e.g., $\hat{P}$).
-OT [2] is a field of mathematics concerned with the transportation of mass at least effort.
-We refer readers to [29] for a computational treatment of the subject, and [30] for applications in machine learning.
+OT [villani2008optimal] is a field of mathematics concerned with the transportation of mass at least effort.
+We refer readers to [peyre2019computational] for a computational treatment of the subject, and [montesuma2024recent] for applications in machine learning.
 Let $P,Q\in\mathcal{P}(\Omega)$, and $c:\Omega^{2}\rightarrow\mathbb{R}$ be a ground-cost.
-OT was originally founded by Monge [31].
+OT was originally founded by Monge [monge1781memoire].
 In this formulation, we seek for a mapping $T:\Omega\rightarrow\Omega$ such that,
 $T_{P\rightarrow Q}^{\star}=\underset{T_{\sharp}P=Q}{\text{arg inf}}\,\int_{\Omega}c(z,T(z))dP(z).$ (2)
-Alternatively, Kantorovich [32] proposed a formulation in terms of a transport plan $\gamma\in\Gamma(P,Q)$, where $\Gamma(P,Q)$ is the set of all joint measures with marginals $P$ and $Q$.
+Alternatively, Kantorovich [kantorovich1942transfer] proposed a formulation in terms of a transport plan $\gamma\in\Gamma(P,Q)$, where $\Gamma(P,Q)$ is the set of all joint measures with marginals $P$ and $Q$.
 In this case,
@@
 where $\theta=\{(\pi_{i}^{(P)},\mu_{i}^{(P)},\Sigma_{i}^{(P)})\}$ are the parameters of the GMM.
-The general theory of OT between GMMs was first presented in [19].
+The general theory of OT between GMMs was first presented in [delon2020wasserstein].
 The main advantage is that, when the OT plan is restricted to the set of GMMs, i.e, $\gamma\in\Gamma(P,Q)\cap\text{GMM}_{\infty}(d)$, there is a discrete equivalent formulation in terms of the GMM components,
@@
 $\omega^{\star}\in\mathbb{R}^{n\times m}$ is a component-to-component OT plan.
-The term $\mathbb{W}_{2}(P_{i},Q_{j})$ is the Wasserstein distance between Gaussian components, which has closed-form [33] in terms of their parameters,
+The term $\mathbb{W}_{2}(P_{i},Q_{j})$ is the Wasserstein distance between Gaussian components, which has closed-form [takatsu2011wasserstein] in terms of their parameters,
 $\mathbb{W}_{2}(P,P^{\prime})^{2}=\lVert\mu-\mu^{\prime}\rVert_{2}^{2}+\text{Tr}(\Sigma+\Sigma^{\prime}-2(\Sigma^{\frac{1}{2}}\Sigma^{\prime}\Sigma^{\frac{1}{2}})^{\frac{1}{2}}).$
@@
 2.3 Wasserstein Barycenter and its Variants
-In the metric setting, the barycenter problem is known as Fréchet [34] or Karcher [35] means.
+In the metric setting, the barycenter problem is known as Fréchet [frechet1948elements] or Karcher [karcher1977riemannian] means.
 In our case, given a finite family of measures $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}$ and a set of barycentric coordinates $\lambda\in\Delta_{K}$, we define the Wasserstein barycenters over $(\mathcal{P}_{p,ac}(\Omega),\mathbb{W}_{p})$ through the following optimization problem,
@@
 In general, equation 7 does not have a closed-form solution.
-However, there are algorithms for computing barycenters for $p=2$, when the measures in $\mathcal{P}$ are either empirical measures [14], Gaussian measures [1, 36] or Gaussian mixtures [19, 12], and a further link can be made to multi-marginal OT [20].
-The work of [14] is of particular interest to us, since theirs was the first algorithm, based on gradient descent, to optimize equation 7 on empirical measures.
+However, there are algorithms for computing barycenters for $p=2$, when the measures in $\mathcal{P}$ are either empirical measures [cuturi2014fast], Gaussian measures [agueh2011barycenters, alvarez2016fixed] or Gaussian mixtures [delon2020wasserstein, montesuma2024lighter], and a further link can be made to multi-marginal OT [pass2015multi].
+The work of [cuturi2014fast] is of particular interest to us, since theirs was the first algorithm, based on gradient descent, to optimize equation 7 on empirical measures.
 The idea is to initialize $z_{0,1}^{(P)},\cdots,z_{0,n}^{(P)}$ randomly (e.g., from a Gaussian measure), and iterate
@@
 where $T_{\gamma_{k}}(z_{i}^{(P)})=n\sum_{j=1}^{n}\gamma_{k,i,j}z_{j}^{(Q_{k})}$ is the barycentric map between $Q_{k}$ and $P$.
-This strategy is reminiscent of the fixed-point approach of [36].
-Furthermore, [11] shows that, for feature-label joint measures, the labels can be propagated through [37],
+This strategy is reminiscent of the fixed-point approach of [alvarez2016fixed].
+Furthermore, [montesuma2023multi] shows that, for feature-label joint measures, the labels can be propagated through [redko2019optimal],
 $y_{\tau+1,i}^{(P)}=(1-\alpha)y_{\tau,i}^{(P)}+\alpha\sum_{k=1}^{K}\lambda_{k}T_{\gamma_{k}}(y_{\tau,i}^{(P)}),$ (9)
@@
 where $P_{t}$ is a curve in $\mathcal{P}_{p,\text{ac}}(\Omega)$, and $\mathbb{F}:\mathcal{P}_{p,\text{ac}}(\Omega)\rightarrow\mathbb{R}$ is a functional over probability measures.
-Usually, these functionals take the following form [38],
+Usually, these functionals take the following form [santambrogio2015optimal],
 $\mathbb{F}(P)=$ $\underbrace{\int G(P(z))dz}_{\mathbb{G}(P)}+\underbrace{\int V(z)dP(z)}_{\mathbb{V}(P)}+$ (13)
@@
 3.3 Flows over Joint Measures
-In [14], the authors considered barycenters for $p=2$, $\Omega=\mathbb{R}^{d}$, and $d(z,z^{\prime})=\lVert z-z^{\prime}\rVert_{2}$.
+In [cuturi2014fast], the authors considered barycenters for $p=2$, $\Omega=\mathbb{R}^{d}$, and $d(z,z^{\prime})=\lVert z-z^{\prime}\rVert_{2}$.
 Therefore, we generalize this setting, since we only need the differentiability of $d$ for our algorithm to work.
@@
 However, for classification, $\mathcal{Y}$ is categorical (e.g., $\mathcal{Y}=\{1,\cdots,n_{\text{classes}}\}$).
-One possible strategy, used in [10] and [39], is fixing the labels and flowing only the features.
+One possible strategy, used in [montesuma2021wasserstein] and [alvarez2021dataset], is fixing the labels and flowing only the features.
 In contrast to these works, we embed $\mathcal{Y}$ into the compact continuous space $\Delta_{K}$, through a one-hot encoding operation.
@@
 where $P_{i,x}=\mathcal{N}(\cdot|\mu_{i}^{(P)},\Sigma_{i}^{(P)})$ is the Gaussian feature marginal.
-This approach was first proposed in [12], but the authors relied on heuristic arguments to justify this modeling choice.
+This approach was first proposed in [montesuma2024lighter], but the authors relied on heuristic arguments to justify this modeling choice.
 We present the following proposition justifying the label term,
@@
 One of the advantages of our proposed method is regularizing the barycenter calculation with internal, interaction, and potential energy functionals.
-This idea was already used in practice by [39] for transfer learning problems.
+This idea was already used in practice by [alvarez2021dataset] for transfer learning problems.
 Here we propose the following functionals,
@@
 3.4 Convergence
-One of the difficulties in analyzing the gradient flow of the functional in equation 14 comes from the fact that $P\mapsto\mathbb{W}_{p}(P,Q)^{p}$ is not geodesically convex in $\mathcal{P}_{p,ac}(\Omega)$ [28, Section 4.4].
+One of the difficulties in analyzing the gradient flow of the functional in equation 14 comes from the fact that $P\mapsto\mathbb{W}_{p}(P,Q)^{p}$ is not geodesically convex in $\mathcal{P}_{p,ac}(\Omega)$ [santambrogio2017euclidean, Section 4.4].
 As a result, we are minimizing a non-convex functional.
@@
 Explicit constants and proofs are available in Section 4 in the Appendix.
-Our analysis relies on a measure-theoretic version of the Polyak Łojasiewicz (PL) inequality [40, 41],
+Our analysis relies on a measure-theoretic version of the Polyak Łojasiewicz (PL) inequality [polyak1964gradient, karimi2016linear],
 $\lVert\nabla\mkern-10.0mu\nabla\mathbb{B}(P)\rVert_{L_{2}(P)}\geq C_{\text{PL}}(\mathbb{B}(P)-\mathbb{B}^{\star}),$ (22)
-which is covered in [42]. Under this assumption, we have the following convergence result,
+which is covered in [chewi2024statistical]. Under this assumption, we have the following convergence result,
 Theorem 3.1.
@@
 Let $P^{\star}\in\mathcal{P}_{2,\text{ac}}(\mathcal{B}(0,R))$ be the barycenter of $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K},Q_{k}\in\mathcal{P}_{2,\text{ac}}(\mathcal{B}(0,R))$ $k=1,\cdots,K$, with barycentric coordinates $\lambda=(\lambda_{1},\cdots,\lambda_{K})\in\Delta_{K}$.
-Assume the inequality 22.
+Assume the inequality 47.
 Then, the following holds,
@@
 Furthermore, we denote by $\hat{Q}_{k}$ the GMM fitted on data from $Q_{k}$ with the expectation-maximization algorithm.
-Our result relies on [43, Theorem 3.1].
+Our result relies on [boite2025differentiable, Theorem 3.1].
 Therefore, we assume,
@@
 In Figure 2, we show the Swiss roll measure $Q_{0}$, alongside four variations obtained via $T_{k,\sharp}Q_{0}$, $k=1,\cdots,4$.
-This example has been used in several state-of-the-art Wasserstein barycenter works [21, 25].
+This example has been used in several state-of-the-art Wasserstein barycenter works [korotin2021continuous, visentin2025computing].
 Originally there are labels associated with the samples of the Swiss-roll measure, corresponding to the position of the points in the underlying manifold.
@@
 First, we compare unsupervised barycenter solvers in Figures 3 (a) through (h).
-Quantitatively, empirical methods, notably [14, Algorithm 2] and our WGF algorithm (cf.
+Quantitatively, empirical methods, notably [cuturi2014fast, Algorithm 2] and our WGF algorithm (cf.
 Algorithm 1) achieve the lowest Wasserstein distance to the ground truth in Figure 2.
 We note that, overall, while the neural network solvers are usually more scalable than discrete methods in terms of number of samples, their optimization problem is more complicated and very sensitive to hyper-parameters.
-(a) CW2B [21].
-(b) NOT [23].
-(c) Discrete [14].
+(a) CW2B [korotin2021continuous].
+(b) NOT [kolesov2024estimating].
+(c) Discrete [cuturi2014fast].
 (d) WGF.
@@
 […]
-(g) U-NOT [24].
-(h) NormFlow [25].
-(i) Discrete [11].
+(g) U-NOT [gazdieva2024robust].
+(h) NormFlow [visentin2025computing].
+(i) Discrete [montesuma2023multi].
 (j) WGF.
@@
 Problem Formulation.
-One of the main applications of Wasserstein barycenters is multi-source domain adaptation (MSDA) [44, 45, 46].
+One of the main applications of Wasserstein barycenters is multi-source domain adaptation (MSDA) [kouw2019review, sun2015survey, pan2009survey].
 In this setting one needs to adapt multiple labeled source measures $Q_{1},\cdots,Q_{K}$ to a single unlabeled target measure $Q_{T}$.
@@
 We obtain the feature extractor by fine-tuning a neural network on the labeled source domain data.
-In addition, similar to [10, 12], we align the barycenter with the target through OT [47].
+In addition, similar to [montesuma2021wasserstein, montesuma2024lighter], we align the barycenter with the target through OT [courty2016optimal].
 More details are available in Section 6 in the Appendix.
 Benchmark Backbone # Samples # Domains # Dim. # Classes
-Office31 ResNet50 [48] 3287 3 2048 31
-BCI-CIV-2a CBraMod [49] 5184 10 200 4
-TEP CNN [50] 17289 6 128 29
-Office Home ResNet101 [48] 15500 4 2048 65
-ISRUC CBraMod [49] 89240 100 512 5
+Office31 ResNet50 [he2016deep] 3287 3 2048 31
+BCI-CIV-2a CBraMod [wang2024cbramod] 5184 10 200 4
+TEP CNN [montesuma2024benchmarking] 17289 6 128 29
+Office Home ResNet101 [he2016deep] 15500 4 2048 65
+ISRUC CBraMod [wang2024cbramod] 89240 100 512 5
 Table 1: Overview of benchmarks used for domain adaptation, sorted by number of samples.
 Experimental Setting.
-We run our experiments on five benchmarks: Office 31 [51], Office Home [52], BCI-CIV-2a [53], ISRUC [54], and TEP [50].
+We run our experiments on five benchmarks: Office 31 [saenko2010adapting], Office Home [venkateswara2017deep], BCI-CIV-2a [brunner2008bci], ISRUC [khalighi2013automatic], and TEP [montesuma2024benchmarking].
 The first two, second two, and last benchmarks correspond to computer vision, neuroscience, and chemical engineering benchmarks, respectively.
@@
 Overall, we compare seven Wasserstein barycenter strategies with ours.
-These methods are: discrete barycenters [14, Algorithm 2], normalizing flows [25], continuous 2-Wasserstein barycenters [21], neural OT barycenters [23], and unbalanced neural OT barycenters [24].
-We also include labeled barycenter strategies [11, 55].
+These methods are: discrete barycenters [cuturi2014fast, Algorithm 2], normalizing flows [visentin2025computing], continuous 2-Wasserstein barycenters [korotin2021continuous], neural OT barycenters [kolesov2024estimating], and unbalanced neural OT barycenters [gazdieva2024robust].
+We also include labeled barycenter strategies [montesuma2023multi, montesuma2024multi].
 A brief review of these methods is available in Section 2 in the Appendix.
 For completeness, we include four other state-of-the-art methods in domain adaptation over embedding vectors.
-Those methods are: WJDOT [56], DaDiL-R and E [11], and GMM-DaDiL [12].
+Those methods are: WJDOT [turrisi2022multi], DaDiL-R and E [montesuma2023multi], and GMM-DaDiL [montesuma2024lighter].
 For each benchmark, we use one domain as the target domain (e.g., Amazon vs.
@@
 Source-Only - 86.40 75.95 50.30 76.63 78.48
-WJDOT [56] - 86.80 76.59 N/A 76.95 86.13
-DaDiL-R [11] - 89.91 77.86 53.41 74.68 86.14
-DaDiL-E [11] - 89.79 78.14 N/A 75.89 85.87
-GMM-DaDiL [12] - 90.63 78.81 57.10 75.47 86.85
-Discrete [14] 81.94 70.42 57.38 70.67 83.81
-NormFlow [25] 85.91 76.73 55.45 78.04 82.89
-CW2B [21] 86.37 76.44 57.25 75.84 85.83
-NOT [23] 86.22 75.36 57.29 76.39 84.93
-U-NOT [24] 86.97 76.85 57.17 77.06 85.43
-Discrete [11] 87.93 77.09 57.43 78.20 86.09
-GMM [12] 88.54 77.87 57.04 74.58 84.67
+WJDOT [turrisi2022multi] - 86.80 76.59 N/A 76.95 86.13
+DaDiL-R [montesuma2023multi] - 89.91 77.86 53.41 74.68 86.14
+DaDiL-E [montesuma2023multi] - 89.79 78.14 N/A 75.89 85.87
+GMM-DaDiL [montesuma2024lighter] - 90.63 78.81 57.10 75.47 86.85
+Discrete [cuturi2014fast] 81.94 70.42 57.38 70.67 83.81
+NormFlow [visentin2025computing] 85.91 76.73 55.45 78.04 82.89
+CW2B [korotin2021continuous] 86.37 76.44 57.25 75.84 85.83
+NOT [kolesov2024estimating] 86.22 75.36 57.29 76.39 84.93
+U-NOT [gazdieva2024robust] 86.97 76.85 57.17 77.06 85.43
+Discrete [montesuma2023multi] 87.93 77.09 57.43 78.20 86.09
+GMM [montesuma2024lighter] 88.54 77.87 57.04 74.58 84.67
 WGF (ours) 88.14 77.83 57.50 79.78 86.21
@@
 Overall, labeled barycenter methods (the four last rows in Table 2) have a clear advantage with respect to the unsupervised methods.
-We argue that label information is pivotal in domain adaptation success, which is consistent with previous research [47, 10, 11, 12].
+We argue that label information is pivotal in domain adaptation success, which is consistent with previous research [courty2016optimal, montesuma2021wasserstein, montesuma2023multi, montesuma2024lighter].
 Among the labeled barycenter methods, our gradient flow framework achieves either the best performance, or remains competitive, surpassing previous methods in MSDA, such as WJDOT and DaDiL in the ISRUC benchmark.
-Figure 4: t-SNE [57] visualization of the barycenter of $\text{Art},\text{Product},\text{Real-World}]$ source domains in the Office home benchmark.
-Colors represent different classes from 11 to $65$.
+Figure 4: t-SNE [maaten2008visualizing] visualization of the barycenter of $\text{Art},\text{Product},\text{Real-World}]$ source domains in the Office home benchmark.
+Colors represent different classes from $1$ to $65$.
 Overall, the classes tend to be more separated when using the repulsion functional $\mathbb{U}$.
@@
 Adding $\mathbb{U}$ leads to classes that are well separated.
-For computer vision experiments we use ResNets [48] as the backbone, which produce $2048$-dimensional embeddings.
+For computer vision experiments we use ResNets [he2016deep] as the backbone, which produce $2048$-dimensional embeddings.
 For that reason, we use the cosine distance $d(x,x^{\prime})=1-\text{cossim}(x,x^{\prime})$ in equation 21 and we fix $\text{margin}=1$.
@@
 Method $\mathcal{X}\times\mathcal{Y}$ $\mathbb{B}$ $\mathbb{B}+\mathbb{V}$ $\mathbb{B}+\mathbb{U}$ $\mathbb{B}+\mathbb{V}+\mathbb{U}$
-Discrete [14] 70.42 N/A N/A N/A
-Discrete [11] 77.09 N/A N/A N/A
+Discrete [cuturi2014fast] 70.42 N/A N/A N/A
+Discrete [montesuma2023multi] 77.09 N/A N/A N/A
 WGF (ours) 77.48 77.07 77.08 77.83
-GMM [12] 77.87 N/A N/A N/A
+GMM [montesuma2024lighter] 77.87 N/A N/A N/A
 WGF-GMM (ours) 77.28 78.28 78.12 78.71
@@
 5 Conclusion
-In this paper we have introduced a new framework for computing Wasserstein barycenters, based on Wasserstein gradient flows [27, 28].
-Our approach addresses scalability issues in discrete barycenter solvers [14], and is capable of regularizing the underlying barycenter through internal, potential, and interaction functionals.
+In this paper we have introduced a new framework for computing Wasserstein barycenters, based on Wasserstein gradient flows [ambrosio2008gradient, santambrogio2017euclidean].
+Our approach addresses scalability issues in discrete barycenter solvers [cuturi2014fast], and is capable of regularizing the underlying barycenter through internal, potential, and interaction functionals.
 We presented gradient flow Algorithms 1 and 2, for empirical and Gaussian mixture measures, respectively.
 We further proved convergence guarantees based on the PL inequality.
-We empirically tested our method against existing discrete [14, 11] and neural-network [21, 23, 24, 25]-based barycenter solvers on computer vision, neuroscience, and chemical engineering benchmarks, demonstrating that our proposed methods consistently achieve state-of-the-art performance.
+We empirically tested our method against existing discrete [cuturi2014fast, montesuma2023multi] and neural-network [korotin2021continuous, kolesov2024estimating, gazdieva2024robust, visentin2025computing]-based barycenter solvers on computer vision, neuroscience, and chemical engineering benchmarks, demonstrating that our proposed methods consistently achieve state-of-the-art performance.
 Our empirical findings (Table 2) further demonstrate that incorporating label information in the OT objective is key for domain adaptation.
@@
 We leave an investigation of this issue for future work.
+Computing Wasserstein Barycenters through Gradient Flows
+Appendix A Introduction
+The goal of this supplementary material is to provide additional background on the main paper, as well as to provide missing proofs, additional experiments, and the complete experimental setting.
+These details were left out of the main paper due to length constraints.
+Organization.
+We divide this supplementary material as follows.
+Section B provides a brief survey on the methods for computing Wasserstein barycenters.
+Section C.1 provides additional background on OT.
+Section C.2 covers additional details on the potential, internal, and interaction energies, as well as the barycenter objectives and their gradients.
+Section D provides missing proofs.
+Finally, section E and F provide additional experiments.
+Appendix B A Brief Survey on Wasserstein Barycenters
+B.1 Introduction
+Given a finite family of measures $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}$, the Wasserstein barycenter problem was first introduced by [agueh2011barycenters] as the following minimization problem,
+$P^{\star}=\underset{P\in\mathcal{P}_{2,\text{ac}}(\Omega)}{\text{arg min}}\,\biggr\{\mathbb{B}_{\mathcal{Q}}=\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2}(P,Q_{k})^{2}\biggr\}.$ (24)
+This minimization problem is the analogous, in Wasserstein space, to the Euclidean barycenter problem $x\mapsto\sum_{k}\lambda_{k}d(x,x_{k})^{2}$, which is made possible since $(\mathcal{P}_{2,\text{ac}}(\Omega),\mathbb{W}_{2})$ is a metric space.
+Solving problem 24 is, in general, challenging.
+Except for the Gaussian case, there are no closed-form solutions.
+In the following, we review a few strategies for computing Wasserstein barycenters.
+B.2 Discrete Barycenter Solvers
+Discrete barycenter solvers are based on the empirical approximation of the underlying measures, that is,
+$\hat{P}(z)=\sum_{i=1}^{n}a_{i}\delta(z-z_{i}^{(P)}).$ (25)
+Here, $\{z_{i}^{(P)}\}_{i=1}^{n}$ is the support of $\hat{P}$ and $a=\{a_{i}:a_{i}\geq 0\text{, and }\sum a_{i}\}_{i=1}^{n}$ are the sample weights.
+There are two prevalent strategies when discretizing $P$.
+First, one can fix a grid/bins over the underlying space $\Omega$, and compute $a_{i}=P(z_{i}^{(P)})$ as the bin weight.
+This strategy is equivalent to estimating the density of $P$ over $\Omega$.
+Second, we can sample $z_{i}^{(P)}\overset{\text{i.i.d.}}{\sim}P$, in which case $a_{i}=\nicefrac{{1}}{{n}}$.
+In this paper, we deal primarily with free-support Wasserstein barycenters.
+Arguably, the first free-support algorithm was proposed by [cuturi2014fast].
+Their strategy comes from plugging equation 25 into 24,
+$Z^{(P),\star}=\underset{\begin{subarray}{c}z_{1}^{(P)},\cdots,z_{n}^{(P)}\in\Omega,\\ \gamma_{k}\in\Gamma(\hat{P},\hat{Q}_{k})\end{subarray}}{\text{arg min}}\,\biggr\{\mathbb{B}_{\mathcal{Q}}(P)=\sum_{k=1}^{K}\lambda_{k}\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}\lVert z_{i}^{(P)}-z_{j}^{(Q_{k})}\rVert_{2}^{2}\biggr\}.$ (26)
+At this point, one should compare equation 26 with equation 17 in the main paper.
+Cuturi and Doucet [cuturi2014fast] propose solving this equation via a block coordinate descent strategy.
+Indeed, assume $\gamma_{k}=\text{OT}(P,Q_{k})$ has been calculated (e.g., through linear programming), then,
+$\nabla_{z_{i}^{(P)}}\mathbb{B}_{\mathcal{Q}}(P)=2\sum_{k=1}^{K}\lambda_{k}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}(z_{i}^{(P)}-z_{j}^{(Q_{k})})=\dfrac{2}{n}z_{i}^{(P)}-2\sum_{k=1}^{K}\lambda_{k}\sum_{j=1}^{n}\gamma_{k,i,j}z_{j}^{(Q_{k})}.$
+As a consequence, $\nabla_{z_{i}^{(P)}}\mathbb{B}_{\mathcal{Q}}(P)=0$ leads to an update rule in terms of the barycentric projection of $P$ to $Q_{k}$, which is a discrete approximation of the Monge map [deb2021rates].
+An important question comes when the samples from $P$ are feature-label joints. This is question was first raised by [montesuma2023multi], who proposed to use the ground-cost,
+$d(z,z^{\prime})=\sqrt{\lVert x-x^{\prime}\rVert_{2}^{2}+\beta\lVert y-y^{\prime}\rVert_{2}^{2}},$ (27)
+where $y$ and $y^{\prime}$ are one-hot encoded labels, and $\beta\geq 0$ strikes a balance between the features and labels terms.
+As we show in the main paper, since this ground-cost is a metric, computing the OT cost with $d^{2}$ as the ground-cost is equivalent to computing the 22-Wasserstein distance on $\mathcal{X}\times\mathcal{Y}$.
+The first paper to remark these computations was [montesuma2023multi], who proposed to couple these mappings,
+$x_{i,\tau+1}^{(P)}=\sum_{k=1}^{K}\lambda_{k}T_{P\rightarrow Q_{k}}^{\star}(x_{i,\tau}^{(P)})\text{, and, }y_{i,\tau+1}^{(P)}=\sum_{k=1}^{K}\lambda_{k}T_{P\rightarrow Q_{k}}^{\star}(y_{i,\tau}^{(P)}).$ (28)
+These iterations have a few nice properties with respect to the label term.
+First, the mappings $T_{P\rightarrow Q_{k}}^{\star}(y_{i,\tau}^{(P)})$ correspond to label propagation terms first defined in [redko2019optimal].
+Second, as a consequence of the first point, the iterations respect the constraint $y_{i,\tau}^{(P)}\in\Delta_{n_{\text{classes}}}$.
+Third, these iterations can be understood as fixed-point iterations in the sense of [alvarez2016fixed].
+The authors use this algorithm as the building block of their dictionary learning strategy.
+The main drawback of the iterations in equation 28 is when the input measures have imbalanced classes.
+As we show in the additional experiments in Section E.2, OT, due the mass preservation constraints, is forced to match samples from different classes, leading to deformed barycenters with fuzzy labels (see Figure 8).
+B.3 Neural Barycenter Solvers
+In this section, we cover a collection of methods that use neural networks as the foundation of computing Wasserstein barycenters.
+These methods are oftentimes referred to as ”continuous”, thus relying on the idea that the calculated barycenter is a continuous measure.
+Note that this is different than claiming that these barycenters solve a continuous problem.
+In fact, most methods are formulated in terms of expectations of the involved measures, which are approximated through mini-batching.
+Generative Modeling.
+One of the first methods was proposed by [cohen2020estimating], who proposed viewing equation 24 through the lens of generative modeling.
+Let $g_{\theta}$ denote a neural network with parameters $\theta$.
+Then,
+$\theta^{\star}=\underset{\theta\in\Theta}{\text{arg min}}\,\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2}(P_{\theta},Q_{k})^{2},$
+where $P_{\theta}=g_{\theta,\sharp}P_{0}$ is the push-forward of a latent measure $P_{0}$ by the generator $g_{\theta}$.
+Therefore, the Wasserstein barycenter is parametrized via the weights of $g_{\theta}$.
+In comparison, methods such as [cuturi2014fast] and [montesuma2023multi] parametrize the barycenter through its samples.
+The underlying idea is sampling from each $Q_{k}$ and $P_{0}=\mathcal{N}(0,\sigma^{2}I)$.
+Input Convex Neural Nets. More refined settings exploit the properties of the 2-Wasserstein distance with the squared Euclidean ground-cost. Indeed, note that,
+$\dfrac{1}{2}\mathbb{W}_{2}(P,Q)^{2}=\underbrace{\int\dfrac{\lVert z\rVert_{2}}{2}dP(z)+\int\dfrac{\lVert z\rVert_{2}}{2}dQ(z)}_{\mathbb{C}(P,Q)}-\underset{f,g\in\Phi}{\text{min}}\,\int f(z)dP(z)+\int g(z)dQ(z),$ (29)
+where $(f,g)\in\Phi=\{f(x)+g(y)\leq\lVert x-y\rVert_{2}^{2}\}$ are called potentials. The main insight of [makkuva2020optimal] is rewritting the minimization problem in the r.h.s. of equation 29 as,
+$\underset{f\in\text{CVX}}{\text{sup}}\,\underset{g\in\text{CVX}}{\text{inf}}\,\underbrace{-\int f(z)dP(z)-\int(\langle z,\nabla g(z)\rangle-f(\nabla g(z)))dQ(z)}_{V_{P,Q}(f,g)}$
+Based on these equations, [fan2020scalable] propose solving,
+[…]
+Here, samples in the Wasserstein barycenter are generated through the forward pass $h(z)$, where $z\sim P_{0}=\mathcal{N}(0,\sigma^{2}I)$.
+This is a min-max-min problem, which, as with many problems involving neural nets (e.g., generative adversarial nets [goodfellow2014generative]), is subject to saddle points.
+An important aspect of equation 30 is the minimization with respect convex functions $f_{k}$ and $g_{k}$.
+Indeed, these are the potentials of the transportation problem between the barycenter $P$ and each $Q_{k}$.
+More specifically, In light of Brenier’s theorem [brenier1991polar], $\nabla g_{k}$ is the Monge map from $Q_{k}$ to the barycenter $P$.
+These elements motivate [fan2020scalable] using ICNNs, as in [makkuva2020optimal], for the optimization problem in equation 30.
+Continuous 2-Wasserstein Barycenter.
+We now cover a family of methods relying on the fixed-point view of Wasserstein barycenters introduced by [alvarez2016fixed].
+Indeed, these authors show that the Wasserstein barycenter is a fixed point of the operator $\Psi:P\mapsto(\sum_{k}\lambda_{k}T_{P\rightarrow Q_{k}})_{\sharp}P$.
+This is actually the principle behind most of the discrete methods described in the previous section.
+Now, with respect to OT potentials, if $\Psi(P)=P$, then one has,
+$\sum_{k=1}^{K}\lambda_{k}T_{Q_{k}\rightarrow P}^{\star}(z)=x\overset{(a)}{\implies}\sum_{k=1}^{K}\lambda_{k}\nabla g_{k}(z)=z\overset{(b)}{\implies}\sum_{k=1}^{K}\lambda_{k}g_{k}(z)=\dfrac{\lVert z\rVert_{2}^{2}}{2}+c$
+where implication (a) follows from Brenier’s theorem, and $g_{k}$ is the potential of the transportation problem from $Q_{k}$ to $P$, and implication (b) follows by integration.
+Here, since $c$ is an integration constant, one can conveniently take $c=0$.
+Now, denote $\text{Congruent}(\lambda)=\{(g_{1},\cdots,g_{K}):\sum_{k}\lambda_{k}g_{k}(z)=\nicefrac{{\lVert z\rVert_{2}^{2}}}{{2}}\}$ be the set of congruent potentials $(g_{1},\cdots,g_{K})$ with respect to coordinates $\lambda$.
+Using the dual formulation in equation 29 and restricting the set of potentials to congruent potentials,
+$\mathbb{B}_{\mathcal{Q}}(P)=\sum_{k=1}^{K}\lambda_{k}\mathbb{E}_{z\sim Q_{k}}\biggr[\dfrac{\lVert z\rVert_{2}^{2}}{2}\biggr]+\mathbb{E}_{z\sim P}\biggr[\dfrac{\lVert z\rVert_{2}^{2}}{2}\biggr]-\underset{\{g_{k}\}_{k=1}^{K}\in\text{Congruent}(\lambda)}{\text{min}}\,\biggr(\sum_{k=1}^{K}\lambda_{k}(\mathbb{E}_{z\sim Q_{k}}[f_{k}(z)]+\mathbb{E}_{z\sim P}[g_{k}])(z)\biggr).$
+Now, taking the expectation on the congruence constraint, we have,
+$\sum_{k=1}^{K}\lambda_{k}\mathbb{E}_{z\sim P}[g_{k}(z)]=\mathbb{E}_{z\sim P}\biggr[\dfrac{\lVert z\rVert_{2}^{2}}{2}\biggr],$
+which allowed [korotin2021continuous] to simplify $\mathbb{B}_{\mathcal{Q}}(P)$ into,
+$\mathbb{B}_{\mathcal{Q}}(P)=\sum_{k=1}^{K}\lambda_{k}\mathbb{E}_{z\sim Q_{k}}\biggr[\dfrac{\lVert z\rVert_{2}^{2}}{2}\biggr]-\underset{\{g_{k}\}_{k=1}^{K}\in\text{Congruent}(\lambda)}{\text{min}}\,\sum_{k=1}^{K}\lambda_{k}\mathbb{E}_{z\sim Q_{k}}[g_{k}(z)],$ (31)
+which does not involve terms with respect to the barycentric measure $P$.
+While the approach of [korotin2021continuous] simplifies the barycenter objective to expectations with respect to the known measures $Q_{k}$, it introduces a challenging constraint – the congruence of potentials.
+This constraint is challenging, as it is a functional inequality over the whole space $\Omega$.
+Furthermore, it adds to the already constrained setting of fitting convex functions.
+[korotin2021continuous] proposes solving it approximately, by penalizing potentials that do not meet the congruence condition.
+There is a circular dependence here, as one needs an estimate of the Wasserstein barycenter to estimate the constraint violation.
+Neural (Weak) Optimal Transport. Weak OT [gozlan2017kantorovich, backhoff2019existence] comes from the interpretation of the Kantorovich formulation as,
+$\gamma^{\star}=\underset{\gamma\in\Gamma(P,Q)}{\text{arg inf}}\,\int_{\Omega}C(z,\gamma(\cdot|z))dP(z)\text{, where, }C(z,Q)=\int_{\Omega}c(z,z^{\prime})dQ(z^{\prime}),$ (32)
+this formulation allows for more general costs $C:\Omega\times\mathcal{P}(\Omega)\rightarrow\mathbb{R}$, i.e., costs that take a sample and a probability measure as inputs.
+In this sense, [kolesov2024estimating] extends equation 31 to,
+$\mathbb{B}_{\mathcal{Q}}(P)=\sum_{k=1}^{K}\lambda_{k}\biggr(\mathbb{E}_{z\sim Q_{k}}[C_{k}(z,\gamma(\cdot|z))]-\mathbb{E}_{z\sim Q_{k}}[\mathbb{E}_{z^{\prime}\sim\gamma(\cdot|z)}(g_{k}(z^{\prime}))]\biggr),$ (33)
+which holds for congruent potentials $\sum_{k}\lambda_{k}g_{k}=0$.
+As with equation 31, $\mathbb{B}_{\mathcal{Q}}$ does not depend explicitly on the unknown barycentric measure $P$.
+As such, the authors compute the barycenter through a max-min optimization problem,
+$\underset{\{g_{1},\cdots,g_{K}:\sum_{k}\lambda_{k}g_{k}=0\}}{\text{max}}\,\underset{\{\gamma_{1},\cdots,\gamma_{K}:\gamma_{k}(\cdot|z)\in\Gamma(P_{k})\}}{\text{min}}\,\biggr\{\sum_{k=1}^{K}\lambda_{k}\biggr(\mathbb{E}_{z\sim Q_{k}}[C_{k}(z,\gamma_{k}(\cdot|z))]-\mathbb{E}_{z\sim Q_{k}}[\mathbb{E}_{z^{\prime}\sim\gamma_{k}(\cdot|z)}(g_{k}(z^{\prime}))]\biggr)\biggr\}.$
+This extension opens up new possibilities for weak OT barycenters. For instance,
+$C_{\text{classical}}(x,\gamma)=\dfrac{1}{n}\sum_{j=1}^{n}c(z_{i}^{(P)},z_{j}^{(Q)})\quad C_{\text{KL}}(x,\gamma)=\dfrac{1}{n}\sum_{j=1}^{n}c(z_{i}^{(P)},z_{j}^{(Q)})+\epsilon\text{KL}(\mathcal{N}(\mu(z),\sigma(z))|P_{0}),$
+where $C_{\text{classical}}$ allows us to retrieve the original OT problem.
+$C_{\text{KL}}$ is an alternative, when modeling $\gamma(\cdot|z)=\mathcal{N}(\mu(z),\sigma(z))$.
+On top of this reformulation, the authors in [kolesov2024estimating] propose approximating $\gamma(\cdot|z)$ through a mapping, as in [korotin2022neural], that is,
+$\gamma(\cdot|z)=T(z,s),\text{ subject to }s\sim\mathbb{S},$
+where $\mathbb{S}$ is some prior measure, such as $\mathbb{S}=\mathcal{N}(0,\text{I})$.
+In this context, [gazdieva2024robust] extends this idea in the semi-unbalanced OT setting [liero2018optimal], where one of the mass preservation constraint of one of the marginals of $\gamma$ is relaxed.
+More specifically, given $\xi:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$,
+$D_{\xi}(P||Q)=\int_{\Omega}\xi\biggr(\dfrac{P(z)}{Q(z)}\biggr)dQ(z)\text{ if }P\ll Q\text{ and }+\infty\text{ otherwise},$
+where $P\ll Q$ means that $P$ is absolutely continuous with respect to $Q$.
+We denote $\bar{\xi}(z^{\prime})=\text{sup}_{z\in\mathbb{R}}\{zz^{\prime}-\xi(z)\}$ to be its convex conjugate.
+With that in mind, [gazdieva2024robust] shows, in Theorem 1 and Corollaries 1 and 2, that minimizing,
+$\underset{P\in\mathcal{P}_{2,\text{ac}}}{\text{arg min}}\,\sum_{k=1}^{K}\lambda_{k}\biggr(\int_{\Omega}\int_{\Omega}c(z,z^{\prime})d\gamma(z,z^{\prime})+D_{\xi}(\gamma(\cdot|z)||Q_{k})\biggr),$
+is equivalent to the following max-min problem,
+$\underset{\{g_{1},\cdots,g_{K}:\sum_{k}\lambda_{k}g_{k}=0\}}{\text{max}}\,\underset{\{\gamma_{1},\cdots,\gamma_{K}:\gamma_{k}(\cdot|z)\in\Gamma(P_{k})\}}{\text{min}}\,\biggr\{\sum_{k=1}^{K}\lambda_{k}\biggr(\mathbb{E}_{z\sim Q_{k}}\biggr[-\bar{\xi}_{k}\biggr(-\mathbb{E}_{z^{\prime}\sim\gamma_{k}(\cdot|z)}[c_{k}(z,z^{\prime})-g_{k}(z^{\prime})]\biggr)\biggr]\biggr)\biggr\}.$
+Normalizing Flows.
+This approach was proposed by [visentin2025computing], and relies on the analysis of [gangbo1996geometry].
+Their main idea is re-writing the Wasserstein distance in terms of functions $f,g$,
+$\mathbb{W}_{p}(P,Q)^{p}=\underset{f\in\text{Borel}(P_{0},P),g\in\text{Borel}(P_{0},Q)}{\text{inf}}\,\lVert f-g\rVert_{L_{p}(P_{0})}^{p},$ (34)
+where $\text{Borel}(P,Q)$ is the set of Borel functions such that $f_{\sharp}P=Q$.
+Naturally, a solution $(f^{\star},g^{\star})$ to the previous problem gives the transport map $T=g^{\star}\circ f^{-1}$, where $f^{-1}$ is the inverse of $f^{\star}$.
+Furthermore, $P_{0}$ is an arbitrary latent measure, which further links the approach with generative modeling (i.e., take $P_{0}=\mathcal{N}(0,\text{I})$).
+Based on equation 34, the insight of [visentin2025computing] is modeling $f$ and $g$ through a conditional function $f(z,s)$, $s\in\{0,1\}$.
+The Wasserstein barycenter problem is a straightforward generalization of the $2$ measures case.
+For instance, the authors consider $S=\{1,\cdots,K\}$, which means that,
+$\mathbb{B}_{\mathcal{Q}}(P)=\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{p}(P,Q_{k})^{p}=\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{p}(g_{\sharp}P_{0},Q_{k})=\sum_{k=1}^{K}\lambda_{k}\lVert g-f(z,k)\rVert_{L_{p}(P_{0})}^{p},$
+where $g$ is the push-forward of $P_{0}$ to the barycenter $P^{\star}$, and $f(z,k)$ is the push-forward of $Q_{k}$ to $P^{\star}$. As the authors show in [visentin2025computing, Theorem 3.3],
+$g(z)=\sum_{k=1}^{K}\lambda_{k}f(z,k).$
+The main challenge in [visentin2025computing] is formulating a conditional model $f:\Omega\times S$.
+As the authors discuss, one possibility would be using $K=|S|$ different neural nets for modeling $f(z,\cdot)$, which is similar to what is done, for instance, in the neural net methods.
+The authors take a different route and use conditional normalizing flows [kobyzev2020normalizing] to model these functions.
+B.4 Future Directions
+From the previous discussion on neural net-based methods for computing Wasserstein barycenters, we can identify a few common design choices.
+For instance, [korotin2021continuous, kolesov2024estimating] and [gazdieva2024robust] use $\mathcal{O}(K)$ neural nets for computing Wasserstein barycenters.
+This design choice results in a parametrization of the underlying barycenter that grows with the number of input measures.
+The only neural net-based method that differs from this idea is [visentin2025computing], which uses a conditional model $f:\Omega\times S$.
+Meanwhile, as we cover in our experiments (C.f., Section 4 in the main paper), there is a clear advantage of incorporating labels in the ground-cost when these are available.
+Unfortunately, it is not clear how to integrate label information in the existing methods.
+For instance, one could consider using the continuous parametrization in terms of logits $\ell$, as we described in Section 3.3.
+However, there are a few caveats,
+  1. 1.
+    Normalizing flows require invertible transformations. However, the softmax operation for getting the labels, i.e., $y=\text{softmax}(\ell)$, is not invertible.
+  2. 2.
+    Using ICNNs, one can enforce the convexity of $z=(x,\ell)$. However, applying the softmax breaks the convexity requirement.
+As a result, adapting neural net based methods for the joint $\mathcal{X}\times\mathcal{Y}$ space is challenging and not straightforward. We leave this question for future work.
+Appendix C Additional Background
+C.1 Calculus in Wasserstein Spaces
+In this section, we consider $\Omega=\mathbb{R}^{d}$, $d(z,z^{\prime})=\lVert z-z^{\prime}\rVert_{2}$ and $p=2$ fixed.
+While our algorithm is more general than this setting, we use these restrictions as they are the most analyzed by current literature.
+We leave the theoretical results for general metrics and $p$ to future work.
+Notation.
+For the following theoretical analysis, we use the following notation.
+Given a measurable map $f,g:\Omega\rightarrow\Omega$.
+Given a measure $P\in\mathcal{P}_{2,\text{ac}}(\Omega)$, we denote,
+$\lVert f\rVert_{L_{2}(P)}=\int_{\Omega}\lVert f(x)\rVert_{2}^{2}dP(x).$
+Likewise, we have a notion of inner product,
+$\langle f,g\rangle_{P}=\int_{\Omega}\langle f(x),g(x)\rangle_{\Omega}dP(x).$
+Definition C.1.
+Given a functional $\mathbb{F}:\mathcal{P}_{2,\text{ac}}(\Omega)(\mathbb{R}^{d})\rightarrow\mathbb{R}$, we call $\delta\mathbb{F}(P):\mathbb{R}^{d}\rightarrow\mathbb{R}$, if it exists, the unique (up to additive constants) function such that
+$\dfrac{d}{d\varepsilon}\mathbb{F}(P+\varepsilon\chi)\biggr|_{\epsilon=0}=\int\delta\mathbb{F}(P)d\chi$
+for every perturbation $\chi$ such that, at least for $\varepsilon\in[0,\varepsilon_{0}]$, the measure $P+\varepsilon\chi\in\mathcal{P}_{p,\text{ac}}(\mathbb{R}^{d})$.
+This functional is called the first variation of $\mathbb{F}$ at $P$.
+Following [chewi2024statistical], given a curve of densities $\{P_{t}\}_{t\geq 0}$, we can take $\chi=\partial_{t}\mathbb{F}(P_{t})$ so that $\partial_{t}\mathbb{F}(P_{t})=\int\delta\mathbb{F}(P_{t})\partial_{t}P_{t}$.
+From Riemannian geometry, we can define the Wasserstein gradient to $P_{t}$ as,
+$\partial_{t}\mathbb{F}(P_{t})=\langle\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t}),v_{t}\rangle_{P_{t}},$ (35)
+where $v_{t}\in\mathcal{T}_{P_{t}}\mathcal{P}_{p,\text{ac}}(\mathbb{R}^{d})$.
+Note, here, that $\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t}):\Omega\rightarrow\Omega$ is a vector field.
+In comparison, the first variation $\delta\mathbb{F}$ is a scalar field.
+These two objects are connected by the following equation [chewi2024statistical, Proposition 5.10],
+$\nabla\mkern-10.0mu\nabla\mathbb{F}(P)=\nabla\delta\mathbb{F}(P),$ (36)
+Meanwhile, recall that, from dynamic OT, the WGF satisfy the continuity equation,
+$\partial_{t}P_{t}+\text{div}(P_{t}v_{t})=0,$
+where $\text{div}(v)=\nabla\cdot v$ is the divergence of a vector field.
+Therefore, taking the direction of steepest descent of the functional $\mathbb{F}$, $v_{t}=-\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t})$, leads to the following connection between $\partial_{t}\mathbb{F}(P_{t})$ and $\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t})$,
+$\partial_{t}\mathbb{F}(P_{t})=-\lVert\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t})\rVert_{L_{2}(P_{t})}^{2},$ (37)
+C.2 Wasserstein Gradient of Functionals
+In this section, we give an overview on the Wasserstein gradient of the terms used in equation 16.
+We refer readers to [chewi2024statistical, Chapter 5] and [santambrogio2017euclidean, Section 4] for further details.
+Potential Energy, $\mathbb{V}$. For reference, this functional corresponds to,
+$\mathbb{V}(P)=\mathbb{E}_{z\sim P}[V(z)]=\int V(z)dP(z),$
+for some potential function $V:\Omega\rightarrow\mathbb{R}$. From the definition of first variation,
+$\mathbb{V}(P+\epsilon\chi)=\int V(z)dP(z)+\epsilon\int V(z)d\chi(z),$
+then, taking the derivative with respect to $\epsilon$,
+$\dfrac{d}{d\epsilon}\mathbb{V}(P+\epsilon\chi)=\int V(z)d\chi,$
+which means that $V=\delta\mathbb{V}(P)$. As a consequence, the Wasserstein gradient of $\mathbb{V}$ is,
+$\nabla\mkern-10.0mu\nabla\mathbb{V}=\nabla V.$
+which is a vector field over $\Omega$.
+Since $\mathbb{V}$ is linear in $P$, it is convex.
+Furthermore, its continuity with respect to the weak convergence depends on the regularity of $V$.
+For instance, if $V$ is bounded, $\mathbb{V}$ is continuous.
+If $V$ is lower semi-continuous and bounded from below, then $\mathbb{V}$ is semi-continuous.
+Internal Energy, $\mathbb{G}$. The internal energy takes the form,
+$\mathbb{G}(P)=\int G(P(z))dz,$
+where $P(z)$ is understood as the density of the measure $P$, with respect the Lebesgue measure. Let us assume $\chi$ admits a density with respect the Lebesgue measure as well. In this case,
+$\mathbb{G}(P+\epsilon\chi)=\int G(P(z)+\epsilon\chi(z))dz.$
+Taking the derivative under the integral sign,
+$\dfrac{d}{d\epsilon}\mathbb{G}(P+\epsilon\chi)\biggr|_{\epsilon=0}=\int G^{\prime}(P(z))\chi(z)dz=\int G^{\prime}(P(z))d\chi(z),$
+which means that $\delta\mathbb{G}=G^{\prime}(P)$ and $\nabla\mkern-10.0mu\nabla\mathbb{G}=\nabla G^{\prime}(P(z))$.
+Interaction Energy, $\mathbb{U}$. Here, let us assume $U(z,z^{\prime})$ symmetric for simplicity,
+$\mathbb{U}(P)=\int\int U(z,z^{\prime})dP(z)dP(z^{\prime}),$
+hence,
+$\mathbb{U}(P+\epsilon\chi)=\int\int U(z,z^{\prime})dP(z)dP(z^{\prime})+2\epsilon\int\int U(z,z^{\prime})dP(z)d\chi(z^{\prime})+\mathcal{O}(\epsilon^{2}).$
+Again, taking derivatives with respect $\epsilon$,
+$\dfrac{d}{d\epsilon}\mathbb{U}(P+\epsilon\chi)\biggr|_{\epsilon=0}=2\int\biggr(\int U(z,z^{\prime})dP(z)\biggr)d\chi(z^{\prime}),$
+which means that,
+$\delta\mathbb{U}(P)=2\int U(z,z^{\prime})dP(z).$
+When $U(z,z^{\prime})=h(z-z^{\prime})$ for an even $h$, $\int U(z,z^{\prime})dP(z)=\int h(z-z^{\prime})P(z)dz=h\star P$, i.e., the convolution of $h$ and $P$.
+Furthermore, $\nabla\mkern-10.0mu\nabla\mathbb{U}=(\nabla h)\star P$
+Barycenter Functional, $\mathbb{B}$.
+For the barycenter functional, we can derive its first variation from the 22-Wasserstein distance.
+Assume there exists an OT map $T_{P\rightarrow Q}^{\star}$.
+In this case,
+$\nabla\mkern-10.0mu\nabla\mathbb{W}_{2}(P,Q)^{2}=2(\text{Id}-T_{P\rightarrow Q}^{\star})$
+Therefore, given $\mathcal{Q}=\{Q_{1},\cdots,Q_{K}\}$, it is straightforward to obtain:
+$\nabla\mkern-10.0mu\nabla\mathbb{B}_{\mathcal{Q}}(P)$ $=\nabla\mkern-10.0mu\nabla\biggr(\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2}(P,Q_{k})^{2}\biggr)=2\sum_{k=1}^{K}\lambda_{k}(\text{Id}-T_{P\rightarrow Q_{k}}^{\star}).$
+A Connection with [cuturi2014fast, Algorithm 2].
+Let us now focus on the empirical case with an Euclidean ground-cost.
+Assume $\gamma^{\star}$ computed between $\hat{P}$ and $\hat{Q}$, which means that,
+$\mathbb{W}_{2}(\hat{P},\hat{Q})^{2}=\sum_{i=1}^{n}\sum_{j=1}^{m}\gamma_{ij}^{\star}\lVert z_{i}^{(P)}-z_{j}^{(Q)}\rVert_{2}^{2}\text{, and }\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})=\sum_{k=1}^{K}\lambda_{k}\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}^{\star}\lVert z_{\tau,i}^{(P)}-z_{j}^{(Q_{k})}\rVert_{2}^{2}.$
+We can compute the Wasserstein gradient $\nabla\mkern-10.0mu\nabla\mathbb{B}_{\mathcal{Q}}$ by computing $\nabla_{z_{\tau,i}^{(P)}}\mathbb{B}_{\mathcal{Q}}$, that is,
+$\nabla_{z_{\tau,i}^{(P)}}\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})=2\sum_{k=1}^{K}\lambda_{k}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}^{\star}(z_{\tau,i}^{(P)}-z_{j}^{(Q_{k})})=\dfrac{2}{n}z_{\tau,i}^{(P)}-2\sum_{k=1}^{K}\lambda_{k}\sum_{j=1}^{K}\gamma_{k,i,j}^{\star}z_{j}^{(Q_{k})}.$
+From this last equation, we can use the barycentric mapping definition,
+$T_{P\rightarrow Q_{k}}^{\star}(z_{i,\tau}^{P})=n\sum_{j=1}^{n_{k}}\gamma_{k,i,j}^{\star}z_{j}^{(Q_{k})},$ (38)
+to write,
+$\nabla_{z_{\tau,i}^{(P)}}\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})=\dfrac{2}{n}\biggr(z_{\tau,i}^{(P)}-\sum_{k=1}^{K}\lambda_{k}T_{P\rightarrow Q_{k}}^{\star}(z_{\tau,i}^{(P)})\biggr),$
+which means we can write the gradient flow update rule as,
+$z_{\tau+1,i}^{(P)}=\biggr(1-\dfrac{2\alpha}{n}\biggr)z_{i,\tau}^{(P)}+\dfrac{2\alpha}{n}\sum_{k=1}^{K}\lambda_{k}T_{P\rightarrow Q_{k}}^{\star}(z_{i,\tau}^{(P)}),$
+Which is the update rule in [cuturi2014fast, Algorithm 2] with $\theta=\nicefrac{{2\alpha}}{{n}}$.
+Furthermore, we can also instantiate the fixed-point algorithm of [montesuma2023multi], for $\alpha=n/2$.
+We now highlight two differences of our work with respect these two strategies.
+First, our setting is more general.
+In fact, our algorithm is capable of handling ground metrics other than the squared Euclidean metric.
+Furthermore, our framework couples the barycenter objective with other functionals.
+Second, while [cuturi2014fast] assume that all samples of $\mathcal{Q}=\{Q_{1},\cdots,Q_{K}\}$ are available at once, our algorithm assumes that these measures are available through sampling.
+This treatment is oftentimes called continuous in the literature [korotin2021continuous, korotin2022wasserstein], however, one still relies on samples from these measures rather than some truly continuous form.
+C.3 Domain Adaptation
+In this section, we offer a brief introduction to domain adaptation in classification problems which are the main application we explore in the main paper.
+Statistical learning theory [vapnik1999overview] deals with the problem of learning a function $h:\mathcal{X}\rightarrow\mathcal{Y}$, such that it minimizes the risk of making a mistake, that is,
+$h^{\star}=\underset{h\in\mathcal{H}}{\text{arg inf}}\,\biggr\{\mathcal{R}_{Q}(h)=\int_{\mathcal{X}\times\mathcal{Y}}\mathcal{L}(y,h(x))dQ(x,y)\biggr\},$ (39)
+where $P\in\mathcal{P}(\mathcal{X}\times\mathcal{Y})$ is a probability measure, and $\mathcal{H}$ is a family of functions from $\mathcal{X}$ to $\mathcal{Y}$.
+$\mathcal{R}_{Q}(h)$ denotes the risk of $h$ under $Q$.
+Equation 39 interprets the learning problem as a minimization problem.
+In practical scenarios, one has a dataset $\{x_{i}^{(Q)},y_{i}^{(P)}\}_{i=1}^{n}$, $(x_{i}^{(P)},y_{i}^{(Q)})\overset{i.i.d.}{\sim}P$.
+As a result, one can estimate the empirical risk, and consequently minimize it,
+$\hat{h}=\underset{h\in\mathcal{H}}{\text{arg inf}}\,\biggr\{\hat{\mathcal{R}}_{Q}(h)=\dfrac{1}{n}\sum_{i=1}^{n}\mathcal{L}(y_{i}^{(Q)},h(x_{i}^{(Q)}))\biggr\},$ (40)
+which is the foundation of the Empirical Risk Minimization (ERM) framework. In this sense, a solution $\hat{h}$ to equation 40 is said to generalize if it achieves low risk $\mathcal{R}_{Q}(h)$.
+There is a fundamental assumption at play in the theory of generalization under the ERM framework.
+Indeed, both $\mathcal{R}_{P}$ and $\hat{\mathcal{R}}_{P}$ evaluate the risk under a fixed probability measure $P$.
+However, machine learning models are oftentimes confronted with data following different, but related probability measures [quinonero2022dataset], a process known as distribution shift.
+The main tool for tackling distribution shift is domain adaptation.
+This problem considers now two datasets, namely, a labeled source dataset $\{x_{i}^{(Q_{S})},y_{i}^{(Q_{S})}\}_{i=1}^{n_{S}}$, and an unlabeled target dataset $\{x_{j}^{(Q_{T})}\}_{j=1}^{n_{T}}$.
+The goal is to obtain a function $h$ that achieves low risk under $Q$, with the available data.
+A natural extension to this problem is multi-source domain adaptation, where multiple source domains are available, i.e., $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}$, each with its own associated dataset $\{(x_{i}^{(Q_{k})},y_{i}^{(Q_{k})})\}_{i=1}^{n_{k}}$.
+Source-Only Training.
+A straightforward baseline in domain adaptation is the source-only training.
+Let $h=g\circ f$ be the composition of a feature extractor $g:\mathcal{X}\rightarrow\mathcal{U}$ and a feature classifier $f:\mathcal{U}\rightarrow\mathcal{Y}$.
+We can train $h$ end-to-end by minimizing,
+$\mathcal{L}(\theta)=\dfrac{1}{K}\sum_{k=1}^{K}\dfrac{1}{n_{k}}\sum_{i=1}^{n_{k}}\mathcal{L}(y_{i}^{(Q_{k})},f(g(x_{i}^{(Q_{k})}))).$
+This training produces an embedding space $\mathcal{U}$ that is discriminative of the classes in the domain adaptation problem.
+Following previous work [montesuma2021wasserstein, montesuma2023multi, montesuma2024distillation], we perform adaptation in this space.
+Wasserstein Barycenters and Domain Adaptation.
+The motivation for using Wasserstein barycenters in multi-source domain adaptation comes from a theoretical result of [redko2019optimal].
+Indeed, by [redko2019optimal, Theorem 2],
+$\mathcal{R}_{Q_{T}}(h)\leq\mathcal{R}_{Q_{S}}(h)+\mathbb{W}_{1}(\hat{Q}_{S},\hat{Q}_{T})+\mathcal{O}(n_{S}^{-1/2}+n_{T}^{-1/2})+\mathbb{R}(Q_{S},Q_{T}),$ (41)
+where $\mathbb{R}(Q_{S},Q_{T})=\text{inf}_{h\in\mathcal{H}}\mathcal{R}_{Q_{S}}(h)+\mathcal{R}_{Q_{T}}(h)$ is the joint risk.
+This result establishes that when source and target measures are close, their risks are close as well.
+Now, assume equation 41 holds for each pair $(Q_{k},Q_{T})$, then, summing over $k$ weighted by $\lambda_{k}$,
+$\mathcal{R}_{Q_{T}}(h)\leq\sum_{k=1}^{K}\lambda_{k}\mathcal{R}_{Q_{k}}(h)+\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{1}(\hat{Q}_{k},\hat{Q}_{T})+\mathcal{O}(n^{-1/2})+\sum_{k=1}^{K}\lambda_{k}\mathbb{R}(Q_{k},Q_{T}),$
+using the triangle inequality on $\mathbb{W}_{1}(\hat{Q}_{k},\hat{Q}_{T})$
+$\mathcal{R}_{Q_{T}}(h)\leq\sum_{k=1}^{K}\lambda_{k}\mathcal{R}_{Q_{k}}(h)+\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{1}(\hat{Q}_{k},P)+\mathbb{W}_{1}(P,\hat{Q}_{T})+\mathcal{O}(n^{-1/2})+\sum_{k=1}^{K}\lambda_{k}\mathbb{R}(Q_{k},Q_{T}),$
+A reasonable choice to make this bound tight is choosing $P$ that minimizes $P\mapsto\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{1}(\hat{Q}_{k},P)+\mathbb{W}_{1}(P,\hat{Q}_{T})$.
+From $p=1$ to $p>1$.
+The argument for $p>1$ is a bit more intricate.
+Since $\mathbb{W}_{1}$ is the weakest of the Wasserstein distances [villani2008optimal, Remark 6.6], we can directly move from $p=1$ to $p>1$,
+$\mathcal{R}_{Q_{T}}(h)\leq\sum_{k=1}^{K}\lambda_{k}\mathcal{R}_{Q_{k}}(h)+\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{p}(\hat{Q}_{k},\hat{Q}_{T})+\mathcal{O}(n^{-1/2})+\sum_{k=1}^{K}\lambda_{k}\mathbb{R}(Q_{k},Q_{T}),$
+which, from the triangle inequality, results in,
+$\mathcal{R}_{Q_{T}}(h)\leq\sum_{k=1}^{K}\lambda_{k}\mathcal{R}_{Q_{k}}(h)+\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{p}(\hat{Q}_{k},P)+\mathbb{W}_{p}(P,\hat{Q}_{T})+\mathcal{O}(n^{-1/2})+\sum_{k=1}^{K}\lambda_{k}\mathbb{R}(Q_{k},Q_{T}),$ (42)
+The term $\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{p}(\hat{Q}_{k},P)+\mathbb{W}_{p}(P,\hat{Q}_{T})$ does not correspond to the $p$-Wasserstein barycenter due the missing power of $p$.
+We can retrieve it through Young’s inequality for products.
+Let $a,b>0$ be real numbers, and $p,q>1$.
+Then $ab\leq\frac{a^{p}}{p}+\frac{b^{q}}{q}.$ We can get the inequality,
+$\mathbb{W}_{p}(\hat{P},\hat{Q})\leq\dfrac{\mathbb{W}_{p}(\hat{P},\hat{Q})^{p}}{p}+\dfrac{p-1}{p},$
+[…]
+$\mathbb{W}_{p}(\hat{Q}_{k},P)\leq\frac{1}{p}\mathbb{W}_{p}(\hat{Q}_{k},P)^{p}+\frac{p-1}{p}\text{, and }\mathbb{W}_{p}(P,\hat{Q}_{T})\leq\frac{1}{p}\mathbb{W}_{p}(P,\hat{Q}_{T})^{p}+\frac{p-1}{p}$
+Plugging this back into inequality 42,
+$\mathcal{R}_{Q_{T}}(h)\leq\sum_{k=1}^{K}\lambda_{k}\mathcal{R}_{Q_{k}}(h)+\dfrac{1}{p}\biggr(\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{p}(\hat{Q}_{k},P)^{p}+\mathbb{W}_{p}(P,\hat{Q}_{T})^{p}\biggr)+\dfrac{2(p-1)}{p}+\mathcal{O}(n^{-1/2})+\sum_{k=1}^{K}\lambda_{k}\mathbb{R}(Q_{k},Q_{T}).$
+Modulo some additional constants depending on $p>1$, we have the same reasoning as $p=1$.
+Algorithm 3 Wasserstein Barycenter Transport
+[…]
+2:Labeled target domain data $\{T_{P\rightarrow Q_{T}}(x_{i}^{(P)}),y_{i}^{(P)}\}_{i=1}^{(P)}$
+3:for $\tau\leftarrow 1$ to $n_{\text{iter}}$ do
+4:  Compute $P=\text{Bar}(\lambda,\mathcal{Q})$
+5:  Compute $\gamma^{\star}$ between barycenter and target using eq. 43
+6:  Transport $(x_{i}^{(P)},y_{i}^{(P)})$ Using eq. 44
+7:end for
+Wasserstein Barycenter Transport $(p=2)$. [montesuma2021wasserstein] uses a two-step procedure for finding $P$. First, one solves,
+$P^{\star}=\underset{P\in\mathcal{P}_{1,\text{ac}}(\Omega)}{\text{arg min}}\,\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2}(Q_{k},P)^{2},$
+which is the 22-Wasserstein barycenter problem. Then, one aligns the barycenter with the target domain through OT. This is done through the barycenter mapping,
+$\gamma^{\star}$ $=\underset{\gamma\in\Gamma(P,\hat{Q}_{T})}{\text{arg min}}\,\sum_{i=1}^{n}\sum_{j=1}^{n_{T}}\gamma_{ij}\lVert x_{i}^{(P)}-x_{j}^{(Q_{T})}\rVert_{2}^{2},$ (43)
+$T_{P\rightarrow Q_{T}}(x_{i}^{(P)})$ $=n_{T}\sum_{j=1}^{n_{T}}\gamma_{ij}^{\star}x_{j}^{(Q_{T})},$ (44)
+which is the minimizer of $P\mapsto\mathbb{W}_{2}(P,\hat{Q}_{T})^{2}$.
+We provide an algorithm for this strategy in Algorithm 3.
+Note that since this procedure generates a labeled dataset on data following the target domain measure, we can train a classifier with the pairs $\{T_{P\rightarrow Q_{T}}(x_{i}^{(P)}),y_{i}^{(P)}\}$.
+Appendix D Missing Proofs and Derivations
+D.1 Connecting equations 12 and 16
+In this section, we the empirical flow in equation 16 from the continuity equation. As a reminder, the continuity equation is,
+$\partial_{t}P_{t}=\text{div}(P_{t}\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t})).$
+Here we want to plug $\hat{P}_{t}=n^{-1}\sum_{i=1}^{n}\delta_{z_{t,i}}$ into this equation. On the one hand, plugging the formula for the empirical measure into the r.h.s. of this equation gives us,
+$\text{div}(\hat{P}_{t}\nabla\mkern-10.0mu\nabla\mathbb{F}(\hat{P}_{t}))=\text{div}\biggr(\dfrac{1}{n}\sum_{i=1}^{n}\delta_{z_{t,i}}\nabla(\delta\mathbb{F}(\hat{P}_{t}))(z_{t,i})\biggr).$ (45)
+Here, we used the definition for the Wasserstein gradient (c.f., equation 36), i.e., $\nabla\mkern-10.0mu\nabla\mathbb{F}(P)=\nabla(\delta\mathbb{F}(P))(z_{t,i})$.
+On the other hand, due to the linearity of derivatives and the chain rule,
+$\partial_{t}\hat{P}_{t}=\dfrac{1}{n}\sum_{i=1}^{n}\partial_{t}\delta_{z_{t,i}^{(P)}}=-\dfrac{1}{n}\sum_{i=1}^{n}\text{div}(\partial_{t}z_{t,i}\delta_{z_{t,i}})=\text{div}\biggr(-\dfrac{1}{n}\sum_{i=1}^{n}\delta_{z_{t,i}}\partial_{t}z_{t,i}\biggr).$ (46)
+Equating the r.h.s.
+of equations 45 and 46 gives us the continuous-time update $\partial_{t}z_{t,i}=-\nabla\mkern-10.0mu\nabla\mathbb{F}(\hat{P}_{t})$.
+We can discretize this differential equation in time, using the Forward Euler scheme,
+$z_{\tau+1,i}=z_{\tau,i}-\alpha\nabla\mkern-10.0mu\nabla\mathbb{F}(\hat{P}_{\tau}),$
+where $\alpha>0$ is the time-step and $\tau$ is the discrete time index.
+D.2 Proposition 3.1
+As a recap, in this section we provide a formal proof that when,
+$P=\sum_{i}\pi_{i}^{(P)}(\mathcal{N}(\cdot|\mu_{i}^{(P)},\Sigma_{i}^{(P)})\otimes\delta_{\nu_{i}^{(P)}}),$
+and under the ground cost,
+$c(z,z^{\prime})=\lVert x-x^{\prime}\rVert_{2}^{2}+\rho(y,y^{\prime}),$
+the $\mathbb{MW}_{2}$ distance becomes,
+$\mathbb{MW}_{2}(P,Q)^{2}=\sum_{i=1}^{n}\sum_{j=1}^{m}\omega^{\star}_{ij}(\mathbb{W}_{2}(P_{i},Q_{j})^{2}+\rho(\nu_{i}^{(P)},\nu_{j}^{(Q)})^{2}).$
+As a reminder, we have the following proposition,
+Proposition D.1.
+Let $P=\sum\pi_{i}^{(P)}(\mathcal{N}(\mu_{i}^{(P)},\Sigma_{i}^{(P)})\otimes\delta_{\nu_{i}^{(P)}})$ and $Q=\sum\pi_{j}^{(Q)}(\mathcal{N}(\mu_{j}^{(Q)},\Sigma_{j}^{(Q)}\otimes\delta_{\nu_{j}^{(Q)}})$ be 2 GMMs over $\Omega=\mathcal{X}\times\mathcal{Y}$.
+Let the ground cost $c$ be,
+$c(z,z^{\prime})=\lVert x-x^{\prime}\rVert_{2}^{2}+\rho(y,y^{\prime})^{2},$
+where $\rho(y,y^{\prime})$ is a metric over $\mathcal{Y}$. Then,
+$\mathbb{MW}_{2}(P,Q)^{2}=\underset{\omega\in\Gamma(\pi^{(P)},\pi^{(Q)})}{\text{min}}\,\sum_{i=1}^{n}\sum_{j=1}^{m}\omega_{ij}C_{ij},$
+where $C_{ij}=\mathbb{W}_{2}(P_{i,x},Q_{j,x})^{2}+\rho(\nu_{i}^{(P)},\nu_{j}^{(P)})^{2}$.
+Proof.
+Our first step is decomposing the Wasserstein distance $\mathbb{W}_{2}(P_{i},Q_{j})^{2}$ into $\mathbb{W}_{2}(P_{i,x},Q_{j,x})^{2}+\rho(\nu_{i}^{(P)},\nu_{j}^{(Q)})^{2}$. Indeed,
+$\mathbb{W}_{2}(P_{i},Q_{j})^{2}=\int\int c(z,z^{\prime})d\gamma(z,z^{\prime}),$
+where $z=(x,y)$ is the feature-label pair.
+Note that since $P_{i}=P_{i,x}\otimes\delta_{\nu_{i}^{(P)}}$, $\gamma$ is supported on the set $\{(x,y,x,y^{\prime}):y=\nu_{i}^{(P)},y^{\prime}=\nu_{j}^{(Q)}\}$.
+Thus, the OT plan has the structure:
+$\gamma=\gamma_{x}\otimes\delta_{(\nu_{i}^{(P)},\nu_{j}^{(Q)})}$
+where $\gamma_{x}$ is the OT plan between Gaussians $P_{i,x}$ and $Q_{j,x}$. This means that the previous integral can be decomposed into,
+$\mathbb{W}_{2}(P_{i},Q_{j})^{2}=\int\int(\lVert x-x^{\prime}\rVert_{2}^{2}+\rho(y,y^{\prime})^{2})d\gamma(z,z^{\prime}),$
+which we can decompose two terms. The first,
+$\int\int\lVert x-x^{\prime}\rVert_{2}^{2}d\gamma(z,z^{\prime})$ $=\int\int\lVert x-x^{\prime}\rVert_{2}^{2}d\gamma_{x}(x,x^{\prime})=\mathbb{W}_{2}(P_{i,x},Q_{j,x})^{2},$
+and,
+$\int\int\rho(y,y^{\prime})^{2}d\gamma(z,z^{\prime})$ $=\int\int\rho(y,y^{\prime})^{2}\delta_{(\nu_{i}^{(P)},\nu_{j}^{(Q)})}dydy^{\prime}=\rho(\nu_{i}^{(P)},\nu_{j}^{(Q)})^{2},$
+which gives the desired decomposition. ∎
+The result in the previous proposition stems from the special structure we defined for the GMMs.
+Since the result is agnostic with respect the label metric, it holds for $\rho(\nu_{i}^{(P)},\nu_{j}^{(Q)})=\lVert\nu_{i}^{(P)}-\nu_{j}^{(Q)}\rVert_{2}$
+D.3 Theorem 3.1
+In this section, we present the proof for theorem 3.1. We mainly use two tools to obtain this result,
+  1. 1.
+    The exponential convergence given by the PL inequality,
+    $\lVert\nabla\mkern-10.0mu\nabla\mathbb{B}(P)\rVert_{L_{2}(P)}\geq C_{\text{PL}}(\mathbb{B}(P)-\mathbb{B}^{\star}),$ (47)
+  2. 2.
+    An uniform bound on $P\mapsto\mathbb{B}_{\mathcal{Q}}(P)$ controlling the error of approximating $Q_{k}$ with $\hat{Q}_{k}$.
+These results are available, respectively, in [chewi2024statistical] and [portales2025sample].
+For completeness, we re-state and demonstrate them.
+We start with a modified version of [chewi2024statistical, Corollary 5.17],
+Proposition D.2.
+(Measure theoretic Gronwall’s Lemma [chewi2024statistical]) Let $\mathbb{F}:\mathcal{P}_{2,\text{ac}}(\Omega)\rightarrow\mathbb{R}$ be a functional that satisfies the PL inequality (c.f.
+47) with constant $C_{\text{PL}}>0$, then,
+$\mathbb{F}(P_{\tau})-\mathbb{F}^{\star}\leq e^{-2C_{\text{PL}}\tau}(\mathbb{F}(P_{0})-\mathbb{F}^{\star}).$ (48)
+Proof.
+Define $\phi(t)=\mathbb{F}(P_{t})-\mathbb{F}^{\star}$. Then,
+$\partial_{t}\phi(t)$ $=\partial_{t}\mathbb{F}(P_{t})=-\lVert\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t})\rVert_{L_{2}(P_{t})}^{2},$
+$\geq-C_{\text{PL}}(\mathbb{F}(P_{0})-\mathbb{F}^{\star})=-C_{\text{PL}}\phi(t).$
+From Grönwall’s Lemma [chewi2024statistical, Lemma 5.16],
+$\underbrace{\mathbb{F}(P_{t})-\mathbb{F}^{\star}}_{\partial_{t}\phi(t)}\leq e^{-C_{\text{PL}}}\underbrace{(\mathbb{F}(P_{0})-\mathbb{F}(P)^{\star})}_{\phi(0)}$
+We now present the uniform approximation bound of [portales2025sample] for empirical barycenters,
+Theorem D.1.
+(Barycenter Empirical Approximation error [portales2025sample]) Let $\mathcal{Q}=\{Q_{1},\cdots,Q_{K}\}$ be a finite family of measures on $\mathcal{P}_{p,\text{ac}}(\mathcal{B}(0,R))$ for $R>0$.
+Let $\{z_{i}^{(Q_{k})}\}_{i=1}^{n_{k}}$, $i=1,\cdots,n_{k}$ be a set of $K$ i.i.d.
+samples from each $Q_{k}$.
+Let $\hat{P}=n^{-1}\sum_{i=1}^{n}\delta_{z_{i}^{(P)}}$.
+Let $p\in[1,+\infty)$.
+Then,
+$\mathbb{E}\biggr[\underset{z_{1}^{(P)},\cdots,z_{n}^{(P)}\in\Omega}{\text{sup}}\,|\mathbb{B}_{\mathcal{Q}}(\hat{P})-\widehat{\mathbb{B}}_{\mathcal{Q}}(\hat{P})|\biggr]\leq C_{p,R}\sqrt{\dfrac{C_{d,m}}{n}},$ (49)
+where,
+$C_{p,R}$ $=8\sqrt{2}\int_{0}^{3(2R)^{p}}\sqrt{\log\biggr(2+\dfrac{32p(2R)^{p}}{s}\biggr)}ds,$
+$C_{d,m}$ $=m(d+1).$
+With these bounds established, we can then prove our Theorem 3.1,
+Proof.
+Let $\tau>0$ be the current optimization step. We have the following,
+[…]
+[…]
+From equation 50, we can bound the expected empirical optimization gap,
+$\mathbb{E}[|\hat{\mathbb{B}}_{\mathcal{Q}}(\hat{P}_{\tau})-\hat{\mathbb{B}}^{\star}_{\mathcal{Q}}|]\leq$ $\mathbb{E}[|\hat{\mathbb{B}}_{\mathcal{Q}}(\hat{P}_{\tau})-\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})|]+\mathbb{E}[|\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})-\mathbb{B}_{\mathcal{Q}}^{\star}|]+\mathbb{E}[|\mathbb{B}_{\mathcal{Q}}^{\star}-\hat{\mathbb{B}}^{\star}_{\mathcal{Q}}|].$ (51)
+Now, we can bound the first and third terms using the approximation error of Portales et al. [portales2025sample] (i.e., equation 49),
+$\mathbb{E}[|\hat{\mathbb{B}}_{\mathcal{Q}}(\hat{P}_{\tau})-\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})|]$ $\leq C_{p,R}\sqrt{\dfrac{C_{d,m}}{n}},$ (52)
+$\mathbb{E}[|\hat{\mathbb{B}}_{\mathcal{Q}}^{\star}-\mathbb{B}_{\mathcal{Q}}^{\star}|]$ $\leq C_{p,R}\sqrt{\dfrac{C_{d,m}}{n}},$
+here, we recall that $\mathbb{B}_{\mathcal{Q}}^{\star}=\mathbb{B}_{\mathcal{Q}}(P^{\star})=\text{inf}_{P}\mathbb{B}_{\mathcal{Q}}(P)$.
+Coming back to the second term, note that the true optimization gap $\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})-\mathbb{B}^{\star}_{\mathcal{Q}}$ is deterministic, as it does not involve sampling from the measures in $\mathcal{Q}$.
+This means that,
+$\mathbb{E}[|\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})-\mathbb{B}_{\mathcal{Q}}^{\star}|]=|\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})-\mathbb{B}_{\mathcal{Q}}^{\star}|\leq e^{-C_{\text{PL}}\tau}(\hat{\mathbb{B}}_{\mathcal{Q}}(\hat{P}_{0})-\hat{\mathbb{B}}_{\mathcal{Q}}^{\star}).$ (53)
+Combining equations 5352 and 51, we get,
+$\mathbb{E}[|\hat{\mathbb{B}}_{\mathcal{Q}}(\hat{P}_{\tau})-\hat{\mathbb{B}}^{\star}_{\mathcal{Q}}|]\leq e^{-C_{\text{PL}}\tau}(\mathbb{B}_{\mathcal{Q}}(\hat{P}_{0})-\mathbb{B}^{\star}_{\mathcal{Q}})+C_{p,R}\sqrt{\dfrac{C_{d,m}}{n}},$
+which is the desired inequality. ∎
+D.4 Theorem 3.2
+The goal of this section is to stablish the proof of theorem 3.2.
+The proof roughly follows the same idea of the empirical case.
+However, since all the involved measures are GMMs by hypothesis, we do not have empirical approximation gaps per se.
+In this regard, we do consider the stochasticity of fitting GMMs, provided data from each underlying $Q_{k}$.
+As a result, we need an approximation tool, in the spirit of theorem D.1 of [portales2025sample].
+We do so through the following result of [boite2025differentiable],
+Proposition D.3.
+(GMM approximation bound [boite2025differentiable]) Let $P$ and $Q$ be GMMs with parameters $\{(\pi^{(P)}_{i},\mu_{i}^{(P)},\Sigma_{i}^{(P)})\}_{i=1}^{n}$ (resp.
+$Q$), and let $\hat{P}$ and $\hat{Q}$ be the corresponding GMMs empirically estimated by the expectation-maximization algorithm.
+Assume the following,
+  • Hypothesis 1 (bounded parameters): There are constants $R_{\mu}>0$ and $R_{\Sigma}>0$ such that,
+    $\lVert\mu_{i}^{(P)}\rVert_{2}\leq R_{\mu},\sqrt{\text{Tr}(\Sigma_{i}^{(P)})}\leq R_{\Sigma},\lVert\hat{\mu}_{i}^{(P)}\rVert_{2}\leq R_{\mu},\sqrt{\text{Tr}(\hat{\Sigma}_{i}^{(P)})}\leq R_{\Sigma}\text{, resp.
+}Q,$
+  • Hypothesis 2 (empirical convergence): Assume that exist $\rho_{\pi}>0$, $\rho_{\mu}$, and $\rho_{\Sigma}>0$ such that,
+    $\mathbb{E}[\lVert\pi^{(P)}-\hat{\pi}^{(P)}\rVert_{1}]\leq\rho_{\pi}\quad\mathbb{E}[\lVert\mu_{i}^{(P)}-\hat{\mu}^{(P)}_{i}\rVert_{2}]$ $\leq\rho_{\mu}\quad\mathbb{E}[d_{\text{bures}}(\Sigma_{i}^{(P)},\hat{\Sigma}_{i}^{(P)})]$ $\leq\rho_{\Sigma}$
+    $i=1,\cdots,n$ (resp. $Q$), where,
+    $d_{\text{bures}}(\Sigma,\Sigma^{\prime})=\sqrt{\text{Tr}(\Sigma+\Sigma^{\prime}-2(\Sigma^{1/2}\Sigma^{\prime}\Sigma^{1/2})^{1/2})}.$
+In these conditions,
+$\mathbb{E}[|\mathbb{MW}_{2}(\hat{P},\hat{Q})^{2}-\mathbb{MW}_{2}(P,Q)^{2}|]\leq\underbrace{8R_{\mu}\rho_{\mu}+8R_{\Sigma}\rho_{\Sigma}+8(R_{\mu}^{2}+R_{\Sigma}^{2})\rho_{\pi}}_{C_{\pi,\mu,\Sigma}},$ (54)
+Note that the same result applies for labeled GMM.
+Indeed, since we fit a GMM per class, the labels are deterministic.
+As a result, $\nu_{i}^{(P)}=\hat{\nu}_{i}^{(P)}$.
+These previous theorems allow us to bound the barycenter functional,
+Corollary D.1.
+Let $\{Q_{k}\}_{k=1}^{K}$ be $K>0$ GMMs over $\Omega$, and let $\{\hat{Q}_{k}\}_{k=1}^{K}$ be their corresponding empirical approximations.
+Assume Hypothesis 1 and 2 of Proposition 4.3.
+Let $P$ be a GMM, and $\lambda\in\Delta_{K}$, then,
+$\mathbb{E}[|\mathbb{B}(P)-\hat{\mathbb{B}}(P)|]\leq\sum_{k=1}^{K}\lambda_{k}C_{k}=C_{\lambda,\mathcal{Q}},$
+where $C_{k}=8R_{\mu_{k}}+8R_{\Sigma_{k}}\rho_{\Sigma_{k}}+8(R_{\mu_{k}}^{2}+R_{\Sigma_{k}}^{2})\rho_{\pi_{k}}$.
+Proof.
+The proof relies on the triangle inequality:
+$|\mathbb{B}(P)-\hat{\mathbb{B}}(P)|$ $\leq\biggr|\sum_{k=1}^{K}\lambda_{k}\mathbb{MW}_{2}(P,Q_{k})^{2}-\sum_{k}\lambda_{k}\mathbb{MW}_{2}(P,\hat{Q}_{k})^{2}\biggr|$
+$\leq\sum_{k=1}^{K}\lambda_{k}|\mathbb{MW}_{2}(P,Q_{k})^{2}-\mathbb{MW}_{2}(P,\hat{Q}_{k})^{2}|,$
+then, taking the expectation on both sides,
+$\mathbb{E}[|\mathbb{B}(P)-\hat{\mathbb{B}}(P)|]\leq\sum_{k=1}^{K}\lambda_{k}\mathbb{E}[|\mathbb{MW}_{2}(P,Q_{k})^{2}-\mathbb{MW}_{2}(P,\hat{Q}_{k})^{2}|]\leq\sum_{k=1}^{K}\lambda_{k}C_{k}=C_{\lambda,\mathcal{Q}}.$
+Furthermore, note that the constant can be written explicitly as,
+$C_{\lambda,\mathcal{Q}}=\sum_{k=1}^{K}\lambda_{k}(8R_{\mu_{k}}+8R_{\Sigma_{k}}\rho_{\Sigma_{k}}+8(R_{\mu_{k}}^{2}+R_{\Sigma_{k}}^{2})\rho_{\pi_{k}})$ (55)
+With these results established, we now prove Theorem 3.2,
+Proof.
+Our arguments largely follow those in section D.3. For instance, the steps taken in equations 50 and 51 directly apply to the GMM setting. Now, in place of equation 52, we have,
+$\mathbb{E}[|\hat{\mathbb{B}}_{\mathcal{Q}}(\hat{P}_{\tau})-\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})|]\leq C_{\lambda,\mathcal{Q}},\quad\mathbb{E}[|\hat{\mathbb{B}}_{\mathcal{Q}}^{\star}-\mathbb{B}_{\mathcal{Q}}^{\star}|]\leq C_{\lambda,\mathcal{Q}}.$ (56)
+Furthermore, equation 53 also holds in the GMM case. Indeed, the optimization gap is deterministic. Therefore, combining equations 56 with 51 and 53, we get,
+$\mathbb{E}[|\hat{\mathbb{B}}_{\mathcal{Q}}(\hat{P}_{\tau})-\hat{\mathbb{B}}^{\star}_{\mathcal{Q}}|]\leq e^{-C_{\text{PL}}\tau}(\mathbb{B}_{\mathcal{Q}}(\hat{P}_{0})-\mathbb{B}^{\star}_{\mathcal{Q}})+C_{\lambda,\mathcal{Q}},$
+which is the desired inequality.
+Appendix E Additional Toy Experiments
+E.1 Two Moons
+Figure 5: Two moons samples alongside the fitted GMMs. Stars and ellipses denote the means and covariance matrices.
+In this experiment, similarly to the Swiss-roll example we explored in the main paper, we deal with Location-Scatter families of measures.
+The idea is that we are given $Q_{0}\in\mathcal{P}_{2,\text{ac}}$, and we define a finite family of measures,
+$\mathcal{Q}=\{T_{k,\sharp}Q_{0}:T_{k}(x)=A_{k}x+b_{k},A_{k}\in\text{PD}(d)\}_{k=1}^{K}$ (57)
+where $\text{PD}(d)$ denotes the set of $d\times d$ positive-definite matrices. The barycenter of $Q_{1},\cdots,Q_{K}$ is itself a push-forward of $Q_{0}$, noted $T$. Furthermore, $T$ has the form,
+$T(x)=\sum_{k=1}^{K}\lambda_{k}T_{k}(x)=\biggr(\sum_{k=1}^{K}\lambda_{k}A_{k}\biggr)x+\sum_{k=1}^{K}\lambda_{k}b_{k}$
+For this experiment, we use $Q_{0}$ as the two moons dataset [^2] More specifically, we use https://scikit-learn.org/stable/modules/generated/sklearn.datasets.make_moons.html to generate data from the two moons measure, and https://scikit-learn.org/stable/modules/generated/sklearn.datasets.make_swiss_roll.html to generate data from the Swiss roll measure..
+For the transformations, we use a family of rotation-shift affine functions, i.e.,
+$A_{k}=\begin{bmatrix}\cos\theta_{k}&-\sin\theta_{k}\\ \sin\theta_{k}&\cos\theta_{k}\end{bmatrix},\quad b_{k}\in\mathbb{R}^{2}.$
+where $\theta_{k}\in\{-30,-15,15,30\}$ degrees, and,
+$b_{k}\in\{\begin{bmatrix}0,0\end{bmatrix},\begin{bmatrix}10,0\end{bmatrix},\begin{bmatrix}0,10\end{bmatrix},\begin{bmatrix}10,10\end{bmatrix}\}$.
+The generated family of measures is shown in Figure 5.
+Figure 6: From left to right, comparison of the empirical barycenter with and without an interaction energy term, and the Gaussian mixture barycenter.
+Overall, due its clustering effect, the GMM barycenter better respects the underlying geometry of the input measures.
+We show a comparison of the empirical and GMM barycenters in Figure 6.
+Note that, without an interaction term, the empirical barycenter mixes a few samples from opposing classes.
+This artifact comes from the rotations in the input measures.
+Adding the interaction energy term $U_{R}$ solves this problem, at the cost of some deformation of the class clusters.
+Finally, the GMM flow is able to compute an accurate barycenter due its clustering effect on the Gaussian components.
+We refer readers to [montesuma2024lighter] and [montesuma2025optimal] for more details on this effect.
+E.2 Handling Class Imbalance
+A challenge classic Wasserstein barycenter algorithms face is handling class imbalance.
+For instance, since the algorithm in [montesuma2023multi] relies on label propagation, the soft-labels in the Wasserstein barycenter might end up mixing classes.
+This phenomenon leads to Wasserstein barycenters with mixed classes and fuzzy labels.
+We tackle these problems with the functionals described in Section 3.3 in the main paper.
+We show, in Figure 7, an example of the imbalance problem.
+Here, $\pi=[\pi_{0},\pi_{1}]$ denotes the class proportions.
+Figure 7: Imbalanced Gaussian blobs for the class imbalance example. We draw $\pi_{0}\sim\mathcal{U}[0,1]$ and let $\pi_{1}=1-\pi_{0}$.
+We start by comparing the joint Wasserstein barycenter algorithm of [montesuma2023multi] (a) with our mini-batch strategy (b).
+Note that using mini-batches already helps with the fuzzy label problem, but many labels are still uncertain in the final barycenter.
+Adding the label entropy penalty $\mathbb{V}$, with underlying functional $V$ defined in equation 20 in the main paper, leads to sharp labels, as we show in (c).
+Finally, by adding the interaction energy term we can separate the classes, as show in Figure 8 (d).
+(a) [montesuma2023multi].
+(b) $\mathbb{B}_{\mathcal{Q}}$ (ours).
+(c) $\mathbb{B}_{\mathcal{Q}}+\mathbb{V}$ (ours).
+(d) $\mathbb{B}_{\mathcal{Q}}+\mathbb{V}+\mathbb{U}$ (ours).
+Figure 8: Discrete (a), and mini-batch Wasserstein barycenters (b and c).
+Overall, the discrete algorithm of [montesuma2023multi] struggles with class imbalances, since the transport plan matches points from different classes.
+This phenomenon leads to fuzzy samples near the class boundary.
+This issue is partly solved through mini-batching (b), but a definitive solution comes with adding the entropy functional defined in equation 20 in the main paper.
+Appendix F Additional details on domain adaptation
+F.1 Evaluation Protocol, Benchmarks, Fine-Grained Results
+Benchmarks.
+In total, we compare five domain adaptation benchmarks, divided into computer vision, neuroscience and chemical engineering.
+An overview of the benchmarks is presented in Table 1 in our main paper.
+Here, we give further details on each of the benchmarks.
+For computer vision, we use the Office 31 [saenko2010adapting] and Office Home [venkateswara2017deep] benchmarks.
+For neuroscience, we use the BCI-CIV-2a [brunner2008bci] and ISRUC [khalighi2013automatic] benchmarks.
+For chemical engineering, we use the TEP benchmark [montesuma2024benchmarking].
+Evaluation Protocol. As we describe in the following, all benchmarks deal with multi-class classification. We measure performance using the standard classification accuracy,
+$\text{accuracy}(y,\hat{y})=\dfrac{1}{n}\sum_{i=1}^{n}\delta(y_{i}-\hat{y}_{i})\times 100\%.$
+For the Office 31, Office Home, Tennessee Eastman Process, and BCI-CIV-2a, we adopt the so-called leave-one-domain-out evaluation.
+For instance, given $3$ domains $\{A,B,C\}$, we perform $3$ experiments, leaving one domain as the target domain at each time.
+For instance, one of these experiments would be $\{A,B\}\rightarrow C$, where $\{A,B\}$ are the sources, and $C$ is the target.
+Backbones.
+As we discuss in the main paper, we perform adaptation over embeddings.
+To that end, we use the same experimental setting as [montesuma2024lighter] for the Office 31, Office Home, and TEP benchmarks.
+More specifically, we train a ResNet 50 and ResNet 101 for the Office 31 and Office Home benchmarks, and a CNN for the TEP benchmark.
+For this latter backbone, we refer readers to [montesuma2024benchmarking] for further details.
+For neuroscience benchmarks, we follow the experimental protocol of [wang2024cbramod], where we fine-tune CBraMod, a transformer-based foundation model, on available source domain data.
+We use the exact same training settings as these authors.
+Algorithm $\mathcal{X}\times\mathcal{Y}$ A D W Avg. $\uparrow$
+ResNet50 - 67.50 95.00 96.83 86.40
+WJDOT [turrisi2022multi] - 67.77 97.32 95.32 86.80
+DaDiL-E [montesuma2023multi] - 70.55 100.00 98.83 89.79
+DaDiL-R [montesuma2023multi] - 70.90 100.00 98.83 89.91
+GMM-DaDiL [montesuma2024lighter] - 72.47 100.00 99.41 90.63
+Discrete [cuturi2014fast] 64.45 88.39 92.98 81.94
+NormFlow [visentin2025computing] 65.67 97.32 94.73 85.91
+CW2B [korotin2021continuous] 69.16 94.64 95.32 86.37
+NOT [kolesov2024estimating] 69.33 92.85 96.49 86.22
+U-NOT [gazdieva2024robust] 68.29 97.32 95.32 86.97
+Discrete [montesuma2023multi] 67.94 98.21 97.66 87.93
+GMM [montesuma2024lighter] 70.13 99.11 96.49 88.54
+WGF (ours) 71.25 98.21 97.07 88.84
+WGF-GMM (ours) 71.77 99.11 97.07 89.34
+Table 4: Office 31.
+Office 31 [saenko2010adapting] contains, in total, 4652 images among 31 classes of objects.
+These images are divided into 3 different domains: Amazon, dSLR, and Webcam.
+Images in the Amazon domain represent products, usually with an homogeneous background, medium resolution, and studio lightning.
+There are 2817 images in this domain.
+The dSLR domain contains high-resolution images obtained through a digital SLR camera in an office environment.
+There are 498 images in this domain.
+Finally, the Webcam domain contains images of objects captured with a webcam, also in an office environment.
+There 795 images in this domain.
+As analyzed in [ringwald2021adaptiope], this benchmark is one of the most commonly used in the literature.
+Despite its moderate size, it presents challenges of its own.
+For instance, since the Amazon domain images were collected automatically, it has label noise.
+We show our results in the Office 31 benchmark in Table 4.
+Overall, dictionary learning methods such as DaDiL [montesuma2023multi] or GMM-DaDiL [montesuma2024lighter] remain state-of-the-art for this benchmark.
+However, our gradient flow algorithm is able to improve previous methods, closing the gap between these two kinds of algorithms.
+In comparison with unsupervised barycenters, using labels in the ground-cost ($\mathcal{X}\times\mathcal{Y}$ = ✓) has a clear advantage.
+Algorithm $\mathcal{X}\times\mathcal{Y}$ Ar Cl Pr Rw Avg. $\uparrow$
+ResNet101 - 72.90 62.20 83.70 85.00 75.95
+WJDOT [turrisi2022multi] - 74.28 63.80 83.78 84.52 76.59
+DaDiL-E [montesuma2023multi] - 77.16 64.95 85.47 84.97 78.14
+DaDiL-R [montesuma2023multi] - 75.92 64.83 85.36 85.32 77.86
+GMM-DaDiL [montesuma2024lighter] - 77.16 66.21 86.15 85.32 78.81
+Discrete [cuturi2014fast] 68.11 59.79 75.56 78.21 70.42
+NormFlow [visentin2025computing] 77.16 62.54 83.10 84.63 76.86
+CW2B [korotin2021continuous] 74.69 63.46 82.54 85.09 76.44
+NOT [kolesov2024estimating] 72.22 63.57 82.77 82.91 75.36
+U-NOT [gazdieva2024robust] 75.30 64.94 82.88 84.28 76.85
+Discrete [montesuma2023multi] 75.72 63.80 84.23 84.63 77.09
+GMM [montesuma2024lighter] 75.31 64.26 86.71 85.21 77.87
+WGF (ours) 74.48 65.17 85.92 85.77 77.84
+WGF-GMM (ours) 74.89 65.97 87.16 86.81 78.71
+Table 5: Office-Home.
+Office-Home [venkateswara2017deep] contains, 15,500 images among 65 classes of objects.
+There are four domains: Art, Clipart, Product, and Real-World.
+The Art domain contains artistic depictions of objects, such as sketches, paintings, and ornamentation.
+It has 2427 images.
+The Clipart domain is a collection of graphic art representing the objects.
+This domain has 4365 images.
+Product contains images of objects without a background.
+It has 4439 images.
+Real-World contains images of objects captured with a camera.
+It contains 4357 images.
+In this sense, the Product domain is similar to the Amazon domain in the Office 31 benchmark, and the Real-World is similar to dSLR or Webcam.
+In comparison with the Amazon benchmark, the Office-Home benchmark adds another factor of domain variation: style (e.g., Art vs.
+Real-World).
+Furthermore, due the larger number of classes, one might expect a harder classification problem.
+We show our results in Table 5.
+In comparison with the Office 31 benchmark, our methods get even closer to the state-of-the-art.
+For instance, WGF-GMM has only a $0.10\%$ gap in performance with GMM-DaDiL, while showing state-of-the-art performance on domains Product and Real-World.
+Again, we verify that using labels in the ground cost ($\mathcal{X}\times\mathcal{Y}$ = ✓) offers a clear advantage.
+Remark F.1.
+We follow previous works on visual domain adaptation, especially [saenko2010adapting] and [venkateswara2017deep], and report the classification accuracy on fixed partitions of the target domains.
+For that reason, Tables 4 and 5 do not have confidence intervals.
+While it would be possible to run these experiments with random partitions (e.g., $k$-fold cross validation), this would mean that our results would not be comparable to previous established results in [montesuma2024lighter], for instance.
+Tennessee-Eastman Process (TEP) [montesuma2024benchmarking] is a benchmark in chemical engineering.
+This benchmark comprises a set of simulations of a chemical reaction that occurs inside a reactor.
+In total, 34 variables are measured over time.
+As a result, we have a set of time-series representing how the reaction progresses over each simulation.
+In some of these simulations, one of 28 possible faults can be introduced.
+The goal is, based on the readings, to determine which kind of fault, or its absence, has occurred.
+As a result, there are 29 classes.
+Each simulation is also associated with a production mode, which drastically changes the statistical properties of the data generation process, thus introducing a distribution shift.
+As a result, each domain corresponds to one of these modes of operation.
+We refer readers to [montesuma2024benchmarking] and [reinartz2021extended] for more details.
+Algorithm $\mathcal{X}\times\mathcal{Y}$ Mode 1 Mode 2 Mode 3 Mode 4 Mode 5 Mode 6 Avg. $\uparrow$
+CNN 80.82 $\pm$ 0.96 63.69 $\pm$ 1.71 87.47 $\pm$ 0.99 79.96 $\pm$ 1.07 74.44 $\pm$ 1.52 84.53 $\pm$ 1.12 78.48
+WJDOT [turrisi2022multi] - 89.06 $\pm$ 1.34 75.60 $\pm$ 1.84 89.99 $\pm$ 0.86 89.38 $\pm$ 0.77 85.32 $\pm$ 1.29 87.43 $\pm$ 1.23 86.13
+DaDiL-E [montesuma2023multi] - 90.45 $\pm$ 1.02 77.08 $\pm$ 1.21 86.79 $\pm$ 2.14 89.01 $\pm$ 1.35 84.04 $\pm$ 3.16 87.85 $\pm$ 1.06 85.87
+DaDiL-R [montesuma2023multi] - 91.97 $\pm$ 1.22 77.15 $\pm$ 1.32 85.41 $\pm$ 1.69 89.39 $\pm$ 1.03 84.49 $\pm$ 1.95 88.44 $\pm$ 1.29 86.14
+GMM-DaDiL [montesuma2024lighter] - 91.72 $\pm$ 1.41 76.41 $\pm$ 1.89 89.68 $\pm$ 1.49 89.18 $\pm$ 1.17 86.05 $\pm$ 1.46 88.02 $\pm$ 1.12 86.85
+Discrete [cuturi2014fast] 90.38 $\pm$ 1.29 71.17 $\pm$ 1.26 85.27 $\pm$ 1.09 87.29 $\pm$ 0.99 83.83 $\pm$ 1.74 84.94 $\pm$ 1.87 83.81
+NormFlow [visentin2025computing] 87.06 $\pm$ 0.57 66.25 $\pm$ 1.11 88.58 $\pm$ 1.00 89.35 $\pm$ 0.34 80.99 $\pm$ 2.21 85.08 $\pm$ 1.45 82.89
+CW2B [korotin2021continuous] 91.72 $\pm$ 0.79 73.71 $\pm$ 1.42 88.37 $\pm$ 1.31 89.32 $\pm$ 1.14 84.66 $\pm$ 1.75 87.19 $\pm$ 1.24 85.83
+NOT [kolesov2024estimating] 90.31 $\pm$ 1.08 72.37 $\pm$ 0.94 87.99 $\pm$ 1.31 89.28 $\pm$ 0.81 84.35 $\pm$ 1.68 85.29 $\pm$ 2.02 84.93
+U-NOT [gazdieva2024robust] 90.82 $\pm$ 0.65 72.68 $\pm$ 1.58 88.65 $\pm$ 0.95 89.00 $\pm$ 1.28 85.22 $\pm$ 1.07 86.19 $\pm$ 1.70 85.43
+Discrete [montesuma2023multi] 92.38 $\pm$ 0.66 73.74 $\pm$ 1.07 88.89 $\pm$ 0.85 89.38 $\pm$ 1.26 85.53 $\pm$ 1.35 86.60 $\pm$ 1.63 86.09
+GMM [montesuma2024lighter] 92.23 $\pm$ 0.70 71.81 $\pm$ 1.78 84.72 $\pm$ 1.92 89.28 $\pm$ 1.55 87.51 $\pm$ 1.73 82.49 $\pm$ 1.81 84.67
+WGF (ours) 92.41 $\pm$ 0.90 74.16 $\pm$ 1.29 88.78 $\pm$ 0.91 89.84 $\pm$ 0.62 84.66 $\pm$ 2.37 87.39 $\pm$ 2.11 86.21
+WGF-GMM (ours) 92.55 $\pm$ 1.12 75.95 $\pm$ 1.51 89.54 $\pm$ 1.11 89.21 $\pm$ 1.10 85.63 $\pm$ 1.17 87.74 $\pm$ 1.48 86.77
+Table 6: TEP.
+We report our results in Table 6.
+For this benchmark, the gap between unlabeled ($\mathcal{X}\times\mathcal{Y}$ = ✗) and labeled ($\mathcal{X}\times\mathcal{Y}$ = ✓) barycenter methods is closer.
+For instance, CW2B [korotin2021continuous] achieves 85.83 for the average target domain performance versus 86.09 for the discrete barycenter.
+Furthermore, it surpasses the GMM method of [montesuma2024lighter].
+In comparison, our methods establish a new state-of-the-art for barycenter-based domain adaptation in this benchmark.
+Note that our WGF-GMM is able to get the second best performance overall, close to GMM-DaDiL.
+Algorithm $\mathcal{X}\times\mathcal{Y}$ 1 2 3 4 5 6 7 8 9 Avg. $\uparrow$
+CBraMod - 52.22 40.45 64.23 40.41 47.74 48.34 53.64 50.52 55.03 50.30
+DaDiL-R [montesuma2023multi] - 58.68 40.27 66.14 41.45 59.54 51.38 53.12 57.46 55.21 53.69
+GMM-DaDiL [montesuma2024lighter] - 60.59 42.36 70.31 48.12 61.11 50.34 62.15 58.85 60.00 57.10
+Discrete [cuturi2014fast] 60.93 43.40 69.96 50.00 61.28 50.69 61.63 61.45 57.11 57.38
+NormFlow [visentin2025computing] 57.11 43.22 67.53 47.71 60.24 51.04 56.25 60.76 55.20 55.45
+CW2B [korotin2021continuous] 60.00 43.57 70.13 48.95 60.93 51.21 61.28 61.63 57.46 57.25
+NOT [kolesov2024estimating] 61.28 43.22 70.31 47.91 60.59 51.21 61.46 61.80 57.29 57.23
+U-NOT [gazdieva2024robust] 60.59 44.09 69.44 47.70 61.11 50.34 61.97 61.11 58.15 57.17
+Discrete [montesuma2023multi] 60.76 42.53 70.66 48.54 61.28 50.69 61.63 61.28 59.54 57.43
+GMM [montesuma2024lighter] 60.59 42.53 69.79 47.08 61.45 50.34 62.32 59.54 59.72 57.04
+WGF (ours) 61.63 43.40 71.00 47.91 61.45 51.38 61.97 61.63 57.12 57.50
+WGF-GMM (ours) 60.59 43.05 70.48 47.29 61.80 51.56 62.15 61.80 57.46 57.35
+Table 7: BCI-CIV-2a.
+BCI-CIV-2a [brunner2008bci] Contains electroencephalogram (EEG) data from nine subjects.
+In brief, an EEG measures the electrical activity of the brain over time.
+As such, like the TEP benchmark, this benchmark has a collection of time series data.
+The BCI-CIV-2a data is concerned with motor imagery tasks, that is, the imagination of movement of the left hand, right hand, both feet, and tongue.
+These correspond to class 1, 2, 3, and 4.
+Therefore, in this benchmark we have four classes.
+Our goal here is to perform cross-subject adaptation, namely, we use data from a given set of source subjects, and try to predict on a target subject.
+Therefore, each domain is a subject, so there are nine domains in total.
+There are, on average, $560$ samples per subject, with a total of $5088$ samples.
+We show our results in Table 7.
+For cross-subject studies, we highlight that, besides improving individual domain performance, it is desirable to not degrade the source-only performance.
+This is indeed the case for almost all methods.
+For instance, DaDiL [montesuma2023multi] degrades performance of subjects 2 and 7 by a small margin.
+For this benchmark, the difference between unlabeled ($\mathcal{X}\times\mathcal{Y}$ = ✗) and labeled ($\mathcal{X}\times\mathcal{Y}$ = ✓) is not as marked as before.
+We hypothesize that this benchmark falls into the covariate shift hypothesis, namely, $Q_{S}(X)\neq Q_{T}(X)$, rather than the more general joint shift $Q_{S}(X,Y)\neq Q_{T}(X,Y)$.
+Algorithm 1 2 3 4 5 6 7 8 9 10 Avg. $\uparrow$
+CBraMod 78.97 80.35 70.46 85.59 80.69 72.85 77.00 73.29 83.41 68.95 76.63
+WJDOT [turrisi2022multi] 80.41 78.25 65.69 82.85 81.74 77.74 76.33 68.75 86.95 70.81 76.95
+DaDiL-E [montesuma2023multi] 76.12 77.90 65.93 83.21 81.28 68.45 75.66 72.61 79.75 65.93 75.14
+DaDiL-R [montesuma2023multi] 78.06 78.13 66.97 84.64 81.51 72.02 76.33 72.84 80.85 67.55 75.89
+GMM-DaDiL [montesuma2024lighter] 76.63 78.95 66.51 82.73 77.44 73.81 77.00 71.70 83.05 66.86 75.47
+Discrete [cuturi2014fast] 78.77 80.58 69.76 85.83 79.06 73.45 78.33 72.50 84.14 68.48 77.09
+NormFlow [visentin2025computing] 81.12 82.90 75.58 84.28 80.34 78.69 75.88 70.34 80.48 70.81 78.05
+CW2B [korotin2021continuous] 79.08 78.25 69.07 86.54 78.48 70.35 76.88 70.90 82.80 66.04 75.84
+NOT [kolesov2024estimating] 78.57 79.88 67.32 85.59 78.72 73.21 77.77 72.50 83.41 67.21 76.44
+U-NOT [gazdieva2024robust] 80.20 83.48 73.72 87.50 76.51 72.62 76.66 70.22 78.17 71.51 77.06
+Discrete [montesuma2023multi] 78.87 80.69 72.09 86.07 78.60 78.45 78.00 72.95 85.48 70.81 78.20
+GMM [montesuma2024lighter] 75.20 75.46 69.88 80.71 78.60 71.90 76.33 68.41 78.17 68.14 74.28
+WGF (ours) 80.20 81.62 76.86 86.42 79.30 80.00 75.88 76.13 86.95 74.42 79.78
+WGF-GMM (ours) 79.89 81.74 72.32 86.42 80.69 79.04 77.55 72.72 85.73 71.04 78.72
+Table 8: ISRUC.
+ISRUC [khalighi2013automatic] is a benchmark for EEG data for sleep staging.
+In brief, sleep staging is the classification of physiological data, such as EEG, recorded during sleep into different sleep stages.
+In this context, the ISRUC benchmark is the most large scale benchmark used in this paper.
+This study comprises 100 subjects, with a total of 89240 samples (approximately 900 samples per domain).
+The sleep staging problem in this benchmark has five classes: awake, non-rapid eye movement sleep (subdivided into N1, N2, N3), and rapid eye movement sleep.
+Due the elevated number of subjects in this benchmark, we adopt a slightly different evaluation protocol.
+Instead of using the leave-of-domain-out strategy, we fix the first $90$ subjects as training subjects and evaluate the performance of adapting to the remaining $10$ subjects.
+This means that we compute Wasserstein barycenters of $K=90$ input measures.
+In this benchmark, using our gradient flow strategy yields the best result overall (e.g., $79.78$ average domain performance for our empirical gradient flow).
+Furthermore, we see an important gap between the GMM barycenter algorithm of [montesuma2024lighter] and ours, which is similar to the gap found in the TEP benchmark (c.f.
+Table 6).
+There are mainly two reasons for this.
+First, we can freely choose the learning rate for updating the GMM components, which leads to an overall more stable algorithm.
+Second, we can regularize the barycenter problem, with additional functionals.
+F.2 Adapting Neural Network Barycenter Methods to Domain Adaptation
+In this section, we cover how we adapted previous neural network-based barycenters to perform domain adaptation.
+There is one main challenge with using unlabeled barycenters for domain adaptation: we cannot directly move the barycenter to the target domain, as we would end up with more unlabeled data.
+A straightforward solution to this problem is moving the available labeled data that we have, that is, the source domain data.
+For this scenario, the barycenter domain works as a pivot domain, as we transport the source to the target through the composition,
+$T_{Q_{k}\rightarrow Q_{T}}(x_{i}^{(Q_{k})})=T_{P\rightarrow Q_{T}}(T_{Q_{k}\rightarrow P}(x_{i}^{(Q_{k})})).$
+In comparison, by computing a labeled barycenter, we are synthesizing data in $P^{\star}$.
+As we demonstrated in Tables 4 through 8, this is generally beneficial for domain adaptation.
+In this way, we can acquire a labeled dataset in the target domain through $\{\{T_{P\rightarrow Q_{T}}(T_{Q_{k}\rightarrow P}(x_{i}^{(Q_{k})})),y_{i}^{(Q_{k})}\}_{i=1}^{n_{k}}\}_{k=1}^{K}$.
+To isolate the effect of the quality of the computed barycenters, we estimate the mappings as we did in equations 43 and $4$, that is,
+$T_{Q_{k}\rightarrow P}(x_{i}^{(Q_{k})})=n\sum_{j=1}^{n}\gamma_{k,i,j}^{\star}x_{j}^{(P)}\quad T_{P\rightarrow Q_{k}}(x_{i}^{(P)})=n_{T}\sum_{j=1}^{n_{T}}\gamma_{ij}^{\star}x_{j}^{(Q_{T})},$
+where $\gamma_{k}^{\star}$ is the OT plan from $Q_{k}$ to $P$, and $\gamma^{\star}$ is the OT plan from $P$ to $Q_{T}$.
+F.3 Complexity Analysis and Running Time
+Complexity Analysis. We start with analyzing the complexity of evaluating the functionals used in our work. We recall that,
+$\mathbb{V}(P)=\dfrac{1}{n}\sum_{i=1}^{n}V(z_{i}^{(P)}),\quad\mathbb{U}(P)=\dfrac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}U(z_{i}^{(P)},z_{j}^{(P)}),$
+where, assuming that the complexity of evaluating $V$ and $U$ is $\mathcal{O}(1)$, the complexity of evaluating $\mathbb{V}$ is $\mathcal{O}(n)$ and for $\mathbb{U}$ is $\mathcal{O}(n^{2})$.
+For $\mathbb{B}_{\mathcal{Q}}$, the complexity depends on the number of elements used in OT, and the number of input measures.
+In other words, we have $\mathcal{O}(Kn_{b}^{3}\log n_{b})$ for the empirical case, and $\mathcal{O}(Kn_{\text{components}}\log n_{\text{components}})$ for WGF-GMM.
+We assume this cost dominates over the evaluation of other functionals.
+Method Complexity per iteration Parameter Complexity Running time per iteration (s) # Iterations Running Time (s)
+Discrete [cuturi2014fast, montesuma2023multi] $Kn^{3}\log n$ $n$ 4.28 50 214.37
+NormFlow [visentin2025computing] N/A $|f|$ 0.17 5000 878.32
+CW2B [korotin2021continuous] N/A $K(|f|+|g|)$ 2.18 1000 2181.00
+NOT [kolesov2024estimating] N/A $K(|f|+|g|)$ 41.63 25 1040.75
+U-NOT [gazdieva2024robust] N/A $K(|f|+|g|)$ 13.29 100 1329.77
+WGF (ours) $Kn_{b}^{3}\log n_{b}$ $n$ 0.58 200 117.24
+WGF-GMM (ours) $Kn_{\text{components}}^{3}\log n_{\text{components}}$ $n_{\text{components}}$ 0.11 150 17.14
+Table 9: Complexity and running time (in seconds) of different barycenter calculation strategies on the ISRUC benchmark.
+$n_{b}$ denotes the batch size, $n$ number of samples, $K$ number of input measures, $d$ number of dimensions, $n_{\text{components}}$ number of components in GMMs.
+For NormFlow, $f$ denotes the neural net implementing the normalizing flow.
+For CW2b, NOT and U-NOT, $f$ and $g$ denote the neural networks approximating the Kantorovich potentials.
+$|f|$ and $|g|$ denote the number of their parameters.
+Running Time.
+We perform a running time analysis of the tested methods on the ISRUC benchmark, which contains the most number of samples and domains.
+For all neural net methods, we fix a batch size of $256$ samples.
+Overall, our proposed methods run significantly faster than previous methods.
+For instance, in comparison with the discrete solver of [cuturi2014fast], our algorithm is approximately seven times faster per iteration.
+It is noteworthy that some neural net methods are relatively faster per iteration (e.g.
+NormFlow).
+However, we empirically verified that these methods sometimes need a large number of iterations to converge to a good barycenter (e.g.
+5000 iterations).
+F.4 Hyperparameter Setting
+In this section, we detail the hyperparameters of our methods. For both flows, we have mainly three parameters,
+  • Number of samples in the barycenter support, $n$. This hyperparameter is replaced by the number of components in the GMM barycenter, $K$.
+  • Learning rate $\alpha$ for the gradient flow,
+  • Number of iterations $n_{\text{iter}}$,
+additionally, for the empirical flow, we have the batch size $n_{b}$ for the input measures.
+We show the complete list of hyperparameters in Table 10.
+Additionally to these parameters, we also have the choice of regularizing functionals, $\mathbb{V}$ and $\mathbb{U}$.
+Remark F.2.
+In all of our domain adaptation experiments, we use a diagonal GMM. We empirically verified that optimizing with respect to complete covariances is numerically unstable and time consuming.
+WGF WGF-GMM
+Benchmark $n$ $n_{b}$ $\alpha$ $n_{\text{iter}}$ $\mathbb{V}$ $\mathbb{U}$ $K$ $\alpha$ $n_{\text{iter}}$ $\mathbb{V}$ $\mathbb{U}$
+Office31 1550 1550 1.0 200 $\mathbb{W}_{2}(P,Q_{T})$ Repulsion 124 1.0 150 - -
+Office-Home 1300 650 1.0 200 $\mathbb{W}_{2}(P,Q_{T})$ Repulsion 975 1.0 200 $\mathbb{W}_{2}(P,Q_{T})$ Repulsion
+TEP 870 290 1.0 150 $\mathbb{W}_{2}(P,Q_{T})$ - 145 0.1 200 - -
+BCI-CIV-2a 800 200 1.0 - - Repulsion 400 0.01 200 Entropy Repulsion
+ISRUC 1000 1000 0.1 200 - - 40 0.1 150 Entropy Repulsion
+Table 10: Hyperparameters used in each benchmark. $n$ denotes number of samples, $\alpha$ denotes learning rate, $n_{\text{iter}}$ denotes number of iterations.
+We offer some additional reasoning on how to choose these parameters.
+The number of samples and components controls the complexity of the calculated barycenter.
+We express these values in terms of the number of classes.
+For instance, $n=1550$ in the Office 31 benchmark translates to $n=31\times 50$ samples, that is, $50$ samples per class.
+The same reasoning applies for the batch size.
+Furthermore, our method is fairly robust to the choice of learning rate.
+For instance, as we discussed in section C.1, the fixed-point iterations in the discrete barycenter are equivalent to setting $\alpha=n/2$, where $n$ is the number of samples in the Wasserstein barycenter.
+This generally strikes a balance on how fast the algorithm converges.

### v3

@@
-Computing Wasserstein Barycenters through Gradient Flows
+Wasserstein Gradient Flows for Scalable and Regularized Barycenter Computation
 Abstract
-Wasserstein barycenters provide a powerful tool for aggregating probability measures, while leveraging the geometry of their ambient space.
-Existing discrete methods suffer from poor scalability, as they require access to the complete set of samples from input measures.
-We address this issue by recasting the original barycenter problem as a gradient flow in the Wasserstein space.
-Our approach offers two advantages.
-First, we achieve scalability by sampling mini-batches from the input measures.
-Second, we incorporate functionals over probability measures, which regularize the barycenter problem through internal, potential, and interaction energies.
-We present two algorithms for empirical and Gaussian mixture measures, providing convergence guarantees under the Polyak-Łojasiewicz inequality.
-Experimental validation on toy datasets and domain adaptation benchmarks show that our methods outperform previous discrete and neural net-based methods for computing Wasserstein barycenters.
+Wasserstein barycenters provide a principled approach for aggregating probability measures, while preserving the geometry of their ambient space.
+Existing discrete methods are not scalable as they assume access to the complete set of samples from the input measures.
+Meanwhile, neural network approaches do scale well, but rely on complex optimization problems and cannot easily incorporate label information.
+We address these limitations through gradient flows in the space of probability measures.
+Through time discretization, we achieve a scalable algorithm that i) relies on mini-batch optimal transport, ii) accepts modular regularization through task-aware functions, and iii) seamlessly integrates supervised information into the ground-cost.
+We empirically validate our approach on domain adaptation benchmarks that span computer vision, neuroscience, and chemical engineering.
+Our method establishes a new state-of-the-art barycenter solver, with labeled barycenters consistently outperforming unlabeled ones.
 1 Introduction
-(a) Static barycenter.
-(b) Barycenter flow.
-Figure 1: In (a), we show the usual static notion of the Wasserstein barycenter (in red), which minimizes the sum of distances to the input measures (in blue).
-In (b), we show our notion of barycenter as a gradient flow, flowing an initial measure $P_{0}$ (purple) to the barycenter $P^{\star}$ (yellow) of the input measures.
-Wasserstein barycenters [agueh2011barycenters] constitute a building block in the analysis of probability measures under Optimal Transport (OT) [villani2008optimal].
-This mathematical object extends the notion of the barycenter of points to Wasserstein spaces, thus defining a notion of the average of probability measures.
-Since it relies on the Wasserstein metric, these barycenters capture the geometry of the underlying space over which the probability measures are defined [kloeckner2010geometric].
-As such, they contributed to several areas of machine learning, including model averaging [singh2020model], ensembling [coz2023barycenter], data augmentation [zhu2023interpolation], distillation [liu2023dataset, montesuma2024multisource], dictionary learning [schmitz2018wasserstein], domain adaptation [montesuma2021wasserstein, montesuma2023multi, montesuma2024lighter], and Bayesian learning [srivastava2018scalable].
-Existing methods for computing Wasserstein barycenters can be divided into three categories.
-The first category relies on discretization of the input measures and the barycentric measure through empirical measures.
-These algorithms may be subdivided into fixed-support [cuturi2014fast, Algorithm 1][benamou2015iterative, janati2020debiased], and free-support [cuturi2014fast, Algorithm 2][montesuma2023multi].
-The second category relies on parametric models for the input measures, such as Gaussian, or Gaussian mixture models.
-In the Gaussian case, these barycenters are known as Bures-Wasserstein barycenters [chewi2020gradient, kroshnin2021statistical].
-For Gaussian mixtures, [delon2020wasserstein] shows connections with multi-marginal OT [pass2015multi] between the components of the input mixtures, while [montesuma2024lighter] shows that, for components with diagonal covariances, there is an iterative algorithm similar to [cuturi2014fast, Algorithm 2].
-The third category of methods for computing Wasserstein barycenters relies on neural nets and comes from the effort to scale OT to measures in high-dimensional spaces with a large number of samples.
-In this sense, [korotin2021continuous] proposes parameterizing the barycentric measure through a specific kind of neural net architecture, known as Input Convex Neural Networks (ICNNs) [amos2017input].
-Furthermore, [kolesov2024estimating] proposes a bi-level adversarial learning approach that can be used for general ground costs, which is made more robust by semi-unbalanced OT [gazdieva2024robust].
-More recently, [visentin2025computing] proposes computing barycenters with normalizing flows.
-Current Wasserstein barycenter algorithms mostly focus on scalability with respect to the number of samples $n$, and dimensions $d$.
-In this sense, neural networks have a clear advantage over discrete methods, since they can operate at the level of mini-batches.
-There is another level of scalability, oftentimes overlooked, which is with respect the number of input measures $K$.
-So far, only [visentin2025computing] has considered experiments in this context.
-In this regard, methods such as [korotin2021continuous, korotin2022wasserstein] are difficult to scale, as they require $\mathcal{O}(K)$ neural networks to compute the Wasserstein barycenter.
-In this paper we propose a new perspective on discrete Wasserstein barycenters, through the lens of Wasserstein Gradient Flows (WGFs) [ambrosio2008gradient, santambrogio2017euclidean], which is an appealing framework for optimization in the space of probability measures.
-This new understanding of Wasserstein barycenters has several advantages.
-First, we can scale the discrete barycenter problem, by drawing mini-batches from the input measures.
-Second, we can scale with respect to the number of input measures, as the representation of the barycentric measure is independent of $K$.
-Third, based on the established literature on gradient flows, we can define regularized Wasserstein barycenters with suitable functionals on the space of probability measures, which were not possible with previous methods.
-Lastly, we can modify the underlying metric to integrate label information between samples, leading to barycenters that respect class structure and provide more accurate and semantically meaningful estimates.
-We show an overview of our approach in Figure 1.
-Our contributions are:
-  • We introduce a unified framework for computing Wasserstein barycenters based on gradient flows, which provides a principled and flexible approach to optimizing in the space of probability measures.
-  • We generalize algorithms from the state-of-the-art [cuturi2014fast, montesuma2023multi, montesuma2024lighter] under a common framework.
-  • We offer theoretical guarantees for the convergence of our algorithms, which also apply to previous works.
-  • We offer a new theoretical result for the mixture-Wasserstein distance between labeled Gaussian mixtures [delon2020wasserstein, montesuma2024lighter].
-  • We conduct extensive experiments benchmarking existing methods, including neural network-based Wasserstein barycenter algorithms [korotin2021continuous, kolesov2024estimating, gazdieva2024robust, visentin2025computing], on both toy datasets and domain adaptation  [^1] In contrast with discrete methods, most experiments in neural network solvers involve generative modeling.
-We thus establish a bridge between these methods and domain adaptation benchmarks, and show that our approach establishes a new state-of-the-art.
+Defining the mean or center of a set of probability measures is a fundamental primitive in geometric probability theory [nielsen2020].
+Given $K\in\mathbb{N}$ probability measures $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}$ on a metric space $(\Omega,d)$, their Wasserstein barycenter [agueh2011barycenters] is given by,
+$P^{\star}=\underset{P\in\mathcal{P}_{2}(\Omega)}{\text{arg min}}\,\biggl\{\mathbb{B}(P)=\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2}(P,Q_{k})^{2}\biggr\},$ (1)
+where $\mathcal{P}_{2}(\Omega)$ is the set of measures over $\Omega$ with finite second moments, and $\lambda=(\lambda_{1},\cdots\lambda_{k})$, $\lambda_{k}\geq 0$, $\forall k$, $\sum_{k}\lambda_{k}=1$ is an array of barycentric coordinates.
+Equation 1 defines a notion of average of the input measures in $\mathcal{Q}$.
+The main advantage of these barycenters stems from Optimal Transport (OT) [villani2008optimal], that is, they lift the geometry of the space $(\Omega,d)$ to the space $(\mathcal{P}_{2}(\Omega),\mathbb{W}_{2})$ [kloeckner2010geometric].
+Wasserstein barycenters have wide applications to machine learning, such as posterior aggregation in Bayesian inference [srivastava2018scalable], model fusion [singh2020model], ensembling in time series forecasting [coz2023barycenter], data augmentation [zhu2023interpolation], and Domain Adaptation (DA) [montesuma2021wasserstein, montesuma2023multi, montesuma2024lighter].
+In each application, the main advantage is geometric fidelity, that is, the Wasserstein barycenter captures and preserves the underlying structure of samples.
+Figure 1 shows a conceptual illustration.
+(a) Conceptual Illustration.
+(b) $\mathbb{B}$.
+(c) $\mathbb{B}+\mathbb{V}$.
+(d) $\mathbb{B}+\mathbb{V}+\mathbb{U}$.
+Figure 1: In (a), we show an illustration of our flow.
+We show the initial barycenter measure $P_{0}$ which flow towards the solution $P^{\star}$.
+In red, green, and orange, we show the input measures’ mini-batches as solid dots.
+The flow is given by the combination of gradients $v_{\tau,i}=\sum_{k=1}^{K}\lambda_{k}\nabla\mathbb{W}_{2}(P_{\tau},Q_{k})$.
+In parallel, (b) shows a pathological barycenter calculation, with fuzzy labels and unclear class boundaries.
+Adding regularizers $\mathbb{V}$ (b) and $\mathbb{U}$ (c) progressively resolves this issue.
+Despite its advantages, Wasserstein barycenters remain challenging beyond the Gaussian setting [altschuler2022wasserstein].
+We have identified three gaps in the current literature.
+First, the seminal algorithm of cuturi2014fast requires having access to the complete set of input measures’ samples at once, making it intractable for large datasets.
+Neural network methods [fan2020scalable, korotin2021continuous, kolesov2024estimating, gazdieva2024robust] alleviate this issue by operating on mini-batches, but parametrize the barycentric measure with $\mathcal{O}(K)$ neural networks, coupling model complexity with the number of input measures.
+Second, neural methods cannot seamlessly incorporate label information into their ground-cost, limiting their performance in supervised tasks (cf.
+Table 2).
+Third, the objective function in equation 1 only accounts for distributional fit.
+In practice, the barycentric measure must satisfy additional structural properties, such as class separation.
+Current barycenter solvers do not offer a principled way of enforcing these properties.
+Recent methods study the barycenter problem through the gradient flow perspective.
+First, chewi2020gradient established a gradient flow algorithm on the Bures-Wasserstein manifold, i.e., on Gaussian measures.
+Second, chizat2025doubly describes the doubly regularized barycenter problem through the same lens.
+Here, in addition to entropic regularization at the level of OT, they proposed an outer regularization that penalizes the entropy of the barycentric measure, which leads to a Noisy Particle Gradient Descent (NPGD) algorithm for empirical measures.
+Although this algorithm incorporates outer regularization, it has the same limitation as cuturi2014fast, as it assumes full-batch access to the samples of input measures.
+We address the aforementioned challenges through the lens of gradient flows in the space of probability measures [ambrosio2008gradient, santambrogio2017euclidean].
+We therefore conceptualize the barycenter problem as the flow from an initial measure $P_{0}=\mathcal{N}(0,\text{Id})$ following the Wasserstein gradient of the functional,
+$\mathbb{F}(P)=\mathbb{B}(P)+\mathbb{R}(P),$ (2)
+where $\mathbb{B}$ denotes the barycenter functional (cf. equation 1), and $\mathbb{R}$ denotes additional regularizing functionals (see Sections 2.3 and 3.3, and Appendix C.2).
+Following the established gradient flow literature [santambrogio2017euclidean], we decompose $\mathbb{R}$ into internal, potential, and interaction energies, denoted $\mathbb{G}$, $\mathbb{V}$, and $\mathbb{U}$, respectively.
+This decomposition generalizes the double regularization strategy of chizat2025doubly.
+Indeed, setting $\mathbb{V}=\mathbb{U}=0$ and $\mathbb{G}(P)=\eta_{G}H(P)$, where $H(P)$ denotes the entropy of $P$, corresponds to their exact framework.
+Similarly, alvarez2021dataset devised an approach for flowing a single dataset towards a known target.
+Our setting differs structurally from that.
+First, we go beyond the single measure objective, optimizing the multi-measure barycentric objective (cf.
+equation 1).
+Second, while alvarez2021dataset transforms an existing dataset towards a known target, ours synthesizes a measure (i.e., the barycentric measure) from noise.
+Our contributions are as follows.
+(i) A mini-batch, time discretized gradient flow algorithm for regularized empirical Wasserstein barycenters (Algorithm 1).
+By randomly sampling the input measures, we reduce the computational complexity of traditional barycenter methods [cuturi2014fast, montesuma2023multi, chizat2025doubly] achieving $2\times$ to $50\times$ speedups with respect to discrete solvers (Figure 5).
+(ii) Modular, task-aware regularizing functionals for Wasserstein barycenters (Section 3.3).
+Following previous gradient flow literature [santambrogio2017euclidean, alvarez2021dataset], we decompose $\mathbb{R}$ into internal, potential, and interaction energies, allowing incorporation of plug-and-play regularizers into the barycenter problem.
+This goes beyond chizat2025doubly, who considered a particular internal energy as an outer regularization strategy.
+(iii) Extensive DA benchmarking (Section 4).
+We empirically show that our algorithm produces barycenters that respect class structure, by incorporating both appropriate functionals (Section 3.3), and labels into the ground-cost of optimal transport (Section 3.2).
+On 5 DA benchmarks, spanning computer vision, neuroscience, and chemical engineering, we establish a comparison between empirical and neural barycenter solvers, showing that structured costs are essential for DA performance.
 Paper organization.
-Section 2 presents the notation and preliminaries of our paper.
-Section 3 covers our proposed algorithms, including the theoretical results.
-Section 4 describes our experiments.
+Section 2 reviews OT, Wasserstein barycenters and gradient flows.
+Section 3 develops our empirical, time-discretized gradient flow framework and algorithms.
+Section 4 presents our experiments.
 Finally, Section 5 concludes this paper.
 2 Background
-Throughout this paper, $(\Omega,d)$ is a metric space.
-We denote by $P$ and $Q$, probability measures on $\mathcal{P}(\Omega)$, which is the space of measures over $\Omega$.
-For $p\in[1,+\infty)$ we further denote by $\mathcal{P}_{p,ac}(\Omega)$, the sub-set of absolutely continuous measures w.r.t.
-the Lebesgue measure such that $\int_{\Omega}d(z,z_{0})^{p}dP(z)<+\infty$.
-In our case, we have access to these measures through their samples, denoted $z^{(P)}\sim P$.
-The set $\Delta_{K}$ denotes the $K$-simplex, i.e., $y\in\Delta_{n}=\{y\in\mathbb{R}^{n}:\sum_{i=1}^{n}y_{i}=1\text{ and }y_{i}\geq 0\text{ }\forall i\}$.
-We denote functionals over $\mathcal{P}_{2,\text{ac}}$ with blackboard bold letters (e.g., $\mathbb{F}$), and their (Wasserstein) gradients by $\nabla\mkern-10.0mu\nabla$.
-We offer a short introduction in Section 3 in the Appendix.
-2.1 Empirical Optimal Transport
-In this section, given $n$ i.i.d. samples from $P$, we can approximate a probability $P$ empirically through,
-$\hat{P}(z)=\dfrac{1}{n}\sum_{i=1}^{n}\delta(z-z_{i}^{(P)}),$ (1)
-which gives a finite parametrization for $P$. Henceforth we use a hat to indicate empirical measures (e.g., $\hat{P}$).
-OT [villani2008optimal] is a field of mathematics concerned with the transportation of mass at least effort.
-We refer readers to [peyre2019computational] for a computational treatment of the subject, and [montesuma2024recent] for applications in machine learning.
-Let $P,Q\in\mathcal{P}(\Omega)$, and $c:\Omega^{2}\rightarrow\mathbb{R}$ be a ground-cost.
-OT was originally founded by Monge [monge1781memoire].
-In this formulation, we seek for a mapping $T:\Omega\rightarrow\Omega$ such that,
-$T_{P\rightarrow Q}^{\star}=\underset{T_{\sharp}P=Q}{\text{arg inf}}\,\int_{\Omega}c(z,T(z))dP(z).$ (2)
-Alternatively, Kantorovich [kantorovich1942transfer] proposed a formulation in terms of a transport plan $\gamma\in\Gamma(P,Q)$, where $\Gamma(P,Q)$ is the set of all joint measures with marginals $P$ and $Q$.
+Throughout this paper $(\Omega,d)$ is a metric space, and its elements are denoted $z\in\Omega$.
+We denote $\mathcal{P}_{2}(\Omega)$ as the set of measures with finite second moments [villani2008optimal, Definition 6.4].
+We work with empirical measures,
+$\text{Emp}_{n}(\Omega)=\biggr\{\hat{P}\in\mathcal{P}_{2}(\Omega):\hat{P}=\frac{1}{n}\sum_{i=1}^{n}\delta_{z_{i}^{(P)}}\biggr\},$ (3)
+where $\delta_{z_{0}}(z)=\delta(z-z_{0})$ denotes the Dirac measure centered at $z_{0}\in\Omega$.
+Given a measure $P\in\mathcal{P}_{2}(\Omega)$, we denote by $\hat{P}\in\text{Emp}_{n}(\Omega)$ its empirical approximation based on i.i.d.
+samples $z_{1}^{(P)},\cdots,z_{n}^{(P)}\sim P$.
+The superscript $\cdot^{(P)}$ indicates the measure from which the samples originate.
+2.1 Optimal Transport
+OT is a field of mathematics [villani2008optimal] concerned with mass transportation at least effort [peyre2019computational, montesuma2024recent].
+It was originally founded by monge1781memoire, and seeks a mapping such that,
+$T_{P\rightarrow Q}^{\star}=\underset{T_{\sharp}P=Q}{\text{arg inf}}\,\mathbb{E}_{z\sim P}[d(z,T(z))^{2}].$ (4)
+$T_{P\rightarrow Q}^{\star}$ is an OT mapping between $P$ and $Q$, and $T_{\sharp}$ is the pushforward mapping $P$ to $Q$.
+Alternatively, kantorovich1942transfer proposed a formulation in terms of a transport plan $\gamma\in\Gamma(P,Q)$, where $\Gamma(P,Q)$ is the set of all joint measures with marginals $P$ and $Q$.
 In this case,
-$\gamma^{\star}=\underset{\gamma\in\Gamma(P,Q)}{\text{arg inf}}\,\int_{\Omega}\int_{\Omega}c(z,z^{\prime})\gamma(z,z)$ (3)
-Given $p\in[1,+\infty)$, when $c(z,z^{\prime})=d(z,z^{\prime})^{p}$, the previous problems define the $p$-Wasserstein distance,
-$\mathbb{W}_{p}(P,Q)^{p}=\underset{\gamma\in\Gamma(P,Q)}{\text{inf}}\,\int_{\Omega}\int_{\Omega}d(z,z^{\prime})^{p}d\gamma(z,z^{\prime}).$ (4)
-Equation 4 defines a metric in $\mathcal{P}_{p,ac}(\Omega)$.
-Conceptually, the Wasserstein distance lifts the metric $d$ on $\Omega$ to $\mathbb{W}_{p}$ on $\mathcal{P}_{p,ac}(\Omega)$.
-Based on this idea, we can define barycenters and gradient flows in the space of probability measures.
-2.2 Gaussian Mixture Optimal Transport
-In comparison with equation 1, we can approximate $P$ through a Gaussian Mixture Model (GMM),
-$P(z)=\sum_{i=1}^{n}\pi_{i}^{(P)}\mathcal{N}(z|\mu_{i}^{(P)},\Sigma_{i}^{(P)}),$ (5)
-where $\theta=\{(\pi_{i}^{(P)},\mu_{i}^{(P)},\Sigma_{i}^{(P)})\}$ are the parameters of the GMM.
-The general theory of OT between GMMs was first presented in [delon2020wasserstein].
-The main advantage is that, when the OT plan is restricted to the set of GMMs, i.e, $\gamma\in\Gamma(P,Q)\cap\text{GMM}_{\infty}(d)$, there is a discrete equivalent formulation in terms of the GMM components,
-$\omega^{\star}=\underset{\omega\in\Gamma(\pi^{(P)},\pi^{(Q)})}{\text{arg min}}\,\sum_{i=1}^{n}\sum_{j=1}^{m}\omega_{ij}\mathbb{W}_{2}(P_{i},Q_{j})^{2}.$
-$\omega^{\star}\in\mathbb{R}^{n\times m}$ is a component-to-component OT plan.
-The term $\mathbb{W}_{2}(P_{i},Q_{j})$ is the Wasserstein distance between Gaussian components, which has closed-form [takatsu2011wasserstein] in terms of their parameters,
-$\mathbb{W}_{2}(P,P^{\prime})^{2}=\lVert\mu-\mu^{\prime}\rVert_{2}^{2}+\text{Tr}(\Sigma+\Sigma^{\prime}-2(\Sigma^{\frac{1}{2}}\Sigma^{\prime}\Sigma^{\frac{1}{2}})^{\frac{1}{2}}).$
-Based on these ideas, one may define a Wasserstein-type distance between GMMs,
-$\mathbb{MW}_{2}(P,Q)^{2}=\underset{\omega\in\Gamma(\pi^{(P)},\pi^{(Q)})}{\text{inf}}\,\sum_{i=1}^{n}\sum_{j=1}^{m}\omega_{ij}\mathbb{W}_{2}(P_{i},Q_{j})^{2}.$ (6)
-2.3 Wasserstein Barycenter and its Variants
-In the metric setting, the barycenter problem is known as Fréchet [frechet1948elements] or Karcher [karcher1977riemannian] means.
-In our case, given a finite family of measures $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}$ and a set of barycentric coordinates $\lambda\in\Delta_{K}$, we define the Wasserstein barycenters over $(\mathcal{P}_{p,ac}(\Omega),\mathbb{W}_{p})$ through the following optimization problem,
-$P^{\star}=\underset{P\in\mathcal{P}(\Omega)}{\text{arg min}}\,\biggr\{\mathbb{B}_{Q}(P)=\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{p}(P,Q_{k})^{p}\biggr\}.$ (7)
-In general, equation 7 does not have a closed-form solution.
-However, there are algorithms for computing barycenters for $p=2$, when the measures in $\mathcal{P}$ are either empirical measures [cuturi2014fast], Gaussian measures [agueh2011barycenters, alvarez2016fixed] or Gaussian mixtures [delon2020wasserstein, montesuma2024lighter], and a further link can be made to multi-marginal OT [pass2015multi].
-The work of [cuturi2014fast] is of particular interest to us, since theirs was the first algorithm, based on gradient descent, to optimize equation 7 on empirical measures.
-The idea is to initialize $z_{0,1}^{(P)},\cdots,z_{0,n}^{(P)}$ randomly (e.g., from a Gaussian measure), and iterate
-$z_{\tau+1,i}^{(P)}=(1-\alpha)z_{\tau,i}^{(P)}+\alpha\sum_{k=1}^{K}\lambda_{k}T_{\gamma_{k}}(z_{\tau,i}^{(P)}),$ (8)
-where $T_{\gamma_{k}}(z_{i}^{(P)})=n\sum_{j=1}^{n}\gamma_{k,i,j}z_{j}^{(Q_{k})}$ is the barycentric map between $Q_{k}$ and $P$.
-This strategy is reminiscent of the fixed-point approach of [alvarez2016fixed].
-Furthermore, [montesuma2023multi] shows that, for feature-label joint measures, the labels can be propagated through [redko2019optimal],
-$y_{\tau+1,i}^{(P)}=(1-\alpha)y_{\tau,i}^{(P)}+\alpha\sum_{k=1}^{K}\lambda_{k}T_{\gamma_{k}}(y_{\tau,i}^{(P)}),$ (9)
-where $y_{\tau,i}^{(P)}$ are one-hot encoded vectors.
-Previous work has extended equation 8 to the GMM setting, when the components are axis-aligned Gaussians, i.e., $\Sigma_{i}^{(P)}=\text{diag}(\sigma_{i}^{(P)})$. In this case,
-$\mu_{i,\tau+1}^{(P)}=(1-\alpha)\mu_{i,\tau}^{(P)}+\alpha\sum_{k=1}^{K}\lambda_{k}T_{\omega_{k}}(\mu_{i,\tau}^{(P)}),$ (10)
-$\sigma_{i,\tau+1}^{(P)}=(1-\alpha)\sigma_{i,\tau}^{(P)}+\alpha\sum_{k=1}^{K}\lambda_{k}T_{\omega_{k}}(\sigma_{i,\tau}^{(P)}),$
-There are a few limitations with equations 8 and 10.
-First, the empirical iterations do not scale well to large datasets.
-Indeed, they assume the availability of all samples of $Q_{k}$ per iteration.
-Second, the GMM iterations in equation 10 are restricted to axis-aligned GMM s.
-This paper addresses these gaps using gradient flows.
+$\gamma^{\star}=\underset{\gamma\in\Gamma(P,Q)}{\text{arg inf}}\,\mathbb{E}_{(z,z^{\prime})\sim\gamma}[d(z,z^{\prime})^{2}],$ (5)
+where $\gamma^{\star}$ denotes an OT plan. OT defines the so-called 22-Wasserstein distance, through the infimum value of the optimization problem in equations 4 and 5,
+$\mathbb{W}_{2}(P,Q)^{2}=\underset{\gamma\in\Gamma(P,Q)}{\text{inf}}\,\mathbb{E}_{(z,z^{\prime})\sim\gamma}[d(z,z^{\prime})^{2}].$ (6)
+Equation 6 defines a metric in $\mathcal{P}_{2}(\Omega)$ which bridges the notions of Euclidean (i.e., center of mass) and Wasserstein barycenters.
+An alternative consists of regularizing the OT problem with the entropy of $\gamma$, resulting in the Sinkhorn divergence cuturi2013sinkhorn,
+$\mathbb{W}_{2,\epsilon}(P,Q)^{2}=\underset{\gamma\in\Gamma(P,Q)}{\text{inf}}\,\mathbb{E}[d(z,z^{\prime})^{2}]+\lambda H(\gamma),$ (7)
+where $H(\gamma)=\text{KL}(\gamma|P\otimes Q)$, for $(P\otimes Q)(z)=P(z)Q(z)$, and KL is the Kullback Leibler divergence.
+Equation 7 enjoys many desirable properties.
+First, by relying on the matrix scaling algorithm of sinkhorn1964relationship, it has complexity $\mathcal{O}(n^{2})$ per iteration.
+Second, it can be vectorized in modern GPUs, leading to dramatic speed ups.
+Third, since $H(\gamma)$ is strictly convex, the overall optimization problem in equation 7 is convex, leading to a unique solution.
+2.2 Wasserstein Barycenter
+In the metric setting, the barycenter problem is known as frechet1948elements or karcher1977riemannian means.
+In our case, given a finite family of measures $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}$ and a set of barycentric coordinates $\lambda\in\Delta_{K}=\{a\in\mathbb{R}^{K}_{\geq 0}:\sum_{k}a_{k}=1\}$, we define the Wasserstein barycenter over $(\mathcal{P}_{2}(\Omega),\mathbb{W}_{2})$ through the following optimization problem,
+$\hat{P}^{\star}=\underset{\hat{P}\in\text{Emp}_{n}(\Omega)}{\text{arg min}}\,\biggr\{\mathbb{B}_{2,\epsilon}(\hat{P}|\mathcal{Q})=\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2,\epsilon}(\hat{P},\hat{Q}_{k})^{2}\biggr\}.$ (8)
+[…]
 2.3 Gradient Flows
-In the Euclidean setting, let $F:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a functional. A gradient flow is the solution to the differential equation,
-$\begin{cases}\dot{x}(t)=-\nabla F(x(t))&\text{for }t>0,\\ x(0)=x_{0},\end{cases}$ (11)
-where $x_{0}$ is the initial condition. In $\mathcal{P}_{p,\text{ac}}(\Omega)$, gradient flows correspond to the continuity equation,
-$\partial_{t}P_{t}=-\text{div}(P_{t}\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t})),$ (12)
-where $P_{t}$ is a curve in $\mathcal{P}_{p,\text{ac}}(\Omega)$, and $\mathbb{F}:\mathcal{P}_{p,\text{ac}}(\Omega)\rightarrow\mathbb{R}$ is a functional over probability measures.
-Usually, these functionals take the following form [santambrogio2015optimal],
-$\mathbb{F}(P)=$ $\underbrace{\int G(P(z))dz}_{\mathbb{G}(P)}+\underbrace{\int V(z)dP(z)}_{\mathbb{V}(P)}+$ (13)
-$\underbrace{\int\int U(z,z^{\prime})dP(z)dP(z^{\prime})}_{\mathbb{U}(P)},$
-where $\mathbb{G}$, $\mathbb{V}$, and $\mathbb{U}$ are the internal, potential, and interaction energies, respectively.
-Here, $G:\mathbb{R}\rightarrow\mathbb{R}$ is convex and superlinear, and $V:\Omega\rightarrow\mathbb{R}$ and $U:\Omega^{2}\rightarrow\mathbb{R}$ are convex and sufficiently smooth.
-Henceforth, we denote $\mathbb{F}^{\star}=\text{inf}_{P\in\mathcal{P}_{2,\text{ac}}(\Omega)}\mathbb{F}(P)$.
+Our method relies on the gradient flow interpretation of the minimization in equation 8.
+Starting from the Euclidean setting, let $F:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a functional.
+A gradient flow is the solution to the Ordinary Differential Equation (ODE),
+$\dot{x}(t)=-\nabla F(x(t))\text{, subject to }x(0)=x_{0}.$ (9)
+This framework extends to probability measures.
+Let $\mathbb{F}:\mathcal{P}_{2}(\Omega)\rightarrow\mathbb{R}$ be a functional over $\mathcal{P}_{2}(\Omega)$.
+The evolution of a curve $\{P_{t}\}_{t\geq 0}$ in $\mathcal{P}_{2}(\Omega)$ is given by the continuity equation
+$\partial_{t}P_{t}=-\text{div}(P_{t}v_{t}),$ (10)
+where $v_{t}:\Omega\rightarrow\Omega$ is a velocity field, and div is the divergence operator, understood in a measure theoretic sense.
+Choosing $v_{t}=-\nabla_{W}\mathbb{F}(P_{t})$ yields the gradient flow of $\mathbb{F}$, that is, the measure theoretic analogue of equation 9.
+Here, $\nabla_{W}$ denotes the Wasserstein gradient of the functional $\mathbb{F}$ (c.f.
+[chewi2024statistical, Section 5.4] and appendix C.1).
+The structure of $\mathbb{F}$ determines the dynamics of $P_{t}$.
+Following the classical OT theory [santambrogio2015optimal, Chapter 7], we consider the functional $\mathbb{F}(P)=\mathbb{B}(P)+\mathbb{R}(P)$, where,
+$\mathbb{R}(P)=$ $\underbrace{\int G(p(z))dz}_{\text{Internal energy }\mathbb{G}(P)}+\underbrace{\int V(z)dP(z)}_{\text{Potential energy }\mathbb{V}(P)}+$ (11)
+$\underbrace{\int\int U(z,z^{\prime})dP(z)dP(z^{\prime})}_{\text{Interaction energy }\mathbb{U}(P)}$
+where $p:\Omega\rightarrow\mathbb{R}$ denotes the density of $P$, $G:\mathbb{R}\rightarrow\mathbb{R}$, $V:\Omega\rightarrow\mathbb{R}$, and $U:\Omega\times\Omega\rightarrow\mathbb{R}$.
+For example, when $(s)=s\log s, 𝔾\mathbb{G}$ is the entropy functional of [chizat2025doubly], associated with diffusion dynamics.
+$\mathbb{G}$, $\mathbb{V}$, and $\mathbb{U}$ introduce diffusion, drift, and pairwise behaviors in the flow, respectively.
+See Section 3.3 for further examples.
 3 Wasserstein Barycenters as Gradient Flows
-In this section, we describe a new method for computing empirical and Gaussian mixture Wasserstein barycenters. Our main idea is using the functional:
-$\mathbb{F}(P)=\mathbb{B}_{\mathcal{Q}}(P)+\mathbb{G}(P)+\mathbb{V}(P)+\mathbb{U}(P),$ (14)
-where $\mathbb{B}_{\mathcal{Q}}(P)$ is the barycenter objective defined in equation 7, and $\mathbb{G},\mathbb{V}$ and $\mathbb{U}$ are the energies defined in equation 13.
-3.1 Empirical Flow
-In the empirical case, the barycenter is approximated through a finite set of particles (see equation 1) $\{z_{i}^{(P)}\}_{i=1}^{n}$. Hence, we may rewrite equation 14 as,
-$\underset{\begin{subarray}{c}z_{1}^{(P)},\cdots,z_{n}^{(P)}\in\Omega,\\ \gamma_{k}\in\Gamma(\hat{P},\hat{Q}_{k})\end{subarray}}{\text{min}}\,\underbrace{\sum_{k=1}^{K}\lambda_{k}\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}d(z_{i}^{(P)},z_{j}^{(Q_{k})})^{p}}_{\mathbb{B}_{\mathcal{Q}}(\hat{P})}$+
-$\underbrace{\dfrac{1}{n}\sum_{i=1}^{n}V(z_{i}^{(P)})}_{\mathbb{V}(\hat{P})}+\underbrace{\dfrac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}U(z_{i}^{(P)},z_{j}^{(P)})}_{\mathbb{U}(\hat{P})}$ (15)
-Here, we omit the internal energy term, $\mathbb{G}(P)$, as it is not defined for empirical measures. We solve the previous optimization problem through a block-coordinate descent strategy, that is,
-$\gamma_{k}$ $=\underset{\gamma\in\Gamma(\hat{P}_{\tau},\hat{Q}_{k})}{\text{arg min}}\,\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{i,j}d(z_{\tau,i}^{(P)},z_{j}^{(Q_{k})})^{p},$
-$z_{\tau+1,i}^{(P)}$ $=z_{\tau,i}^{(P)}-\alpha\nabla\mkern-10.0mu\nabla\mathbb{F}(\hat{P}_{\tau}).$ (16)
-In equation 16, $\gamma_{k}$ is used to compute $\mathbb{B}_{\mathcal{Q}}$ and is updated for each $\tau$.
-We provide in Algorithm 1 the pseudo-code for this strategy.
-Our algorithm assumes access to $Q_{k}$ through sampling, that is, we do not necessarily have access to all samples of $Q_{k}$ at once.
-Algorithm 1 Empirical barycenter Wasserstein flow using Gradient Descent.
+In this section, we describe a new method for computing empirical OT barycenters. We use,
+[…]
+[…]
+First, we derive the empirical version of the continuity equation (c.f., Equation 10). For $\hat{P}_{t}\in\text{Emp}_{n}(\Omega)$:
+$\partial_{t}\hat{P}_{t}$ $=\frac{1}{n}\sum_{i=1}^{n}\partial_{t}\delta_{z_{t,i}^{(P)}}=-\dfrac{1}{n}\sum_{i=1}^{n}\text{div}(\dot{z}_{t,i}^{(P)}\delta_{z_{t,i}^{(P)}}),$ (13)
+$=-\text{div}(\hat{P}_{t}v_{t}),$
+where we identified $v_{t}(z_{t,i}^{(P)})=\dot{z}_{t,i}^{(P)}$. Second, we discretize the gradient flow in time, through an index $\tau=\alpha t$, $\alpha>0$. Using the forward Euler scheme,
+$\dot{z}_{t,i}^{(P)}=\frac{z_{\tau+1,i}^{(P)}-z_{\tau,i}^{(P)}}{\alpha}\rightarrow z_{\tau+1,i}^{(P)}=z_{\tau,i}^{(P)}+\alpha v_{\tau,i},$ (14)
+where $v_{\tau,i}=v_{\tau}(z_{\tau,i}^{(P)})$. We can obtain different update strategies based on the velocity field,
+$v_{\tau,i}$ $=-\nabla_{z_{\tau,i}^{(P)}}\mathbb{F}(\hat{P}_{\tau}),$ (15)
+$v_{\tau,i}$ $=-\nabla_{z_{\tau,i}^{(P)}}\mathbb{F}(\hat{P}_{\tau})+\sqrt{2\alpha\eta_{G}}\eta_{\tau,i},$ (16)
+$v_{\tau,i}$ $=u_{\tau,i},\quad u_{\tau+1,i}=\beta u_{\tau,i}-\alpha\nabla_{z_{\tau,i}^{(P)}}\mathbb{F}(\hat{P}_{\tau}).$ (17)
+Equations 15, 16 and 17 correspond to the steepest descent, Langevin dynamics, and momentum descent strategies, respectively.
+$\eta_{G}\geq 0$ denotes the diffusion coefficient.
+The Langevin dynamics version of the flow corresponds to the NPGD dynamics of chizat2025doubly.
+Furthermore, for empirical measures, the functionals in Equation 12 are tractable, and take the following form,
+[…]
+[…]
+[…]
+[…]
+where $\hat{p}$ denotes the density of $\hat{P}$.
+Since $\hat{P}$ is empirical it is, in general, difficult to compute $\mathbb{G}$.
+A notable exception is the entropy functional (i.e., $G(s)=s\log s$), which adds a diffusion term to the gradient flow dynamics.
+This corresponds to Langevin dynamics [chizat2025doubly] (Equation 16).
+For general choices of $G$, one needs to estimate the density $\hat{p}$ at each iteration of the empirical flow.
+Algorithm 1 Empirical Time Discretized Wasserstein Gradient Flow
 […]
 2:Barycenter support $\{z_{i}^{(P)}\}_{i=1}^{n}$
-3:for $\tau\leftarrow 1$ to $n_{\text{iter}}$ do
-4:  $\mathbb{F}(\hat{P}_{\tau})\leftarrow\mathbb{V}(\hat{P}_{\tau})+\mathbb{U}(\hat{P}_{\tau})$
-5:  for $k\leftarrow 1$ to $K$ do
-6:   Sample $\{z_{i}^{(Q_{k})}\}_{i=1}^{m}$ i.i.d. from $Q_{k}$.
-7:   $\mathbb{F}(\hat{P}_{\tau})\leftarrow\mathbb{F}(\hat{P}_{\tau})+\lambda_{k}\mathbb{W}_{p}(\hat{P}_{\tau},\hat{Q}_{k})^{p}$
-8:  end for
-9:  $z_{\tau+1,i}^{(P)}\leftarrow z_{\tau,i}^{(P)}-\alpha\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{\tau})$
-10:end for
-Remark 3.1.
-In line 5 of Algorithm 1, $\hat{Q}_{k}$ denotes the empirical measure (c.f., equation 1) supported on the mini-batch $\{z_{i}^{(Q_{k})}\}_{i=1}^{m}$ sampled from $Q_{k}$.
-3.2 Gaussian Mixture Flow
-In the case of GMMs, we flow the parameters $\theta=\{(\pi_{k},\mu_{k},\Sigma_{k})\}_{k=1}^{K}$ of $P_{\theta}=\sum\pi_{i}\mathcal{N}(\cdot|\mu_{i},\Sigma_{i})$. Similarly to equation 16, we have:
-$\omega_{k}$ $=\underset{\omega\in\Gamma(\pi,\pi_{k})}{\text{arg min}}\,\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\omega_{i,j}\mathbb{W}_{2}(P_{\theta_{\tau},i},Q_{k,j})^{2},$
-$\theta_{\tau+1,i}$ $=\theta_{\tau,i}-\alpha\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{\theta_{\tau}}).$ (17)
-Here, $\omega_{k}$ is a transport plan between GMM components.
-In general, the energy functionals do not have closed form with respect to the GMM parameters.
-However, we can estimate them using Monte-Carlo and the reparametrization trick, for instance,
-$z_{i}$ $=L_{i}\varepsilon+\mu_{i},\quad i\sim\pi,\varepsilon\sim\mathcal{N}(\cdot|0,\text{Id}),$ (18)
-This strategy allows us to propagate the gradients to the parameters of $P_{\tau}$. We show the overall approach in Algorithm 2.
-Algorithm 2 GMM barycenter flow using Gradient Descent.
 […]
-2:Barycenter support $\{z_{i}^{(P)}\}_{i=1}^{n}$
-3:for $\tau\leftarrow 1$ to $n_{\text{iter}}$ do
-4:  Sample $\{z_{i}=L_{i}\varepsilon+\mu_{i}:i\sim\pi,\varepsilon\sim\mathcal{N}(\cdot|0,\text{Id})\}$,
-5:  $\mathbb{F}(\hat{P}_{\tau})\leftarrow\mathbb{G}(\hat{P}_{\tau})+\mathbb{V}(\hat{P}_{\tau})+\mathbb{U}(\hat{P}_{\tau})$
+4:for $\tau\leftarrow 1$ to $n_{\text{iter}}$ do
+5:  $\mathbb{F}(\hat{P}_{\tau})\leftarrow\mathbb{V}(\hat{P}_{\tau})+\mathbb{U}(\hat{P}_{\tau})+\mathbb{G}(\hat{P}_{\tau})$
 6:  for $k\leftarrow 1$ to $K$ do
-7:   $\mathbb{F}(P_{\theta_{\tau}})\leftarrow\mathbb{F}(P_{\theta_{\tau}})+\lambda_{k}\mathbb{MW}_{2}(P_{\theta_{\tau}},Q_{k})^{2}$
-8:  end for
-9:  $z_{\tau+1,i}^{(P)}\leftarrow z_{\tau,i}^{(P)}-\alpha\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{\theta_{\tau}})$
-10:end for
-Covariance matrix parametrization.
-Optimizing $\mathbb{F}$ with respect to $\Sigma_{k}$ is challenging, due to the symmetric positive definite constraint.
-We enforce this constraint by writing $\Sigma_{k}=L_{k}L_{k}^{T}$, where $L_{k}$ is the lower-triangular Cholesky factor.
-There are two advantages to this strategy.
-First, optimizing with respect to $L_{k}$ is empirically more stable than $\Sigma_{k}$.
-Second, we can use $L_{k}$ directly in the reparametrization trick in equation 18.
+7:   Sample $\{z_{i}^{(Q_{k})}\}_{i=1}^{m}$ i .i.d. from $Q_{k}$.
+8:   $\mathbb{F}(\hat{P}_{\tau})\leftarrow\mathbb{F}(\hat{P}_{\tau})+\lambda_{k}\mathbb{W}_{2,\epsilon}(\hat{P}_{\tau},\hat{Q}_{k})^{2}$
+9:  end for
+10:  $z_{\tau+1,i}^{(P)}\leftarrow z_{\tau,i}^{(P)}+\alpha v_{\tau,i}$
+11:end for
+An immediate consequence of Equation 18 is that our gradient flow formulation encompasses several previous works.
+For instance, with $G=U=V=0$, we have the classical free-support barycenters of cuturi2014fast.
+Furthermore, for the entropy functional and $\epsilon>0$, we have the NPGD algorithm of chizat2025doubly.
+Finally, as we show in Section 3.2 below, with a specific metric on $\Omega=\mathcal{X}\times\mathcal{Y}$, we retrieve the algorithm of montesuma2023multi.
+3.1 Scaling with Mini-Batch OT
+The computation of $\mathbb{B}_{\epsilon}$ in the discrete approach [cuturi2014fast] involves an OT problem between the complete set of samples in the input measures.
+Therefore, when the input measures are large scale, the barycenter problem becomes infeasible.
+We circumvent this issue with mini-batch OT [fatras2021minibatch].
+The main idea behind our scalable barycenter algorithm is accessing $\hat{Q}_{k}\in\mathcal{Q}$ through sampling.
+At each iteration of Algorithm 1, we sample a mini-batch $\{z_{i}^{(Q_{k})}\}_{i=1}^{m}$ of $m$ samples from each input measure (cf.
+line 5 in Algorithm 1).
+Since we sample the same number of samples from each measure, we can actually vectorize the computation of the $K$ OT problems in the evaluation of $\mathbb{W}_{2,\epsilon}$.
+Let $\{z_{\tau,i}^{(P)}\}_{i=1}^{n}$ be the barycenter support, and $\{z_{j}^{(Q_{k})}\}_{j=1}^{m}$ be a mini-batch from $Q_{k}$.
+The Gibbs kernel $E_{k,i,j}=e^{-(C_{k,i,j}/\epsilon)}$, and $C_{k,i,j}=d(z_{\tau,i}^{(P)},z_{j}^{(Q_{k})})^{2}$ are arrays with shape $(K,n,m)$.
+The Sinkhorn iterations become,
+$\psi_{k}^{(\ell+1)}=\dfrac{b}{E_{k}^{T}\phi_{k}^{(\ell)}}\text{, and }\phi_{k}^{(\ell+1)}=\dfrac{a_{k}}{E_{k}\psi_{k}^{(\ell+1)}},$ (19)
+where the division operation should be understood coordinate-wise.
+Furthermore, $b_{j}=n^{-1}$ and $a_{k,i}=m^{-1}$ is the importance of each sample in the barycenter, and input measure support, respectively.
+For simplicity, we assume uniform weights.
+For $L\in\mathbb{N}$ iterations, we can retrieve $\gamma_{\epsilon,k}^{\star}=\text{diag}(\phi_{k}^{(L)})E_{k}\text{diag}(\psi_{k}^{(L)})$.
+Since each $\gamma_{\epsilon,k}^{\star}$ has shape $(n,m)$, we can actually stack it in an array of shape $(K,n,m)$, similarly to $E$ and $C$.
+The main advantage of equation 19 is that the $K$ OT problems can be vectorized.
+This is possible because we draw the same number of samples from each input measure, making it possible to vectorize the Sinkhorn iterations.
+For instance, a similar strategy is possible for measures over a fixed grid [peyre2019computational, Remark 4.16].
+The OT plan $\gamma_{\epsilon}$, obtained at the end of the Sinkhorn iterations, is used to compute $\mathbb{B}_{\epsilon}$, in lines 4 through 7 in Algorithm 1.
 3.2 Flows over Joint Measures
-In [cuturi2014fast], the authors considered barycenters for $p=2$, $\Omega=\mathbb{R}^{d}$, and $d(z,z^{\prime})=\lVert z-z^{\prime}\rVert_{2}$.
-Therefore, we generalize this setting, since we only need the differentiability of $d$ for our algorithm to work.
-As we show in our experiments, in machine learning applications one encounters $\Omega=\mathcal{X}\times\mathcal{Y}$, i.e., measures over the joint space of features and labels.
-Whenever we deal with labeled measures (i.e., $z=(x,y)$), we adopt,
+An interesting advantage of the gradient flow formalism is that the only necessary ingredient is a differentiable metric over $\Omega$.
+Therefore, we generalize the setting of cuturi2014fast, who assumed $\Omega=\mathbb{R}^{d}$, and $d(z,z^{\prime})=\lVert z-z^{\prime}\rVert_{2}$.
+As we show in our experiments, in machine learning applications one has $\Omega=\mathcal{X}\times\mathcal{Y}$, i.e., measures over the joint space of features and labels.
+Let $z=(x,y),x\in\mathcal{X},y\in\mathcal{Y}$,
 $d(z,z^{\prime})=\sqrt{\lVert x-x^{\prime}\rVert_{2}^{2}+\beta\lVert y-y^{\prime}\rVert_{2}^{2}},$ (20)
@@
 However, for classification, $\mathcal{Y}$ is categorical (e.g., $\mathcal{Y}=\{1,\cdots,n_{\text{classes}}\}$).
-One possible strategy, used in [montesuma2021wasserstein] and [alvarez2021dataset], is fixing the labels and flowing only the features.
-In contrast to these works, we embed $\mathcal{Y}$ into the compact continuous space $\Delta_{K}$, through a one-hot encoding operation.
+One possible strategy, used in montesuma2021wasserstein and alvarez2021dataset, is fixing the labels and flowing only the features.
+In contrast to these works, we embed $\mathcal{Y}$ into the compact continuous space $\Delta_{n_{\text{classes}}}$, through a one-hot encoding operation.
 For our flow, we parametrize labels through a change of variables,
 $y_{i,c}^{(P)}=\text{softmax}(\ell_{i,1}^{(P)},\cdots,\ell_{i,n_{c}}^{(P)})=\dfrac{\exp(\ell_{i,c}^{(P)})}{\sum_{c=1}^{n_{c}}\exp(\ell_{i,c}^{(P)})},$
-thus, instead of optimizing over $z=(x,y)$, we optimize over $z=(x,\ell)$.
-From the soft probabilities, we can retrieve the actual discrete labels with an argmax, $y_{i}^{(P)}=\text{argmax}_{c=1,\cdots,n_{c}}y_{i,c}^{(P)}$.
-For GMMs, we equip their components with labels $\nu_{k}\in\Delta_{n_{\text{classes}}}$. Using the ground-cost in equation 27, we are able to write the $\mathbb{MW}_{2}(P,Q)^{2}$ as,
-$\sum_{i=1}^{n}\sum_{j=1}^{m}\omega_{ij}^{\star}(\mathbb{W}_{2}(P_{i,x},Q_{j,x})^{2}+\beta\lVert\nu_{i}^{(P)}-\nu_{j}^{(Q)}\rVert_{2}^{2}),$
-where $P_{i,x}=\mathcal{N}(\cdot|\mu_{i}^{(P)},\Sigma_{i}^{(P)})$ is the Gaussian feature marginal.
-This approach was first proposed in [montesuma2024lighter], but the authors relied on heuristic arguments to justify this modeling choice.
-We present the following proposition justifying the label term,
-Proposition 3.1.
-Let $P=\sum\pi_{i}^{(P)}(\mathcal{N}(\mu_{i}^{(P)},\Sigma_{i}^{(P)})\otimes\delta_{\nu_{i}^{(P)}})$ and $Q=\sum\pi_{j}^{(Q)}(\mathcal{N}(\mu_{j}^{(Q)},\Sigma_{j}^{(Q)}\otimes\delta_{\nu_{j}^{(Q)}})$ be two GMMs over $\Omega=\mathcal{X}\times\mathcal{Y}$.
-Let the ground cost $c$ be,
-$c(z,z^{\prime})=\lVert x-x^{\prime}\rVert_{2}^{2}+\rho(y,y^{\prime})^{2},$
-where $\rho(y,y^{\prime})$ is a metric over $\mathcal{Y}$. Then,
-$\mathbb{MW}_{2}(P,Q)^{2}=\underset{\omega\in\Gamma(\pi^{(P)},\pi^{(Q)})}{\text{min}}\,\sum_{i=1}^{n}\sum_{j=1}^{m}\omega_{ij}C_{ij},$
-where $C_{ij}=\mathbb{W}_{2}(P_{i,x},Q_{j,x})^{2}+\rho(\nu_{i}^{(P)},\nu_{j}^{(P)})^{2}$.
-Functionals for joint measures.
+thus, instead of optimizing over $z=(x,y)$, we optimize over $z=(x,\ell)$, $x\in\mathbb{R}^{d}$, $\ell\in\mathbb{R}^{n_{\text{classes}}}$.
+From the soft probabilities, we can retrieve the actual discrete labels with an argmax, $y_{\text{hard},i}^{(P)}=\text{argmax}_{c=1,\cdots,n_{c}}y_{i,c}^{(P)}$.
+3.3 Task-Aware Regularizing Functionals
 One of the advantages of our proposed method is regularizing the barycenter calculation with internal, interaction, and potential energy functionals.
-This idea was already used in practice by [alvarez2021dataset] for transfer learning problems.
+This idea was already used in practice by alvarez2021dataset for transfer learning problems.
 Here we propose the following functionals,
-$V_{E}(z)$ $=-\sum_{c=1}^{n_{\text{classes}}}y_{c}\log y_{c},$ (20)
-$U_{R}(z,z^{\prime})$ $=\begin{cases}h(d(x,x^{\prime}))&\text{if }y\neq y^{\prime},\\ 0&\text{otherwise}\end{cases},$ (21)
-where $h:\mathbb{R}\rightarrow\mathbb{R}$ is lower semi-continuous and bounded from below.
+$V_{E}(z)$ $=-\eta_{V}\sum_{c=1}^{n_{\text{classes}}}y_{c}\log y_{c},$ (21)
+$U_{R}(z,z^{\prime})$ $=\begin{cases}\eta_{U}h(d(x,x^{\prime}))&\text{if }y_{\text{hard}}\neq y_{\text{hard}}^{\prime},\\ 0&\text{otherwise}\end{cases},$ (22)
+where $h:\mathbb{R}\rightarrow\mathbb{R}$ is lower semi-continuous and bounded from below, and $\eta_{V}>0$ and $\eta_{U}>0$ are scaling parameters.
 For instance, in our experiments we use the hinge loss, $h(u)=\text{max}(0,\text{margin}-d)$, where $\text{margin}\geq 0$ is a margin parameter.
-Equations 20 and 21 correspond to entropy and repulsion terms.
+Equations 21 and 22 correspond to entropy and repulsion, respectively.
 The first functional penalizes barycenters that have fuzzy labels.
@@
 3.4 Convergence
-One of the difficulties in analyzing the gradient flow of the functional in equation 14 comes from the fact that $P\mapsto\mathbb{W}_{p}(P,Q)^{p}$ is not geodesically convex in $\mathcal{P}_{p,ac}(\Omega)$ [santambrogio2017euclidean, Section 4.4].
+One of the difficulties in analyzing the gradient flow of the functional in equation 12 comes from the fact that $P\mapsto\mathbb{W}_{2}(P,Q)^{2}$ is not geodesically convex in $\mathcal{P}_{2}(\Omega)$ [santambrogio2017euclidean, Section 4.4].
 As a result, we are minimizing a non-convex functional.
-Henceforth, we focus on $p=2$, as this is the setting where most results are available.
-Explicit constants and proofs are available in Section 4 in the Appendix.
+Explicit constants and proofs are available in Appendix D.
 Our analysis relies on a measure-theoretic version of the Polyak Łojasiewicz (PL) inequality [polyak1964gradient, karimi2016linear],
-$\lVert\nabla\mkern-10.0mu\nabla\mathbb{B}(P)\rVert_{L_{2}(P)}\geq C_{\text{PL}}(\mathbb{B}(P)-\mathbb{B}^{\star}),$ (22)
-which is covered in [chewi2024statistical]. Under this assumption, we have the following convergence result,
+$\lVert\nabla_{W}\mathbb{B}(P)\rVert_{L_{2}(P)}^{2}\geq C_{\text{PL}}(\mathbb{B}(P)-\mathbb{B}^{\star}).$ (23)
+When this inequality holds, we have the following convergence result.
 Theorem 3.1.
-Let $\Omega=\mathcal{B}(0,R)$ be the closed ball on $\mathbb{R}^{d}$ with radius $R>0$.
-Let $P^{\star}\in\mathcal{P}_{2,\text{ac}}(\mathcal{B}(0,R))$ be the barycenter of $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K},Q_{k}\in\mathcal{P}_{2,\text{ac}}(\mathcal{B}(0,R))$ $k=1,\cdots,K$, with barycentric coordinates $\lambda=(\lambda_{1},\cdots,\lambda_{K})\in\Delta_{K}$.
-Assume the inequality 47.
-Then, the following holds,
-$\begin{aligned} \mathbb{E}[\hat{\mathbb{B}}(P_{\tau})-\hat{\mathbb{B}}^{\star}]\leq e^{-C_{\text{PL}}\tau}(\mathbb{B}(P_{0})-\mathbb{B}^{\star})+C_{R}\sqrt{\dfrac{C_{d,m}}{n}}\end{aligned},$ (23)
-where the expectation on the l.h.s. is taken with respect to samples from $Q_{k}$. The constants $C_{R}$ and $C_{d,m}$ depend on the radius $R$, number of barycenter samples $m$, and dimensions $d$.
-The right hand side of inequality 23 is composed of two terms.
-The first term comes from the PL inequality, and covers the convergence towards the minimizer of $\mathbb{B}$.
-The second term covers the error of empirical approximations.
-In this sense, letting $\tau\rightarrow+\infty$ leads to an error governed by the empirical approximation.
-Next, we derive a new result for the convergence of the GMM gradient flow (cf.
-Algorithm 2).
-Here, we assume $P$ and $Q_{k}\in\mathcal{Q}$ are GMMs.
-Furthermore, we denote by $\hat{Q}_{k}$ the GMM fitted on data from $Q_{k}$ with the expectation-maximization algorithm.
-Our result relies on [boite2025differentiable, Theorem 3.1].
-Therefore, we assume,
-  1. 1.
-    bounded GMM parameters, i.e., $\lVert\mu_{i}^{(Q_{k})}\lVert_{2}\leq R_{\mu}$, $\sqrt{\text{Tr}(\Sigma_{i}^{(Q_{k})})}\leq R_{\Sigma}$ $\forall i,\forall k$ (resp.
-$\hat{\mu}_{i}^{(Q_{k})}$ and $\hat{\Sigma}_{i}^{(Q_{k})}$, the estimated parameters); furthermore,
-  2. 2.
-    Empirical convergence rates, i.e., $\mathbb{E}[\lVert\pi^{(Q_{k})}-\hat{\pi}^{(Q_{k})}\rVert_{1}]\leq\rho_{\pi}$, $\mathbb{E}[\lVert\mu_{i}^{(Q_{k})}-\hat{\mu}_{i}^{(Q_{k})}\rVert_{2}]\leq\rho_{\mu}$, and $\mathbb{E}[d_{\text{Bures}}(\Sigma_{i}^{(Q_{k})},\hat{\Sigma}_{i}^{(Q_{k})})]\leq\rho_{\Sigma}$, $\forall i$ and $\forall k$.
-Here the expectations are taken with respect to samples drawn from $Q_{k}$. Under these conditions we have,
-Theorem 3.2.
-Let $P$ and $Q$ be two labeled GMMs with bounded parameters and satisfying the empirical convergence rates. Then, the following holds,
-$\mathbb{E}[\hat{\mathbb{B}}(P_{\tau})-\hat{\mathbb{B}}^{\star}]\leq e^{-C_{\text{PL}\tau}}(\mathbb{B}(P_{0})-\mathbb{B}^{\star})+C_{\lambda,\mathcal{Q}},$
-where $C_{\lambda,\mathcal{Q}}$ is a constant that depends on the coordinates $\lambda$ and GMMs in $\mathcal{Q}$.
-Here $C_{\lambda,\mathcal{Q}}$ depends on the $(\pi^{(Q_{k})},\mu^{(Q_{k})},\Sigma^{(Q_{k})})$ of each GMM, but not on $\nu^{(Q_{k})}$.
-Indeed, since we fit a GMM per class, the labels are estimated exactly (i.e., $P_{i,x}$ and $\hat{P}_{i,x}$ belong to the same class).
+Let $\Omega=\mathcal{B}(0,R)$ be the closed ball in $\mathbb{R}^{d}$ with radius $R>0$.
+Let $P^{\star}\in\mathcal{P}_{2}(\Omega)$ be the barycenter of $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K},Q_{k}\in\mathcal{P}_{2}(\Omega)$ $k=1,\cdots,K$, with barycentric coordinates $\lambda=(\lambda_{1},\cdots,\lambda_{K})\in\Delta_{K}$.
+Approximate each $Q_{k}$ with $m$ i.i.d.
+samples $\{z_{i}^{(Q_{k})}\}_{i=1}^{m}$, $z_{i}^{(Q_{k})}\sim Q_{k}$ and $\{z_{\tau,i}^{(P)}\}_{i=1}^{n}$ be the barycenter support.
+Under inequality 42 the following holds,
+$\begin{aligned} \mathbb{E}[\hat{\mathbb{B}}(\hat{P}_{\tau})-\hat{\mathbb{B}}^{\star}]\leq e^{-C_{\text{PL}}\tau}(\mathbb{B}(\hat{P}_{0})-\mathbb{B}^{\star})+C_{R}\sqrt{\dfrac{C_{d,n}}{m}}\end{aligned},$ (24)
+where the expectation is taken with respect to samples from $Q_{k}$.
+The constants $C_{R}$ and $C_{d,n}$ depend on the radius $R$, number of barycenter samples $n$, and dimensions $d$.
+$m$ denotes the number of samples drawn from each $Q_{k}.$
+The first term in the r.h.s.
+of equation 24 decays exponentially with the iterations $\tau$, and is a direct consequence of the PL inequality in equation 23.
+The second term captures the empirical approximation error of the involved measures, that is, $Q_{k}\in\mathcal{Q}$ and $P_{\tau}$.
+In practice, $\tau\rightarrow+\infty$ leads to an error governed by the empirical approximations.
+On the PL Inequality.
+Equation 23 is a standard tool used in proving the convergence of non-convex optimization algorithms, however, it is an open question whether it holds in general settings.
+For instance, chewi2020gradient shows that this inequality holds for Guassian measures under specific spectral conditions.
+In the Appendix, we show that the same result holds for location-scatter families of measures (Appendix D.2), based on the isometry between these sets, and the Bures-Wasserstein manifold established by [alvarez2016fixed, Theorem 2.3].
+On top of this theoretical justification, we provide an experiment validating the bound on the Swiss roll measure (Appendix D.1).
+(a) $Q_{0}$.
+(b) $Q_{1}$.
+(c) $Q_{2}$.
+(d) $Q_{3}$.
+(e) $Q_{4}$.
+(f) $P^{\star}$.
+Figure 2: Swiss roll measures in a location-scatter family, where a source measure is transformed under an affine function $T_{k}(x)=A_{k}x+b_{k}$.
+Colors reflect the relative position of points in the manifold.
 4 Experiments
-We divide our experiments into a toy example (Section 4.1) to illustrate our methods, and multi-source domain adaptation experiments (Section 4.2).
-Due to length constraints, we include additional experiments in Sections 5 and 6 in the Appendix.
-In particular, in Table 1 in the Appendix, we show a running-time comparison demonstrating the scalability of our method.
-4.1 Toy Example
-In Figure 2, we show the Swiss roll measure $Q_{0}$, alongside four variations obtained via $T_{k,\sharp}Q_{0}$, $k=1,\cdots,4$.
-This example has been used in several state-of-the-art Wasserstein barycenter works [korotin2021continuous, visentin2025computing].
-Originally there are labels associated with the samples of the Swiss-roll measure, corresponding to the position of the points in the underlying manifold.
-We show a summary of our results in Figure 3.
-Figure 2: Location-scatter family generated by a Swiss-roll measure, $Q_{0}$. Each measure $Q_{k}=T_{k,\sharp}Q_{0}$, for $T_{k}(x)=A_{k}x+b_{k}$.
-First, we compare unsupervised barycenter solvers in Figures 3 (a) through (h).
-Quantitatively, empirical methods, notably [cuturi2014fast, Algorithm 2] and our WGF algorithm (cf.
+In the following, all experiments were conducted on a virtual machine with 8 CPUs (AMD EPYC 7413, 48GB of RAM) and an NVIDIA L4 GPU (24GB of VRAM).
+Appendix F includes additional details, fine-grained results, complexity analysis and ablations.
+4.1 Swiss Roll Measures
+In Figure 2, we show the Swiss roll measure [korotin2021continuous] $Q_{0}$, alongside four variations obtained via $T_{k,\sharp}Q_{0}$, $k=1,\cdots,4$, $T_{k}(x)=A_{k}x+b_{k}$, for which the ground-truth barycenter is known in closed-form [alvarez2016fixed], that is, $P^{\star}=(\sum_{k}\lambda_{k}T_{k})_{\sharp}Q_{0}$.
+These measures are shown in Figure 2.
+(a) Discrete.
+(b) CW2B.
+(c) U-NOT.
+(d) NormFlow.
+(e) WGF.
+[…]
+Figure 3: Comparison of unsupervised barycenter methods. Solvers (b, c, d) are neural, and (e, f) are ours.
+We compare 5 barycenter solvers, including discrete solvers [cuturi2014fast, montesuma2023multi], and neural solvers: CW2B [korotin2021continuous], U-NOT gazdieva2024robust, and NormFlow visentin2025computing.
+First, we benchmark unsupervised solvers in Figures 3 (a) through (f).
+Quantitatively, empirical methods, notably [cuturi2014fast, Algorithm 2] and our Wasserstein Gradient Flow (WGF) algorithm (cf.
 Algorithm 1) achieve the lowest Wasserstein distance to the ground truth in Figure 2.
-We note that, overall, while the neural network solvers are usually more scalable than discrete methods in terms of number of samples, their optimization problem is more complicated and very sensitive to hyper-parameters.
-(a) CW2B [korotin2021continuous].
-(b) NOT [kolesov2024estimating].
-(c) Discrete [cuturi2014fast].
-(d) WGF.
+Overall, while the neural network solvers are usually more scalable than discrete methods in terms of number of samples, their optimization is more complicated and very sensitive to hyper-parameters.
+Second, we experiment with integrating labels in the ground cost as described in Section 3.2.
+These methods are shown in Figures 4 (a), (b) and (c).
+In all cases, integrating labels produces barycenters that are closer to the ground-truth.
+We conclude that using the labels gives a strong inductive bias in the barycenter computation, which explains the gain in performance of labeled barycenters in the next section.
+(a) Discrete.
+(b) WGF.
 […]
-[…]
-(g) U-NOT [gazdieva2024robust].
-(h) NormFlow [visentin2025computing].
-(i) Discrete [montesuma2023multi].
-(j) WGF.
-[…]
-[…]
-Figure 3: Comparison between Wasserstein barycenter solvers.
-Colored scatter plots indicate labeled barycenters.
-For each solver, we compute the Wasserstein distance between its solution $\hat{P}$ and the ground-truth $P^{\star}$, shown in the title of each sub figure (best seen on screen).
-Overall, using label information leads to barycenters that better approximate the ground-truth barycenter.
-Second, we experiment with integrating labels in the ground cost as described in Section 3.3.
-These methods are shown in Figures 3 (i) through (l).
-In all cases, integrating labels produces barycenters that are closer to the ground-truth.
-We conclude that using the labels gives a strong inductive bias in the barycenter computation, which helps explain the gain in performance of labeled barycenters in the next section.
+Figure 4: Comparison of supervised barycenter methods.
+Next, we compare the running time of our WGF solver with the discrete solver of cuturi2014fast, for an increasing barycenter support size $n\in\{2^{10},\cdots,2^{16}\}$ and batch size $m\in\{2^{8},2^{9},2^{10}\}$ for $200$ iterations.
+The discrete solver is run for $n\in\{2^{10},\cdots,2^{14}\}$.
+For this algorithm, larger support sizes results in memory overflow.
+Our WGF is able to compute larger barycenters via mini-batching.
+We summarize our results in Figure 5.
+First, in Figure 5 (a), our WGF leads to speedups ranging from $2\times$ $(n=2^{12},m=2^{10})$ to $50\times$ $(n=2^{14},m=2^{8})$.
+Similar gains were obtained using Exact OT on CPU (Appendix F.3).
+Second, in Figure 5 (b), we isolate the gain of using GPU acceleration and our vectorization strategy (Section 3.1), showing consistent gains ranging from $2.1\times$ to $8.3\times$.
+Overall, our WGF strategy provides compounding speedups from three sources: (i) minibatching, (ii) entropic regularization, and (iii) GPU parallelism.
+(a) Scaling with $m$ ($\epsilon>0$)
+(b) Scaling with $\epsilon$ ($m=256$)
+Figure 5: Running time analysis for our WGF algorithm.
+In (a), we compare the running time of our technique in comparison with the discrete solver of cuturi2014fast.
+In (b), we isolate the effect of GPU acceleration, showing consitent gains with increasing support size.
 4.2 Multi-Source Domain Adaptation
-Problem Formulation.
 One of the main applications of Wasserstein barycenters is multi-source domain adaptation (MSDA) [kouw2019review, sun2015survey, pan2009survey].
 In this setting one needs to adapt multiple labeled source measures $Q_{1},\cdots,Q_{K}$ to a single unlabeled target measure $Q_{T}$.
-The goal is to learn, from samples $\{\{x_{i}^{(Q_{k})},y_{i}^{(Q_{k})}\}_{i=1}^{n_{k}}\}_{k=1}^{K}$ and $\{x_{i}^{(Q_{T})}\}_{i=1}^{n_{T}}$, a classifier $h$ that achieves low risk in the target domain measure,
+The goal is to learn, from samples $\{\{x_{i}^{(Q_{k})},y_{i}^{(Q_{k})}\}_{i=1}^{n_{k}}\}_{k=1}^{K}$ and $\{x_{i}^{(Q_{T})}\}_{i=1}^{n_{T}}$, a classifier $h$ that achieves low risk or error in the target domain measure,
 $\mathcal{R}_{Q_{T}}(h)=\mathbb{E}_{(x,y)\sim Q_{T}}[\mathcal{L}(y,h(x))],$
@@
 Thus, we assume that a meaningful feature extractor $\phi$, called the backbone, has been previously learned.
-We obtain the feature extractor by fine-tuning a neural network on the labeled source domain data.
-In addition, similar to [montesuma2021wasserstein, montesuma2024lighter], we align the barycenter with the target through OT [courty2016optimal].
-More details are available in Section 6 in the Appendix.
-Benchmark Backbone # Samples # Domains # Dim. # Classes
-Office31 ResNet50 [he2016deep] 3287 3 2048 31
-BCI-CIV-2a CBraMod [wang2024cbramod] 5184 10 200 4
-TEP CNN [montesuma2024benchmarking] 17289 6 128 29
-Office Home ResNet101 [he2016deep] 15500 4 2048 65
-ISRUC CBraMod [wang2024cbramod] 89240 100 512 5
-Table 1: Overview of benchmarks used for domain adaptation, sorted by number of samples.
-Experimental Setting.
+We obtain the feature extractor by fine-tuning a neural network on the labeled source domain data (Appendix E.2).
+Experimental Setup.
+The unsupervised WGF is fit with unlabeled measures, i.e., $Q_{k}=n_{k}^{-1}\sum_{i}\delta_{x_{i}^{(Q_{k})}}$, where $x_{i}^{(Q_{k})}$ are embedding vectors.
+The supervised WGF is fit with labeled measures, that is, $Q_{k}=n_{k}^{-1}\sum_{i=1}^{n_{k}}\delta_{(x_{i}^{(Q_{k})},y_{i}^{(Q_{k})})}$.
+In both cases, $\hat{P}=n^{-1}\sum_{i=1}^{n}\delta_{(x_{i}^{(P)},y_{i}^{(P)})}$ which allows us to apply the regularizing functionals defined in section 3.3.
+Wasserstein Barycenter Transport.
+montesuma2021wasserstein proposed a technique based on Wasserstein barycenters for domain adaptation.
+The idea is to map the synthesized barycenter support to the target domain through the barycentric mapping [courty2016optimal].
+Then, a classifier can be learned on the target domain based on the transported data.
+We adapt this idea to unsupervised barycenter algorithms by using the computed barycenter as a pivot domain.
+See appendix E.4 for more details.
+Benchmarks.
 We run our experiments on five benchmarks: Office 31 [saenko2010adapting], Office Home [venkateswara2017deep], BCI-CIV-2a [brunner2008bci], ISRUC [khalighi2013automatic], and TEP [montesuma2024benchmarking].
@@
 We show in Table 1 an overview of our experimental setting.
+Benchmark Backbone # Samples # Domains # Dim. # Classes
+Office31 ResNet50 3287 3 2048 31
+BCI-CIV-2a CBraMod 5184 10 200 4
+TEP CNN 17289 6 128 29
+Office Home ResNet101 15500 4 2048 65
+ISRUC CBraMod 89240 100 512 5
+Table 1: Overview of benchmarks used for domain adaptation, sorted by number of samples.
 Compared Methods.
 Overall, we compare seven Wasserstein barycenter strategies with ours.
-These methods are: discrete barycenters [cuturi2014fast, Algorithm 2], normalizing flows [visentin2025computing], continuous 2-Wasserstein barycenters [korotin2021continuous], neural OT barycenters [kolesov2024estimating], and unbalanced neural OT barycenters [gazdieva2024robust].
-We also include labeled barycenter strategies [montesuma2023multi, montesuma2024multi].
-A brief review of these methods is available in Section 2 in the Appendix.
+Besides those used in the previous section, we include the NOT approach of kolesov2024estimating, and the Gaussian Mixture Model (GMM) barycenter of montesuma2024lighter.
 For completeness, we include four other state-of-the-art methods in domain adaptation over embedding vectors.
-Those methods are: WJDOT [turrisi2022multi], DaDiL-R and E [montesuma2023multi], and GMM-DaDiL [montesuma2024lighter].
+Those methods are: WJDOT turrisi2022multi, DaDiL-R, and E montesuma2023multi, and GMM-DaDiL montesuma2024lighter.
 For each benchmark, we use one domain as the target domain (e.g., Amazon vs.
-$\{\text{dSLR, Webcam}\}$ in the Office 31 benchamrk), and we measure the classification accuracy, i.e., the percentage of correct predictions.
-Benchmark $\mathcal{X}\times\mathcal{Y}$ Office31 OfficeHome BCI-CIV-2a ISRUC TEP
-Backbone - ResNet-50 ResNet-101 CBraMod CBraMod CNN
-Source-Only - 86.40 75.95 50.30 76.63 78.48
-WJDOT [turrisi2022multi] - 86.80 76.59 N/A 76.95 86.13
-DaDiL-R [montesuma2023multi] - 89.91 77.86 53.41 74.68 86.14
-DaDiL-E [montesuma2023multi] - 89.79 78.14 N/A 75.89 85.87
-GMM-DaDiL [montesuma2024lighter] - 90.63 78.81 57.10 75.47 86.85
-Discrete [cuturi2014fast] 81.94 70.42 57.38 70.67 83.81
-NormFlow [visentin2025computing] 85.91 76.73 55.45 78.04 82.89
-CW2B [korotin2021continuous] 86.37 76.44 57.25 75.84 85.83
-NOT [kolesov2024estimating] 86.22 75.36 57.29 76.39 84.93
-U-NOT [gazdieva2024robust] 86.97 76.85 57.17 77.06 85.43
-Discrete [montesuma2023multi] 87.93 77.09 57.43 78.20 86.09
-GMM [montesuma2024lighter] 88.54 77.87 57.04 74.58 84.67
-WGF (ours) 88.14 77.83 57.50 79.78 86.21
-WGF-GMM (ours) 89.31 78.71 57.35 78.72 86.77
-Table 2: Average classification accuracy on target domains.
-In total, we compare eight methods against our WGF strategy.
-For barycenter solvers, we indicate by ✓methods that compute barycenters on $\Omega=\mathcal{X}\times\mathcal{Y}$.
-Bold numbers represent the best barycenter method, and underlined numbers represent second best.
+$\{\text{dSLR, Webcam}\}$ in the Office 31 benchmark), and we measure the classification accuracy, i.e., the percentage of correct predictions.
 Main Results.
 We present our main results in Table 2, which reports the average performance per domain on each benchmark.
-We provide fine-grained results in Section 6 in the Appendix.
-Overall, labeled barycenter methods (the four last rows in Table 2) have a clear advantage with respect to the unsupervised methods.
-We argue that label information is pivotal in domain adaptation success, which is consistent with previous research [courty2016optimal, montesuma2021wasserstein, montesuma2023multi, montesuma2024lighter].
-Among the labeled barycenter methods, our gradient flow framework achieves either the best performance, or remains competitive, surpassing previous methods in MSDA, such as WJDOT and DaDiL in the ISRUC benchmark.
-Figure 4: t-SNE [maaten2008visualizing] visualization of the barycenter of $\text{Art},\text{Product},\text{Real-World}]$ source domains in the Office home benchmark.
-Colors represent different classes from $1$ to $65$.
-Overall, the classes tend to be more separated when using the repulsion functional $\mathbb{U}$.
-Visualization.
-We now visualize the effect of adding the repulsion interaction energy functional on the Office-Home benchmark.
-Adding $\mathbb{U}$ leads to classes that are well separated.
-For computer vision experiments we use ResNets [he2016deep] as the backbone, which produce $2048$-dimensional embeddings.
-For that reason, we use the cosine distance $d(x,x^{\prime})=1-\text{cossim}(x,x^{\prime})$ in equation 21 and we fix $\text{margin}=1$.
-This choice encourages class clusters to live in orthogonal sub-spaces of the latent space.
-We show a comparison of the obtained barycenters in Figure 4.
-Overall, the classes tend to have less overlap when the interaction energy is added to the barycenter functional.
-Ablation.
-We ablate the effect of $\mathbb{V}$ and $\mathbb{U}$ in the Office Home benchmark.
-For this benchmark, we use $\mathbb{V}(P)=\mathbb{W}_{2}(\hat{P},\hat{Q}_{T})^{2}$.
-The intent here is to isolate the contribution of each component to the state-of-the-art results in Table 2.
-Our results are summarized in Table 3, where we further compare our results with other related barycenter methods.
-In the empirical setting, the use of Algorithm 1 alone already presents a gain in performance compared to its discrete counterpart.
-Overall, the combination of $\mathbb{B}+\mathbb{V}+\mathbb{U}$ consistently gives the best performance, showing the importance of adding underlying structure to the Wasserstein barycenter in domain adaptation.
-Method $\mathcal{X}\times\mathcal{Y}$ $\mathbb{B}$ $\mathbb{B}+\mathbb{V}$ $\mathbb{B}+\mathbb{U}$ $\mathbb{B}+\mathbb{V}+\mathbb{U}$
-Discrete [cuturi2014fast] 70.42 N/A N/A N/A
-Discrete [montesuma2023multi] 77.09 N/A N/A N/A
-WGF (ours) 77.48 77.07 77.08 77.83
-GMM [montesuma2024lighter] 77.87 N/A N/A N/A
-WGF-GMM (ours) 77.28 78.28 78.12 78.71
-Table 3: Ablation of average target domain classification accuracy on the Office Home benchmark.
+We provide fine-grained results in Appendix F.
+In general, labeled barycenter methods (the 3 last rows in Table 2) have a clear advantage over unsupervised methods.
+We argue that label information is essential in domain adaptation success, which is consistent with previous research [courty2016optimal, montesuma2021wasserstein, montesuma2023multi, montesuma2024lighter].
+Among barycenter methods, our WGF method achieves the best performance in all benchmarks, outperforming previous methods in MSDA, such as WJDOT and DaDiL in the ISRUC, BCI-CIV, and TEP benchmarks.
+Benchmark $\mathcal{X}\times\mathcal{Y}$ Office31 OfficeHome BCI-CIV-2a ISRUC TEP Avg. Rank
+Backbone - ResNet-50 ResNet-101 CBraMod CBraMod CNN -
+Source-Only - 86.40 75.95 50.30 76.63 78.48 -
+WJDOT - 86.80 76.59 N/A 76.95 86.13 7.8
+DaDiL-R - 89.91 77.86 53.41 74.68 86.14 6.4
+DaDiL-E - 89.79 78.14 N/A 75.89 85.87 5.0
+GMM-DaDiL - 90.63 78.81 57.10 75.47 86.85 4.4
+Discrete 87.82 77.17 57.38 70.67 83.81 8.8
+NormFlow 85.91 76.73 55.45 78.04 82.89 10.0
+CW2B 86.37 76.44 57.25 75.84 85.83 9.0
+NOT 86.22 75.36 57.29 76.39 84.93 9.4
+U-NOT 86.97 76.85 57.17 77.06 85.43 7.8
+[…]
+Discrete 87.93 77.09 57.43 78.20 86.09 5.0
+GMM 88.54 77.87 57.04 74.58 84.67 8.2
+[…]
+Table 2: Average classification accuracy on target domains.
+The top group (WJDOT through GMM-DaDiL) include general domain adaptation methods.
+The middle and bottom groups show barycenter solvers without and with access to source labels ($\mathcal{X}\times\mathcal{Y}=$✗, $\mathcal{X}\times\mathcal{Y}=$✓), respectively.
+Bold indicates best among barycenter methods.
+Unsupervised vs.
+Supervised Barycenters.
+An intriguing finding of Table 2 is that, on some benchmarks, unsupervised WGF achieves decent performance, even outperforming supervised baselines (e.g., WJDOT of turrisi2022multi).
+We probe the reason why these methods have unreasonably good performance, by performing a t-SNE [maaten2008visualizing] visualization of the barycenter against the source domain samples, shown in Figure 6.
+We see that even without labels, the barycenter samples cluster around the classes.
+This effect stems from the geometry of the embeddings, which were optimized (through supervision on source domains) to separate classes.
+As a result, unsupervised OT already captures, to some extent, the cluster structure of the barycentric measure.
+Figures 6 (b) and (c) investigate the remaining performance gap, as we transport the measures towards the target.
+Figure 6 (b) shows the transported sources to the barycentric domain (left) and the synthesized barycenter (right), and Figure 6 shows the final transported samples in the target domain for the unsupervised strategy (left) and supervised one (right).
+We conclude that using a synthesized supervised barycenter is better than using an unsupervised barycenter as pivot domain (Appendix E.4).
+Therefore, the inductive bias of using labels in the ground cost is essential for domain adaptation.
+This further explains the performance of DaDiL methods, which directly synthesize the target domain through an approximate barycenter.
+(a) Unsupervised vs. Supervised barycenters.
+(b) Transported sources vs. supervised barycenter.
+(c) Transported samples in the target domain.
+Figure 6: Comparison of unsupervised vs. supervised barycenter transports on the TEP benchmark.
 5 Conclusion
-In this paper we have introduced a new framework for computing Wasserstein barycenters, based on Wasserstein gradient flows [ambrosio2008gradient, santambrogio2017euclidean].
-Our approach addresses scalability issues in discrete barycenter solvers [cuturi2014fast], and is capable of regularizing the underlying barycenter through internal, potential, and interaction functionals.
-We presented gradient flow Algorithms 1 and 2, for empirical and Gaussian mixture measures, respectively.
-We further proved convergence guarantees based on the PL inequality.
-We empirically tested our method against existing discrete [cuturi2014fast, montesuma2023multi] and neural-network [korotin2021continuous, kolesov2024estimating, gazdieva2024robust, visentin2025computing]-based barycenter solvers on computer vision, neuroscience, and chemical engineering benchmarks, demonstrating that our proposed methods consistently achieve state-of-the-art performance.
-Our empirical findings (Table 2) further demonstrate that incorporating label information in the OT objective is key for domain adaptation.
-This finding highlights an interesting gap in the literature, where neural net-based solvers are not capable of exploiting this information, and thus have sub-optimal performance in domain adaptation.
-We leave an investigation of this issue for future work.
-Computing Wasserstein Barycenters through Gradient Flows
+We have presented a new Wasserstein barycenter algorithm (Algorithm 1), inspired by the gradient flow literature santambrogio2015optimal, that addresses two persistent challenges: scalability and regularization.
+We have derived a new algorithm that is amenable to mini-batch OT fatras2021minibatch (Section 3.1).
+Under entropic OT, we further demonstrate that OT computations can be vectorized across the $K$ input measures, leading to further speedups (Figure 5).
+Our framework is flexible, and accommodates modular plug-in regularization from internal, potential, and interaction energies (Section 3.3).
+Extensive experiments on domain adaptation demonstrate the value of our approach and that incorporating labels in the OT ground-cost is necessary for state-of-the-art performance (Table 2).
+Finally, our work opens interesting research questions.
+First, our framework can be extended to more complex differentiable structures, such as Riemannian manifolds.
+Second, it opens the question on how to effectively integrate labels and other regularization schemes into neural network solvers.
 Appendix A Introduction
@@
 Section D provides missing proofs.
-Finally, section E and F provide additional experiments.
+Finally, section F provide additional experiments.
 Appendix B A Brief Survey on Wasserstein Barycenters
 B.1 Introduction
-Given a finite family of measures $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}$, the Wasserstein barycenter problem was first introduced by [agueh2011barycenters] as the following minimization problem,
-$P^{\star}=\underset{P\in\mathcal{P}_{2,\text{ac}}(\Omega)}{\text{arg min}}\,\biggr\{\mathbb{B}_{\mathcal{Q}}=\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2}(P,Q_{k})^{2}\biggr\}.$ (24)
-This minimization problem is the analogous, in Wasserstein space, to the Euclidean barycenter problem $x\mapsto\sum_{k}\lambda_{k}d(x,x_{k})^{2}$, which is made possible since $(\mathcal{P}_{2,\text{ac}}(\Omega),\mathbb{W}_{2})$ is a metric space.
-Solving problem 24 is, in general, challenging.
+Given a finite family of measures $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}$, the Wasserstein barycenter problem was first introduced by agueh2011barycenters as the following minimization problem,
+$P^{\star}=\underset{P\in\mathcal{P}_{2}(\Omega)}{\text{arg min}}\,\biggr\{\mathbb{B}_{\epsilon}(P|\mathcal{Q})=\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2,\epsilon}(P,Q_{k})^{2}\biggr\},$ (25)
+where $\epsilon\geq 0$ denotes the entropic regularization weight.
+This minimization problem is the analogous, in Wasserstein space, to the Euclidean barycenter problem $x\mapsto\sum_{k}\lambda_{k}d(x,x_{k})^{2}$, which is made possible since $(\mathcal{P}_{2}(\Omega),\mathbb{W}_{2})$ is a metric space.
+Solving problem 25 is, in general, challenging.
 Except for the Gaussian case, there are no closed-form solutions.
@@
 Discrete barycenter solvers are based on the empirical approximation of the underlying measures, that is,
-$\hat{P}(z)=\sum_{i=1}^{n}a_{i}\delta(z-z_{i}^{(P)}).$ (25)
+$\hat{P}(z)=\sum_{i=1}^{n}a_{i}\delta(z-z_{i}^{(P)})\in\text{Emp}_{n}(\Omega).$ (26)
 Here, $\{z_{i}^{(P)}\}_{i=1}^{n}$ is the support of $\hat{P}$ and $a=\{a_{i}:a_{i}\geq 0\text{, and }\sum a_{i}\}_{i=1}^{n}$ are the sample weights.
@@
 In this paper, we deal primarily with free-support Wasserstein barycenters.
-Arguably, the first free-support algorithm was proposed by [cuturi2014fast].
-Their strategy comes from plugging equation 25 into 24,
-$Z^{(P),\star}=\underset{\begin{subarray}{c}z_{1}^{(P)},\cdots,z_{n}^{(P)}\in\Omega,\\ \gamma_{k}\in\Gamma(\hat{P},\hat{Q}_{k})\end{subarray}}{\text{arg min}}\,\biggr\{\mathbb{B}_{\mathcal{Q}}(P)=\sum_{k=1}^{K}\lambda_{k}\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}\lVert z_{i}^{(P)}-z_{j}^{(Q_{k})}\rVert_{2}^{2}\biggr\}.$ (26)
+Arguably, the first free-support algorithm was proposed by cuturi2014fast.
+Their strategy comes from plugging equation 26 into 25,
+$Z^{(P),\star}$ $=\underset{z_{1}^{(P)},\cdots,z_{n}^{(P)}\in\Omega}{\text{arg min}}\,\biggr\{\mathbb{B}_{\epsilon}(P|\mathcal{Q})=\sum_{k=1}^{K}\lambda_{k}\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{\epsilon,k,i,j}\lVert z_{i}^{(P)}-z_{j}^{(Q_{k})}\rVert_{2}^{2}\biggr\},$ (27)
+subject to $\gamma_{\epsilon,k}^{\star}=\underset{\gamma\in\Gamma(P,Q_{k})}{\text{arg min}}\,\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{ij}\lVert z_{i}^{(P)}-z_{j}^{(Q)}\rVert_{2}^{2}+\epsilon\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{ij}\log\gamma_{ij}$
 At this point, one should compare equation 27 with equation 18 in the main paper.
-Cuturi and Doucet [cuturi2014fast] propose solving this equation via a block coordinate descent strategy.
+cuturi2014fast propose solving this equation via a block coordinate descent strategy.
 Indeed, assume $\gamma_{k}=\text{OT}(P,Q_{k})$ has been calculated (e.g., through linear programming), then,
-$\nabla_{z_{i}^{(P)}}\mathbb{B}_{\mathcal{Q}}(P)=2\sum_{k=1}^{K}\lambda_{k}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}(z_{i}^{(P)}-z_{j}^{(Q_{k})})=\dfrac{2}{n}z_{i}^{(P)}-2\sum_{k=1}^{K}\lambda_{k}\sum_{j=1}^{n}\gamma_{k,i,j}z_{j}^{(Q_{k})}.$
-As a consequence, $\nabla_{z_{i}^{(P)}}\mathbb{B}_{\mathcal{Q}}(P)=0$ leads to an update rule in terms of the barycentric projection of $P$ to $Q_{k}$, which is a discrete approximation of the Monge map [deb2021rates].
-An important question comes when the samples from $P$ are feature-label joints. This is question was first raised by [montesuma2023multi], who proposed to use the ground-cost,
+$\nabla_{z_{i}^{(P)}}\mathbb{B}_{\epsilon}(P|\mathcal{Q})=2\sum_{k=1}^{K}\lambda_{k}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}(z_{i}^{(P)}-z_{j}^{(Q_{k})})=\dfrac{2}{n}z_{i}^{(P)}-2\sum_{k=1}^{K}\lambda_{k}\sum_{j=1}^{n}\gamma_{k,i,j}z_{j}^{(Q_{k})}.$
+As a consequence, $\nabla_{z_{i}^{(P)}}\mathbb{B}_{\epsilon}(P|\mathcal{Q})=0$ leads to an update rule in terms of the barycentric projection of $P$ to $Q_{k}$, which is a discrete approximation of the Monge map deb2021rates.
+An important question comes when the samples from $P$ are feature-label joints, first raised by montesuma2023multi, who proposed to use the ground-cost,
 $d(z,z^{\prime})=\sqrt{\lVert x-x^{\prime}\rVert_{2}^{2}+\beta\lVert y-y^{\prime}\rVert_{2}^{2}},$ (28)
@@
 As we show in the main paper, since this ground-cost is a metric, computing the OT cost with $d^{2}$ as the ground-cost is equivalent to computing the 22-Wasserstein distance on $\mathcal{X}\times\mathcal{Y}$.
-The first paper to remark these computations was [montesuma2023multi], who proposed to couple these mappings,
-$x_{i,\tau+1}^{(P)}=\sum_{k=1}^{K}\lambda_{k}T_{P\rightarrow Q_{k}}^{\star}(x_{i,\tau}^{(P)})\text{, and, }y_{i,\tau+1}^{(P)}=\sum_{k=1}^{K}\lambda_{k}T_{P\rightarrow Q_{k}}^{\star}(y_{i,\tau}^{(P)}).$ (28)
+The first paper to remark these computations was montesuma2023multi, who proposed to couple these mappings,
+$x_{\tau+1,i}^{(P)}=\sum_{k=1}^{K}\lambda_{k}\biggr(\underbrace{n\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}x_{j}^{(Q_{k})}}_{T_{P\rightarrow Q_{k}}^{\star}(x_{\tau,i}^{(P)})}\biggr)\text{, and, }y_{\tau+1,i}^{(P)}=\sum_{k=1}^{K}\lambda_{k}\biggr(\underbrace{n\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}y_{j}^{(Q_{k})}}_{T_{P\rightarrow Q_{k}}^{\star}(y_{\tau,i}^{(P)})}\biggr).$ (29)
 These iterations have a few nice properties with respect to the label term.
-First, the mappings $T_{P\rightarrow Q_{k}}^{\star}(y_{i,\tau}^{(P)})$ correspond to label propagation terms first defined in [redko2019optimal].
-Second, as a consequence of the first point, the iterations respect the constraint $y_{i,\tau}^{(P)}\in\Delta_{n_{\text{classes}}}$.
-Third, these iterations can be understood as fixed-point iterations in the sense of [alvarez2016fixed].
+First, the mappings $T_{P\rightarrow Q_{k}}^{\star}(y_{i,\tau}^{(P)})$ correspond to label propagation terms first defined in redko2019optimal.
+Second, since the summations in equation 29 are convex combinations of the labels, the update rules naturally suffice $y_{\tau,i}^{(P)}\in\Delta_{n_{\text{classes}}}$.
+Third, these iterations can be understood as fixed-point iterations in the sense of alvarez2016fixed.
 The authors use this algorithm as the building block of their dictionary learning strategy.
-The main drawback of the iterations in equation 28 is when the input measures have imbalanced classes.
-As we show in the additional experiments in Section E.2, OT, due the mass preservation constraints, is forced to match samples from different classes, leading to deformed barycenters with fuzzy labels (see Figure 8).
 B.3 Neural Barycenter Solvers
 In this section, we cover a collection of methods that use neural networks as the foundation of computing Wasserstein barycenters.
-These methods are oftentimes referred to as ”continuous”, thus relying on the idea that the calculated barycenter is a continuous measure.
+These methods are oftentimes referred to as "continuous", thus relying on the idea that the calculated barycenter is a continuous measure.
 Note that this is different than claiming that these barycenters solve a continuous problem.
 In fact, most methods are formulated in terms of expectations of the involved measures, which are approximated through mini-batching.
+Indeed, mini-batching is the main tool for scalability in Wasserstein barycenters.
 Generative Modeling.
-One of the first methods was proposed by [cohen2020estimating], who proposed viewing equation 24 through the lens of generative modeling.
+One of the first methods was proposed by cohen2020estimating, who proposed viewing equation 25 through the lens of generative modeling.
 Let $g_{\theta}$ denote a neural network with parameters $\theta$.
@@
 Therefore, the Wasserstein barycenter is parametrized via the weights of $g_{\theta}$.
-In comparison, methods such as [cuturi2014fast] and [montesuma2023multi] parametrize the barycenter through its samples.
+In comparison, methods such as cuturi2014fast and montesuma2023multi parametrize the barycenter through its samples.
 The underlying idea is sampling from each $Q_{k}$ and $P_{0}=\mathcal{N}(0,\sigma^{2}I)$.
-Input Convex Neural Nets. More refined settings exploit the properties of the 2-Wasserstein distance with the squared Euclidean ground-cost. Indeed, note that,
+Input Convex Neural Nets.
+More refined settings exploit the properties of the 2-Wasserstein distance with the squared Euclidean ground-cost.
+Indeed, from [makkuva2020optimal, Equation 4], exploiting the decomposition
+$\lVert z-z^{\prime}\rVert_{2}^{2}=\lVert z\rVert_{2}^{2}+\lVert z^{\prime}\rVert_{2}^{2}-2\langle z,z^{\prime}\rangle,$
+we can obtain,
 $\dfrac{1}{2}\mathbb{W}_{2}(P,Q)^{2}=\underbrace{\int\dfrac{\lVert z\rVert_{2}}{2}dP(z)+\int\dfrac{\lVert z\rVert_{2}}{2}dQ(z)}_{\mathbb{C}(P,Q)}-\underset{f,g\in\Phi}{\text{min}}\,\int f(z)dP(z)+\int g(z)dQ(z),$ (30)
-where $(f,g)\in\Phi=\{f(x)+g(y)\leq\lVert x-y\rVert_{2}^{2}\}$ are called potentials. The main insight of [makkuva2020optimal] is rewritting the minimization problem in the r.h.s. of equation 29 as,
+where $(f,g)\in\Phi=\{f(x)+g(y)\leq\lVert x-y\rVert_{2}^{2}\}$ are called potentials. The main insight of makkuva2020optimal is rewritting the minimization problem in the r.h.s. of equation 30 as,
 $\underset{f\in\text{CVX}}{\text{sup}}\,\underset{g\in\text{CVX}}{\text{inf}}\,\underbrace{-\int f(z)dP(z)-\int(\langle z,\nabla g(z)\rangle-f(\nabla g(z)))dQ(z)}_{V_{P,Q}(f,g)}$
-Based on these equations, [fan2020scalable] propose solving,
+Based on these equations, fan2020scalable propose solving,
 […]
@@
 Indeed, these are the potentials of the transportation problem between the barycenter $P$ and each $Q_{k}$.
-More specifically, In light of Brenier’s theorem [brenier1991polar], $\nabla g_{k}$ is the Monge map from $Q_{k}$ to the barycenter $P$.
-These elements motivate [fan2020scalable] using ICNNs, as in [makkuva2020optimal], for the optimization problem in equation 30.
+More specifically, In light of Brenier’s theorem brenier1991polar, $\nabla g_{k}$ is the Monge map from $Q_{k}$ to the barycenter $P$.
+These elements motivate fan2020scalable using Input Convex Neural Networks (ICNNs), as in makkuva2020optimal, for the optimization problem in equation 31.
 Continuous 2-Wasserstein Barycenter.
-We now cover a family of methods relying on the fixed-point view of Wasserstein barycenters introduced by [alvarez2016fixed].
+We now cover a family of methods relying on the fixed-point view of Wasserstein barycenters introduced by alvarez2016fixed.
 Indeed, these authors show that the Wasserstein barycenter is a fixed point of the operator $\Psi:P\mapsto(\sum_{k}\lambda_{k}T_{P\rightarrow Q_{k}})_{\sharp}P$.
@@
 Now, with respect to OT potentials, if $\Psi(P)=P$, then one has,
-$\sum_{k=1}^{K}\lambda_{k}T_{Q_{k}\rightarrow P}^{\star}(z)=x\overset{(a)}{\implies}\sum_{k=1}^{K}\lambda_{k}\nabla g_{k}(z)=z\overset{(b)}{\implies}\sum_{k=1}^{K}\lambda_{k}g_{k}(z)=\dfrac{\lVert z\rVert_{2}^{2}}{2}+c$
+$\sum_{k=1}^{K}\lambda_{k}T_{Q_{k}\rightarrow P}^{\star}(z)=z\overset{(a)}{\implies}\sum_{k=1}^{K}\lambda_{k}\nabla g_{k}(z)=z\overset{(b)}{\implies}\sum_{k=1}^{K}\lambda_{k}g_{k}(z)=\dfrac{\lVert z\rVert_{2}^{2}}{2}+c$
 where implication (a) follows from Brenier’s theorem, and $g_{k}$ is the potential of the transportation problem from $Q_{k}$ to $P$, and implication (b) follows by integration.
 Here, since $c$ is an integration constant, one can conveniently take $c=0$.
-Now, denote $\text{Congruent}(\lambda)=\{(g_{1},\cdots,g_{K}):\sum_{k}\lambda_{k}g_{k}(z)=\nicefrac{{\lVert z\rVert_{2}^{2}}}{{2}}\}$ be the set of congruent potentials $(g_{1},\cdots,g_{K})$ with respect to coordinates $\lambda$.
+Now, denote $\text{Congruent}(\lambda)=\{g_{1:K}=(g_{1},\cdots,g_{K}):\sum_{k}\lambda_{k}g_{k}(z)=\nicefrac{{\lVert z\rVert_{2}^{2}}}{{2}}\}$ be the set of congruent potentials $g_{1:K}=(g_{1},\cdots,g_{K})$ with respect to coordinates $\lambda$.
 Using the dual formulation in equation 30 and restricting the set of potentials to congruent potentials,
-$\mathbb{B}_{\mathcal{Q}}(P)=\sum_{k=1}^{K}\lambda_{k}\mathbb{E}_{z\sim Q_{k}}\biggr[\dfrac{\lVert z\rVert_{2}^{2}}{2}\biggr]+\mathbb{E}_{z\sim P}\biggr[\dfrac{\lVert z\rVert_{2}^{2}}{2}\biggr]-\underset{\{g_{k}\}_{k=1}^{K}\in\text{Congruent}(\lambda)}{\text{min}}\,\biggr(\sum_{k=1}^{K}\lambda_{k}(\mathbb{E}_{z\sim Q_{k}}[f_{k}(z)]+\mathbb{E}_{z\sim P}[g_{k}])(z)\biggr).$
-Now, taking the expectation on the congruence constraint, we have,
+$\mathbb{B}(P|\mathcal{Q})=\sum_{k=1}^{K}\lambda_{k}\mathbb{E}_{z\sim Q_{k}}\biggr[\dfrac{\lVert z\rVert_{2}^{2}}{2}\biggr]+\mathbb{E}_{z\sim P}\biggr[\dfrac{\lVert z\rVert_{2}^{2}}{2}\biggr]-\underset{g_{1:K}\in\text{Congruent}(\lambda)}{\text{min}}\,\biggr(\sum_{k=1}^{K}\lambda_{k}(\mathbb{E}_{z\sim Q_{k}}[f_{k}(z)]+\mathbb{E}_{z\sim P}[g_{k}])(z)\biggr),$
+where $g_{1:K}=(g_{1},\cdots,g_{K})$. Now, taking the expectation on the congruence constraint, we have,
 $\sum_{k=1}^{K}\lambda_{k}\mathbb{E}_{z\sim P}[g_{k}(z)]=\mathbb{E}_{z\sim P}\biggr[\dfrac{\lVert z\rVert_{2}^{2}}{2}\biggr],$
-which allowed [korotin2021continuous] to simplify $\mathbb{B}_{\mathcal{Q}}(P)$ into,
-$\mathbb{B}_{\mathcal{Q}}(P)=\sum_{k=1}^{K}\lambda_{k}\mathbb{E}_{z\sim Q_{k}}\biggr[\dfrac{\lVert z\rVert_{2}^{2}}{2}\biggr]-\underset{\{g_{k}\}_{k=1}^{K}\in\text{Congruent}(\lambda)}{\text{min}}\,\sum_{k=1}^{K}\lambda_{k}\mathbb{E}_{z\sim Q_{k}}[g_{k}(z)],$ (31)
+which allowed korotin2021continuous to simplify $\mathbb{B}(P|\mathcal{Q})$ into,
+$\mathbb{B}(P|\mathcal{Q})=\sum_{k=1}^{K}\lambda_{k}\mathbb{E}_{z\sim Q_{k}}\biggr[\dfrac{\lVert z\rVert_{2}^{2}}{2}\biggr]-\underset{g_{1:K}\in\text{Congruent}(\lambda)}{\text{min}}\,\sum_{k=1}^{K}\lambda_{k}\mathbb{E}_{z\sim Q_{k}}[g_{k}(z)],$ (32)
 which does not involve terms with respect to the barycentric measure $P$.
-While the approach of [korotin2021continuous] simplifies the barycenter objective to expectations with respect to the known measures $Q_{k}$, it introduces a challenging constraint – the congruence of potentials.
+While the approach of korotin2021continuous simplifies the barycenter objective to expectations with respect to the known measures $Q_{k}$, it introduces a challenging constraint – the congruence of potentials.
 This constraint is challenging, as it is a functional inequality over the whole space $\Omega$.
 Furthermore, it adds to the already constrained setting of fitting convex functions.
-[korotin2021continuous] proposes solving it approximately, by penalizing potentials that do not meet the congruence condition.
+korotin2021continuous proposes solving it approximately, by penalizing potentials that do not meet the congruence condition.
 There is a circular dependence here, as one needs an estimate of the Wasserstein barycenter to estimate the constraint violation.
-Neural (Weak) Optimal Transport. Weak OT [gozlan2017kantorovich, backhoff2019existence] comes from the interpretation of the Kantorovich formulation as,
+Neural (Weak) Optimal Transport. Weak OT gozlan2017kantorovich, backhoff2019existence comes from the interpretation of the Kantorovich formulation as,
 $\gamma^{\star}=\underset{\gamma\in\Gamma(P,Q)}{\text{arg inf}}\,\int_{\Omega}C(z,\gamma(\cdot|z))dP(z)\text{, where, }C(z,Q)=\int_{\Omega}c(z,z^{\prime})dQ(z^{\prime}),$ (33)
 this formulation allows for more general costs $C:\Omega\times\mathcal{P}(\Omega)\rightarrow\mathbb{R}$, i.e., costs that take a sample and a probability measure as inputs.
-In this sense, [kolesov2024estimating] extends equation 31 to,
-$\mathbb{B}_{\mathcal{Q}}(P)=\sum_{k=1}^{K}\lambda_{k}\biggr(\mathbb{E}_{z\sim Q_{k}}[C_{k}(z,\gamma(\cdot|z))]-\mathbb{E}_{z\sim Q_{k}}[\mathbb{E}_{z^{\prime}\sim\gamma(\cdot|z)}(g_{k}(z^{\prime}))]\biggr),$ (33)
+In this sense, kolesov2024estimating extends equation 32 to,
+$\mathbb{B}(P|\mathcal{Q})=\sum_{k=1}^{K}\lambda_{k}\biggr(\mathbb{E}_{z\sim Q_{k}}[C_{k}(z,\gamma(\cdot|z))]-\mathbb{E}_{z\sim Q_{k}}[\mathbb{E}_{z^{\prime}\sim\gamma(\cdot|z)}(g_{k}(z^{\prime}))]\biggr),$ (34)
 which holds for congruent potentials $\sum_{k}\lambda_{k}g_{k}=0$.
-As with equation 31, $\mathbb{B}_{\mathcal{Q}}$ does not depend explicitly on the unknown barycentric measure $P$.
+As with equation 32, $\mathbb{B}$ does not depend explicitly on the unknown barycentric measure $P$.
 As such, the authors compute the barycenter through a max-min optimization problem,
-$\underset{\{g_{1},\cdots,g_{K}:\sum_{k}\lambda_{k}g_{k}=0\}}{\text{max}}\,\underset{\{\gamma_{1},\cdots,\gamma_{K}:\gamma_{k}(\cdot|z)\in\Gamma(P_{k})\}}{\text{min}}\,\biggr\{\sum_{k=1}^{K}\lambda_{k}\biggr(\mathbb{E}_{z\sim Q_{k}}[C_{k}(z,\gamma_{k}(\cdot|z))]-\mathbb{E}_{z\sim Q_{k}}[\mathbb{E}_{z^{\prime}\sim\gamma_{k}(\cdot|z)}(g_{k}(z^{\prime}))]\biggr)\biggr\}.$
+$\underset{\{g_{1:K}:\sum_{k}\lambda_{k}g_{k}=0\}}{\text{max}}\,\underset{\{\gamma_{1},\cdots,\gamma_{K}:\gamma_{k}(\cdot|z)\in\Gamma(P_{k})\}}{\text{min}}\,\biggr\{\sum_{k=1}^{K}\lambda_{k}\biggr(\mathbb{E}_{z\sim Q_{k}}[C_{k}(z,\gamma_{k}(\cdot|z))]-\mathbb{E}_{z\sim Q_{k}}[\mathbb{E}_{z^{\prime}\sim\gamma_{k}(\cdot|z)}(g_{k}(z^{\prime}))]\biggr)\biggr\}.$
 This extension opens up new possibilities for weak OT barycenters. For instance,
@@
 $C_{\text{KL}}$ is an alternative, when modeling $\gamma(\cdot|z)=\mathcal{N}(\mu(z),\sigma(z))$.
-On top of this reformulation, the authors in [kolesov2024estimating] propose approximating $\gamma(\cdot|z)$ through a mapping, as in [korotin2022neural], that is,
+On top of this reformulation, the authors in kolesov2024estimating propose approximating $\gamma(\cdot|z)$ through a mapping, as in korotin2022neural, that is,
 $\gamma(\cdot|z)=T(z,s),\text{ subject to }s\sim\mathbb{S},$
 where $\mathbb{S}$ is some prior measure, such as $\mathbb{S}=\mathcal{N}(0,\text{I})$.
-In this context, [gazdieva2024robust] extends this idea in the semi-unbalanced OT setting [liero2018optimal], where one of the mass preservation constraint of one of the marginals of $\gamma$ is relaxed.
+In this context, gazdieva2024robust extends this idea in the semi-unbalanced OT setting liero2018optimal, where one of the mass preservation constraint of one of the marginals of $\gamma$ is relaxed.
 More specifically, given $\xi:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$,
@@
 We denote $\bar{\xi}(z^{\prime})=\text{sup}_{z\in\mathbb{R}}\{zz^{\prime}-\xi(z)\}$ to be its convex conjugate.
-With that in mind, [gazdieva2024robust] shows, in Theorem 1 and Corollaries 1 and 2, that minimizing,
-$\underset{P\in\mathcal{P}_{2,\text{ac}}}{\text{arg min}}\,\sum_{k=1}^{K}\lambda_{k}\biggr(\int_{\Omega}\int_{\Omega}c(z,z^{\prime})d\gamma(z,z^{\prime})+D_{\xi}(\gamma(\cdot|z)||Q_{k})\biggr),$
+With that in mind, gazdieva2024robust shows, in Theorem 1 and Corollaries 1 and 2, that minimizing,
+$\underset{P\in\mathcal{P}_{2}}{\text{arg min}}\,\sum_{k=1}^{K}\lambda_{k}\biggr(\int_{\Omega}\int_{\Omega}c(z,z^{\prime})d\gamma(z,z^{\prime})+D_{\xi}(\gamma(\cdot|z)||Q_{k})\biggr),$
 is equivalent to the following max-min problem,
-$\underset{\{g_{1},\cdots,g_{K}:\sum_{k}\lambda_{k}g_{k}=0\}}{\text{max}}\,\underset{\{\gamma_{1},\cdots,\gamma_{K}:\gamma_{k}(\cdot|z)\in\Gamma(P_{k})\}}{\text{min}}\,\biggr\{\sum_{k=1}^{K}\lambda_{k}\biggr(\mathbb{E}_{z\sim Q_{k}}\biggr[-\bar{\xi}_{k}\biggr(-\mathbb{E}_{z^{\prime}\sim\gamma_{k}(\cdot|z)}[c_{k}(z,z^{\prime})-g_{k}(z^{\prime})]\biggr)\biggr]\biggr)\biggr\}.$
+$\underset{\{g_{1:K}:\sum_{k}\lambda_{k}g_{k}=0\}}{\text{max}}\,\underset{\{\gamma_{1},\cdots,\gamma_{K}:\gamma_{k}(\cdot|z)\in\Gamma(P_{k})\}}{\text{min}}\,\biggr\{\sum_{k=1}^{K}\lambda_{k}\biggr(\mathbb{E}_{z\sim Q_{k}}\biggr[-\bar{\xi}_{k}\biggr(-\mathbb{E}_{z^{\prime}\sim\gamma_{k}(\cdot|z)}[c_{k}(z,z^{\prime})-g_{k}(z^{\prime})]\biggr)\biggr]\biggr)\biggr\}.$
 Normalizing Flows.
-This approach was proposed by [visentin2025computing], and relies on the analysis of [gangbo1996geometry].
+This approach was proposed by visentin2025computing, and relies on the analysis of gangbo1996geometry.
 Their main idea is re-writing the Wasserstein distance in terms of functions $f,g$,
-$\mathbb{W}_{p}(P,Q)^{p}=\underset{f\in\text{Borel}(P_{0},P),g\in\text{Borel}(P_{0},Q)}{\text{inf}}\,\lVert f-g\rVert_{L_{p}(P_{0})}^{p},$ (34)
+$\mathbb{W}_{2}(P,Q)^{2}=\underset{f\in\text{Borel}(P_{0},P),g\in\text{Borel}(P_{0},Q)}{\text{inf}}\,\lVert f-g\rVert_{L_{2}(P_{0})}^{2},$ (35)
 where $\text{Borel}(P,Q)$ is the set of Borel functions such that $f_{\sharp}P=Q$.
@@
 Furthermore, $P_{0}$ is an arbitrary latent measure, which further links the approach with generative modeling (i.e., take $P_{0}=\mathcal{N}(0,\text{I})$).
-Based on equation 34, the insight of [visentin2025computing] is modeling $f$ and $g$ through a conditional function $f(z,s)$, $s\in\{0,1\}$.
+Based on equation 35, the insight of visentin2025computing is modeling $f$ and $g$ through a conditional function $f(z,s)$, $s\in\{0,1\}$.
 The Wasserstein barycenter problem is a straightforward generalization of the $2$ measures case.
 For instance, the authors consider $S=\{1,\cdots,K\}$, which means that,
-$\mathbb{B}_{\mathcal{Q}}(P)=\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{p}(P,Q_{k})^{p}=\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{p}(g_{\sharp}P_{0},Q_{k})=\sum_{k=1}^{K}\lambda_{k}\lVert g-f(z,k)\rVert_{L_{p}(P_{0})}^{p},$
+$\mathbb{B}(P|\mathcal{Q})=\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2}(P,Q_{k})^{2}=\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2}(g_{\sharp}P_{0},Q_{k})^{2}=\sum_{k=1}^{K}\lambda_{k}\lVert g-f(z,k)\rVert_{L_{p}(P_{0})}^{2},$
 where $g$ is the push-forward of $P_{0}$ to the barycenter $P^{\star}$, and $f(z,k)$ is the push-forward of $Q_{k}$ to $P^{\star}$. As the authors show in [visentin2025computing, Theorem 3.3],
 $g(z)=\sum_{k=1}^{K}\lambda_{k}f(z,k).$
-The main challenge in [visentin2025computing] is formulating a conditional model $f:\Omega\times S$.
+The main challenge in visentin2025computing is formulating a conditional model $f:\Omega\times S$.
 As the authors discuss, one possibility would be using $K=|S|$ different neural nets for modeling $f(z,\cdot)$, which is similar to what is done, for instance, in the neural net methods.
-The authors take a different route and use conditional normalizing flows [kobyzev2020normalizing] to model these functions.
+The authors take a different route and use conditional normalizing flows kobyzev2020normalizing to model these functions.
 B.4 Future Directions
 From the previous discussion on neural net-based methods for computing Wasserstein barycenters, we can identify a few common design choices.
-For instance, [korotin2021continuous, kolesov2024estimating] and [gazdieva2024robust] use $\mathcal{O}(K)$ neural nets for computing Wasserstein barycenters.
+For instance, korotin2021continuous, kolesov2024estimating and gazdieva2024robust use $\mathcal{O}(K)$ neural nets for computing Wasserstein barycenters.
 This design choice results in a parametrization of the underlying barycenter that grows with the number of input measures.
-The only neural net-based method that differs from this idea is [visentin2025computing], which uses a conditional model $f:\Omega\times S$.
+The only neural net-based method that differs from this idea is visentin2025computing, which uses a conditional model $f:\Omega\times S$.
 Meanwhile, as we cover in our experiments (C.f., Section 4 in the main paper), there is a clear advantage of incorporating labels in the ground-cost when these are available.
@@
 We leave the theoretical results for general metrics and $p$ to future work.
-Notation.
-For the following theoretical analysis, we use the following notation.
-Given a measurable map $f,g:\Omega\rightarrow\Omega$.
-Given a measure $P\in\mathcal{P}_{2,\text{ac}}(\Omega)$, we denote,
-$\lVert f\rVert_{L_{2}(P)}=\int_{\Omega}\lVert f(x)\rVert_{2}^{2}dP(x).$
+Notation. For the following theoretical analysis, we use the following notation. Given a measurable map $f,g:\Omega\rightarrow\Omega$. Given a measure $P\in\mathcal{P}_{2}(\Omega)$, we denote,
+$\lVert f\rVert_{L_{2}(P)}^{2}=\int_{\Omega}\lVert f(x)\rVert_{2}^{2}dP(x).$
 Likewise, we have a notion of inner product,
@@
 Definition C.1.
-Given a functional $\mathbb{F}:\mathcal{P}_{2,\text{ac}}(\Omega)(\mathbb{R}^{d})\rightarrow\mathbb{R}$, we call $\delta\mathbb{F}(P):\mathbb{R}^{d}\rightarrow\mathbb{R}$, if it exists, the unique (up to additive constants) function such that
-$\dfrac{d}{d\varepsilon}\mathbb{F}(P+\varepsilon\chi)\biggr|_{\epsilon=0}=\int\delta\mathbb{F}(P)d\chi$
-for every perturbation $\chi$ such that, at least for $\varepsilon\in[0,\varepsilon_{0}]$, the measure $P+\varepsilon\chi\in\mathcal{P}_{p,\text{ac}}(\mathbb{R}^{d})$.
+Given a functional $\mathbb{F}:\mathcal{P}_{2}(\Omega)(\mathbb{R}^{d})\rightarrow\mathbb{R}$, we call $\delta\mathbb{F}(P):\mathbb{R}^{d}\rightarrow\mathbb{R}$, if it exists, the unique (up to additive constants) function such that
+$\dfrac{d}{d\varepsilon}\mathbb{F}(P+\varepsilon\chi)\biggr|_{\epsilon=0}=\int\delta\mathbb{F}(P)d\chi,$ (36)
+for every perturbation $\chi$ such that, at least for $\varepsilon\in[0,\varepsilon_{0}]$, the measure $P+\varepsilon\chi\in\mathcal{P}_{2}(\mathbb{R}^{d})$.
 This functional is called the first variation of $\mathbb{F}$ at $P$.
-Following [chewi2024statistical], given a curve of densities $\{P_{t}\}_{t\geq 0}$, we can take $\chi=\partial_{t}\mathbb{F}(P_{t})$ so that $\partial_{t}\mathbb{F}(P_{t})=\int\delta\mathbb{F}(P_{t})\partial_{t}P_{t}$.
-From Riemannian geometry, we can define the Wasserstein gradient to $P_{t}$ as,
-$\partial_{t}\mathbb{F}(P_{t})=\langle\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t}),v_{t}\rangle_{P_{t}},$ (35)
-where $v_{t}\in\mathcal{T}_{P_{t}}\mathcal{P}_{p,\text{ac}}(\mathbb{R}^{d})$.
-Note, here, that $\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t}):\Omega\rightarrow\Omega$ is a vector field.
-In comparison, the first variation $\delta\mathbb{F}$ is a scalar field.
-These two objects are connected by the following equation [chewi2024statistical, Proposition 5.10],
-$\nabla\mkern-10.0mu\nabla\mathbb{F}(P)=\nabla\delta\mathbb{F}(P),$ (36)
-Meanwhile, recall that, from dynamic OT, the WGF satisfy the continuity equation,
-$\partial_{t}P_{t}+\text{div}(P_{t}v_{t})=0,$
-where $\text{div}(v)=\nabla\cdot v$ is the divergence of a vector field.
-Therefore, taking the direction of steepest descent of the functional $\mathbb{F}$, $v_{t}=-\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t})$, leads to the following connection between $\partial_{t}\mathbb{F}(P_{t})$ and $\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t})$,
-$\partial_{t}\mathbb{F}(P_{t})=-\lVert\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t})\rVert_{L_{2}(P_{t})}^{2},$ (37)
+We refer readers to [chewi2024statistical, Chapter 5] and [ambrosio2008gradient, Chapter 8] for further details on gradient flows in Wasserstein spaces.
+The main tool we will use here is the continuity equation,
+$\partial_{t}P_{t}=-\text{div}(P_{t}v_{t}),$
+where the divergence operator should be understood in the measure theoretic sense.
+Here, $v_{t}:\Omega\rightarrow\Omega$ is a vector field.
+Now, given $\mathbb{F}:\mathcal{P}_{2}(\Omega)\to\mathbb{R}$ with first variation $\delta\mathbb{F}(P):\Omega\to\mathbb{R}$, the time derivative of $\mathbb{F}$ along the curve $\{P_{t}\}_{t\geq 0}$ can be computed as,
+$\partial_{t}\mathbb{F}(P_{t})\overset{(1)}{=}\int\delta\mathbb{F}(P_{t})d(\partial_{t}P_{t})\overset{(2)}{=}-\int\delta\mathbb{F}(P_{t})\text{div}(P_{t}v_{t})dz\overset{(3)}{=}\int\langle\nabla\delta\mathbb{F}(P_{t}),v_{t}\rangle dP_{t}\overset{(4)}{=}\langle\nabla\delta\mathbb{F}(P_{t}),v_{t}\rangle_{P_{t}},$
+where (1) is obtained by taking $\chi=\partial_{t}P_{t}$ in equation 36; (2) is obtained by applying the continuity equation; (3) is obtained by using integration by parts; finally, (4) is obtained via the definition of inner product.
+This motivates the equality,
+$\nabla_{W}\mathbb{F}(P)=\nabla\delta\mathbb{F}(P).$ (37)
+We refer readers to [chewi2024statistical, Proposition 5.10], for a more formal derivation.
+This discussion characterizes the Wasserstein gradient $\nabla_{W}\mathbb{F}(P):\Omega\to\Omega$ as a vector field.
+Therefore, taking the direction of steepest descent of the functional $\mathbb{F}$, $v_{t}=-\nabla_{W}\mathbb{F}(P_{t})$, leads to the following connection between $\partial_{t}\mathbb{F}(P_{t})$ and $\nabla_{W}\mathbb{F}(P_{t})$,
+$\partial_{t}\mathbb{F}(P_{t})=-\lVert\nabla_{W}\mathbb{F}(P_{t})\rVert_{L_{2}(P_{t})}^{2},$ (38)
 C.2 Wasserstein Gradient of Functionals
@@
 which means that $V=\delta\mathbb{V}(P)$. As a consequence, the Wasserstein gradient of $\mathbb{V}$ is,
-$\nabla\mkern-10.0mu\nabla\mathbb{V}=\nabla V.$
+$\nabla_{W}\mathbb{V}=\nabla V.$
 which is a vector field over $\Omega$.
@@
 $\dfrac{d}{d\epsilon}\mathbb{G}(P+\epsilon\chi)\biggr|_{\epsilon=0}=\int G^{\prime}(P(z))\chi(z)dz=\int G^{\prime}(P(z))d\chi(z),$
-which means that $\delta\mathbb{G}=G^{\prime}(P)$ and $\nabla\mkern-10.0mu\nabla\mathbb{G}=\nabla G^{\prime}(P(z))$.
+which means that $\delta\mathbb{G}=G^{\prime}(P)$ and $\nabla_{W}\mathbb{G}=\nabla G^{\prime}(P(z))$.
 Interaction Energy, $\mathbb{U}$. Here, let us assume $U(z,z^{\prime})$ symmetric for simplicity,
@@
 When $U(z,z^{\prime})=h(z-z^{\prime})$ for an even $h$, $\int U(z,z^{\prime})dP(z)=\int h(z-z^{\prime})P(z)dz=h\star P$, i.e., the convolution of $h$ and $P$.
-Furthermore, $\nabla\mkern-10.0mu\nabla\mathbb{U}=(\nabla h)\star P$
+Furthermore, $\nabla_{W}\mathbb{U}=(\nabla h)\star P$
 Barycenter Functional, $\mathbb{B}$.
@@
 In this case,
-$\nabla\mkern-10.0mu\nabla\mathbb{W}_{2}(P,Q)^{2}=2(\text{Id}-T_{P\rightarrow Q}^{\star})$
+$\nabla_{W}\mathbb{W}_{2}(P,Q)^{2}=2(\text{Id}-T_{P\rightarrow Q}^{\star}).$ (39)
 Therefore, given $\mathcal{Q}=\{Q_{1},\cdots,Q_{K}\}$, it is straightforward to obtain:
-$\nabla\mkern-10.0mu\nabla\mathbb{B}_{\mathcal{Q}}(P)$ $=\nabla\mkern-10.0mu\nabla\biggr(\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2}(P,Q_{k})^{2}\biggr)=2\sum_{k=1}^{K}\lambda_{k}(\text{Id}-T_{P\rightarrow Q_{k}}^{\star}).$
+$\nabla_{W}\mathbb{B}(P|\mathcal{Q})$ $=\nabla_{W}\biggr(\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2}(P,Q_{k})^{2}\biggr)=2\sum_{k=1}^{K}\lambda_{k}(\text{Id}-T_{P\rightarrow Q_{k}}^{\star}).$ (40)
 A Connection with [cuturi2014fast, Algorithm 2].
@@
 Assume $\gamma^{\star}$ computed between $\hat{P}$ and $\hat{Q}$, which means that,
-$\mathbb{W}_{2}(\hat{P},\hat{Q})^{2}=\sum_{i=1}^{n}\sum_{j=1}^{m}\gamma_{ij}^{\star}\lVert z_{i}^{(P)}-z_{j}^{(Q)}\rVert_{2}^{2}\text{, and }\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})=\sum_{k=1}^{K}\lambda_{k}\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}^{\star}\lVert z_{\tau,i}^{(P)}-z_{j}^{(Q_{k})}\rVert_{2}^{2}.$
-We can compute the Wasserstein gradient $\nabla\mkern-10.0mu\nabla\mathbb{B}_{\mathcal{Q}}$ by computing $\nabla_{z_{\tau,i}^{(P)}}\mathbb{B}_{\mathcal{Q}}$, that is,
-$\nabla_{z_{\tau,i}^{(P)}}\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})=2\sum_{k=1}^{K}\lambda_{k}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}^{\star}(z_{\tau,i}^{(P)}-z_{j}^{(Q_{k})})=\dfrac{2}{n}z_{\tau,i}^{(P)}-2\sum_{k=1}^{K}\lambda_{k}\sum_{j=1}^{K}\gamma_{k,i,j}^{\star}z_{j}^{(Q_{k})}.$
+$\mathbb{W}_{2}(\hat{P},\hat{Q})^{2}=\sum_{i=1}^{n}\sum_{j=1}^{m}\gamma_{ij}^{\star}\lVert z_{i}^{(P)}-z_{j}^{(Q)}\rVert_{2}^{2}\text{, and }\mathbb{B}(\hat{P}_{\tau}|\mathcal{Q})=\sum_{k=1}^{K}\lambda_{k}\sum_{i=1}^{n}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}^{\star}\lVert z_{\tau,i}^{(P)}-z_{j}^{(Q_{k})}\rVert_{2}^{2}.$
+We can compute the Wasserstein gradient $\nabla_{W}\mathbb{B}$ by computing $\nabla_{z_{\tau,i}^{(P)}}\mathbb{B}$, that is,
+$\nabla_{z_{\tau,i}^{(P)}}\mathbb{B}(\hat{P}_{\tau}|\mathcal{Q})=2\sum_{k=1}^{K}\lambda_{k}\sum_{j=1}^{n_{k}}\gamma_{k,i,j}^{\star}(z_{\tau,i}^{(P)}-z_{j}^{(Q_{k})})=\dfrac{2}{n}z_{\tau,i}^{(P)}-2\sum_{k=1}^{K}\lambda_{k}\sum_{j=1}^{K}\gamma_{k,i,j}^{\star}z_{j}^{(Q_{k})}.$
 From this last equation, we can use the barycentric mapping definition,
@@
 to write,
-$\nabla_{z_{\tau,i}^{(P)}}\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})=\dfrac{2}{n}\biggr(z_{\tau,i}^{(P)}-\sum_{k=1}^{K}\lambda_{k}T_{P\rightarrow Q_{k}}^{\star}(z_{\tau,i}^{(P)})\biggr),$
+$\nabla_{z_{\tau,i}^{(P)}}\mathbb{B}(\hat{P}_{\tau}|\mathcal{Q})=\dfrac{2}{n}\biggr(z_{\tau,i}^{(P)}-\sum_{k=1}^{K}\lambda_{k}T_{P\rightarrow Q_{k}}^{\star}(z_{\tau,i}^{(P)})\biggr),$
 which means we can write the gradient flow update rule as,
@@
 Which is the update rule in [cuturi2014fast, Algorithm 2] with $\theta=\nicefrac{{2\alpha}}{{n}}$.
-Furthermore, we can also instantiate the fixed-point algorithm of [montesuma2023multi], for $\alpha=n/2$.
+Furthermore, we can also instantiate the fixed-point algorithm of montesuma2023multi, for $\alpha=n/2$.
 We now highlight two differences of our work with respect these two strategies.
@@
 Furthermore, our framework couples the barycenter objective with other functionals.
-Second, while [cuturi2014fast] assume that all samples of $\mathcal{Q}=\{Q_{1},\cdots,Q_{K}\}$ are available at once, our algorithm assumes that these measures are available through sampling.
-This treatment is oftentimes called continuous in the literature [korotin2021continuous, korotin2022wasserstein], however, one still relies on samples from these measures rather than some truly continuous form.
-C.3 Domain Adaptation
-In this section, we offer a brief introduction to domain adaptation in classification problems which are the main application we explore in the main paper.
-Statistical learning theory [vapnik1999overview] deals with the problem of learning a function $h:\mathcal{X}\rightarrow\mathcal{Y}$, such that it minimizes the risk of making a mistake, that is,
+Second, while cuturi2014fast assume that all samples of $\mathcal{Q}=\{Q_{1},\cdots,Q_{K}\}$ are available at once, our algorithm assumes that these measures are available through sampling.
+This treatment is oftentimes called continuous in the literature korotin2021continuous, korotin2022wasserstein, however, one still relies on samples from these measures rather than some truly continuous form.
+Appendix D Proof of Theorem 3.1
+In this section, we present the proof for theorem 3.1. We mainly use two tools to obtain this result,
+  1. 1.
+    The exponential convergence given by the PL inequality,
+    $\lVert\nabla_{W}\mathbb{B}(P)\rVert_{L_{2}(P)}^{2}\geq C_{\text{PL}}(\mathbb{B}(P)-\mathbb{B}^{\star}),$ (42)
+  2. 2.
+    An uniform bound on $P\mapsto\mathbb{B}(P|\mathcal{Q})$ controlling the error of approximating $Q_{k}$ with $\hat{Q}_{k}$.
+These results are available, respectively, in chewi2024statistical and portales2025sample.
+For completeness, we re-state and demonstrate them.
+We start with a modified version of [chewi2024statistical, Corollary 5.17],
+Proposition D.1.
+(Measure theoretic Gronwall’s Lemma chewi2024statistical) Let $\mathbb{F}:\mathcal{P}_{2}(\Omega)\rightarrow\mathbb{R}$ be a functional that satisfies the PL inequality (c.f.
+42) with constant $C_{\text{PL}}>0$, then,
+$\mathbb{F}(P_{\tau})-\mathbb{F}^{\star}\leq e^{-2C_{\text{PL}}\tau}(\mathbb{F}(P_{0})-\mathbb{F}^{\star}).$ (43)
+Proof.
+Define $\phi(t)=\mathbb{F}(P_{t})-\mathbb{F}^{\star}$. Then,
+$\partial_{t}\phi(t)$ $=\partial_{t}\mathbb{F}(P_{t})=-\lVert\nabla_{W}\mathbb{F}(P_{t})\rVert_{L_{2}(P_{t})}^{2},$
+$\geq-C_{\text{PL}}(\mathbb{F}(P_{0})-\mathbb{F}^{\star})=-C_{\text{PL}}\phi(t).$
+From Grönwall’s Lemma [chewi2024statistical, Lemma 5.16],
+$\underbrace{\mathbb{F}(P_{t})-\mathbb{F}^{\star}}_{\partial_{t}\phi(t)}\leq e^{-C_{\text{PL}}}\underbrace{(\mathbb{F}(P_{0})-\mathbb{F}(P)^{\star})}_{\phi(0)}$
+We now present the uniform approximation bound of portales2025sample for empirical barycenters,
+Theorem D.1.
+(Barycenter Empirical Approximation error portales2025sample) Let $\mathcal{Q}=\{Q_{1},\cdots,Q_{K}\}$ be a finite family of measures on $\mathcal{P}_{2}(\mathcal{B}(0,R))$ for $R>0$.
+Approximate each $Q_{k}$ with $m$ i.i.d.
+samples $\{z_{i}^{(Q_{k})}\}_{i=1}^{m}$.
+Let $\hat{P}=n^{-1}\sum_{i=1}^{n}\delta_{z_{i}^{(P)}}$.
+Let $p\in[1,+\infty)$ and $\epsilon\geq 0$.
+Then,
+$\mathbb{E}\biggr[\underset{z_{1}^{(P)},\cdots,z_{n}^{(P)}\in\Omega}{\text{sup}}\,|\mathbb{B}_{\epsilon}(\hat{P}|\mathcal{Q})-\widehat{\mathbb{B}}_{\epsilon}(\hat{P}|\mathcal{Q})|\biggr]\leq C_{p,R}\sqrt{\dfrac{C_{d,n}}{m}},$ (44)
+where,
+$C_{p,R}$ $=\begin{cases}8\sqrt{2}\int_{0}^{3(2R)^{p}}\sqrt{\log\biggr(2+\dfrac{32p(2R)^{p}}{s}\biggr)}ds&\epsilon=0,\\ 8\sqrt{2}\int_{0}^{(4p+1)(2R)^{p}}\sqrt{\log(2+\frac{64p(2R)^{p}}{s})}ds&\epsilon>0,\end{cases},$
+$C_{d,n}$ $=n(d+1).$
+With these bounds established, we can then prove our Theorem 3.1.
+Proof.
+Let $\tau>0$ be the current optimization step. We have the following,
+[…]
+[…]
+From Equation 45 we can bound the expected empirical optimization gap,
+$\mathbb{E}[|\hat{\mathbb{B}}(\hat{P}_{\tau})-\hat{\mathbb{B}}^{\star}|]\leq$ $\mathbb{E}[|\hat{\mathbb{B}}(\hat{P}_{\tau})-\mathbb{B}(\hat{P}_{\tau})|]+\mathbb{E}[|\mathbb{B}(\hat{P}_{\tau})-\mathbb{B}^{\star}|]+\mathbb{E}[|\mathbb{B}^{\star}-\hat{\mathbb{B}}^{\star}|].$ (46)
+Now, we can bound the first and third terms using the approximation error of portales2025sample (i.e., Equation 44),
+$\mathbb{E}[|\hat{\mathbb{B}}(\hat{P}_{\tau})-\mathbb{B}(\hat{P}_{\tau})|]$ $\leq C_{p,R}\sqrt{\dfrac{C_{d,n}}{m}},$ (47)
+$\mathbb{E}[|\hat{\mathbb{B}}^{\star}-\mathbb{B}^{\star}|]$ $\leq C_{p,R}\sqrt{\dfrac{C_{d,n}}{m}},$
+here, we recall that $\mathbb{B}^{\star}=\mathbb{B}(P^{\star})=\text{inf}_{P}\mathbb{B}(P)$.
+Returning to the second term, note that the true optimization gap $\mathbb{B}(\hat{P}_{\tau}|\mathcal{Q})-\mathbb{B}^{\star}$ is deterministic, as it does not involve sampling from the measures in $\mathcal{Q}$.
+This means that,
+$\mathbb{E}[|\mathbb{B}(\hat{P}_{\tau})-\mathbb{B}^{\star}|]=|\mathbb{B}(\hat{P}_{\tau})-\mathbb{B}^{\star}|\leq e^{-C_{\text{PL}}\tau}(\hat{\mathbb{B}}(\hat{P}_{0})-\hat{\mathbb{B}}^{\star}).$ (48)
+Combining Equations 46, 47 and 48 we get,
+$\mathbb{E}[|\hat{\mathbb{B}}(\hat{P}_{\tau})-\hat{\mathbb{B}}^{\star}|]\leq e^{-C_{\text{PL}}\tau}(\mathbb{B}(\hat{P}_{0})-\mathbb{B}^{\star})+C_{p,R}\sqrt{\dfrac{C_{d,n}}{m}},$
+which is the desired inequality. ∎
+D.1 Empirical Verification of the PL Inequality
+An open question in the literature is the verification of the PL inequality (Equation 42) for general measures, such as empirical ones.
+To give further context, current results prove that it indeed holds for Gaussian measures, under specific spectral conditions chewi2020gradient.
+Therefore, we do an empirical validation on whether this bound holds in the empirical case.
+In a nutshell, to test Equation 42, we need to compute the PL ratio,
+$\text{PL}_{\text{ratio}}(\hat{P}_{\tau})=\dfrac{\lVert\nabla_{W}\mathbb{B}(\hat{P}_{\tau})\rVert_{L_{2}(\hat{P}_{\tau})}}{\mathbb{B}(\hat{P}_{\tau})-\mathbb{B}^{\star}},$ (49)
+and verify that it indeed remains strictly positive over the course of our optimization.
+Computing the PL ratio in Equation 49 requires knowledge of $\mathbb{B}^{\star}$, which is, in general, intractable.
+However, for location-scatter families, it is known in closed form, since $P^{\star}$ itself is known.
+Therefore, we verify it in the context of the Swiss-roll measure, where all quantities are known.
+We show our results in Figure 7, which shows that the ratio converges with iterations to approximately $0.56$.
+Figure 7: Empirical test of the PL inequality. From left to right, we show: (i) the optimality gap, (ii) the gradient norm, and (iii) the PL ratio as specified in Equation 49.
+D.2 PL Inequality on Location-Scatter Family
+An open question in the literature is the verification of the PL inequality (Equation 42) for general measures.
+Current results establish that it holds for Gaussian measures under specific spectral conditions [chewi2020gradient].
+In this section, we show that the PL inequality transfers from Gaussians to the broader class of location-scatter families, which includes our Swiss roll experiments (Section 4).
+Location-Scatter Families.
+Following alvarez2016fixed, let $Q_{0}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ be a measure with mean $\mu_{0}$ and covariance $\Sigma_{0}$.
+The location-scatter family generated by $Q_{0}$ is,
+$\mathcal{L}(Q_{0}):=\left\{T_{\sharp}Q_{0}:T(x)=Ax+m,b\in\mathbb{R}^{d},A\in\mathcal{S}^{++}_{d}\right\}$, (50)
+where $\mathcal{S}^{++}_{d}$ denotes the cone of $d\times d$ symmetric positive definite matrices.
+Each $Q\in\mathcal{L}(Q_{0})$ has mean $b$ and covariance $A$, but inherits the higher-order structure of $Q_{0}$.
+Isometry to the Bures-Wasserstein Manifold. The key insight comes from alvarez2016fixed: for any $Q_{1},Q_{2}\in\mathcal{L}(Q_{0})$, the 2-Wasserstein distance equals
+$\mathbb{W}_{2}^{2}(Q_{1},Q_{2})=\|\mu_{1}-\mu_{2}\|^{2}+\text{Tr}(\Sigma_{1}+\Sigma_{2}-2(\Sigma_{2}^{1/2}\Sigma_{1}\Sigma_{2}^{1/2})^{1/2}),$ (51)
+This is precisely the Wasserstein distance on the Bures-Wasserstein manifold: $\mathcal{BW}=\{\mathcal{N}(\mu,\Sigma):\mu\in\mathbb{R}^{d},\Sigma\in\mathcal{S}_{d}^{++}\}$ [takatsu2011wasserstein].
+Henceforth, we note the elements of $\mathcal{BW}$ by $G_{\mu,\Sigma}$.
+In that sense $G_{\mu_{1},\Sigma_{1}}$ and $G_{\mu_{2},\Sigma_{2}}$ denote the Gaussian measures matching the moments of $Q_{1},Q_{2}\in\mathcal{L}(Q_{0})$.
+As a consequence, the restriction of $\mathbb{W}_{2}$ to $\mathcal{L}(Q_{0})$ is isometric to the Bures-Wasserstein manifold.
+Moreover, the optimal transport map between elements of $\mathcal{L}(Q_{0})$ takes the same form as for Gaussians:
+$T^{\star}_{Q_{1}\to Q_{2}}(x)=\mu_{2}+A(x-\mu_{1}),\quad A=\Sigma_{1}^{-1/2}\left(\Sigma_{1}^{1/2}\Sigma_{2}\Sigma_{1}^{1/2}\right)^{1/2}\Sigma_{1}^{-1/2}.$ (52)
+Barycenter Functional Equivalence.
+Let $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}\subset\mathcal{L}(Q_{0})$ with $Q_{k}=T_{k}Q_{0}$, such that $\mu_{k}$ and $\Sigma_{k}$ are the mean and covariance of $Q_{k}$.
+Let $\mathcal{G}=\{G_{\mu_{k},\Sigma_{k}}\}_{k=1}^{K}$ be the corresponding Gaussian measures.
+Equation 51 establishes:
+$\mathbb{B}(P|\mathcal{Q})=\mathbb{B}(G_{\mu,\Sigma}|\mathcal{G})\text{, where }P\in\mathcal{L}(Q_{0})\text{, and }G_{\mu,\Sigma}\in\mathcal{BW},$ (53)
+where $\mu,\Sigma$ are the mean and covariance of $P\in\mathcal{L}(Q_{0})$.
+Furthermore, by alvarez2016fixed, the barycenter $P^{\star}\in\mathcal{L}(Q_{0})$ corresponds to the Bures-Wasserstein barycenter: if $P^{\star}$ is the minimizer of $P\mapsto\mathbb{B}(P|\mathcal{Q})$ over $\mathcal{L}(Q_{0})$ with mean $\mu^{\star}$ and covariance $\Sigma^{\star}$, then $G_{\mu^{\star},\Sigma^{\star}}$ is the minimizer of $G_{\mu,\Sigma}\mapsto\mathbb{B}(G_{\mu,\Sigma}|\mathcal{G})$.
+Wasserstein Gradient Structure. The Wasserstein gradient of $\mathbb{B}$ at $P$ with mean and covariance $\mu$ and $\Sigma$ is (cf. Equation 40)
+$\nabla_{W}\mathbb{B}(P|\mathcal{Q})=2\sum_{k=1}^{K}\lambda_{k}\left(\mathrm{Id}-T^{\star}_{P\to Q_{k}}\right).$ (54)
+Since each $T^{\star}_{P\to Q_{k}}$ is affine (Equation 52), the gradient $\nabla_{W}\mathbb{B}(P|\mathcal{Q})$ is itself an affine map:
+$\nabla_{W}\mathbb{B}(P|\mathcal{Q})(x)$ $=2\sum_{k=1}^{K}\lambda_{k}(x-(A_{k}x+b_{k})),$
+$=2\biggr(x-(\sum_{k=1}^{K}\lambda_{k}A_{k})(x)+\sum_{k=1}^{K}\lambda_{k}b_{k}\biggr),$
+$=\underbrace{\biggr(2(\text{Id}-\sum_{k=1}^{K}\lambda_{k}A_{k})\biggr)}_{M}(x)+\underbrace{2\sum_{k=1}^{K}\lambda_{k}b_{k}}_{c},$
+$=Mx+c,$ (55)
+where $M=2(\text{Id}-\bar{A})$, $c=2(\bar{A}\mu-\bar{\mu})$, where $\bar{A}=\sum_{k=1}^{K}\lambda_{k}A_{k}$ and $\bar{\mu}=\sum_{k=1}^{K}\lambda_{k}\mu_{k}$.
+More importantly, by Equation (52), this is the same affine map as for the Gaussian case, resulting in,
+$\nabla_{W}\mathbb{B}(P|\mathcal{Q})=\nabla_{W}\mathbb{B}(G_{\mu,\Sigma}|\mathcal{G})\quad\text{as vector fields }\mathbb{R}^{d}\to\mathbb{R}^{d}.$ (56)
+We now establish that the $L^{2}$-norms of these gradients are also equal, which is the key step for transferring the PL inequality.
+Lemma D.1 ($L^{2}$-norm of Affine Maps).
+Let $T(x)=Mx+c$ be an affine map with $M\in\mathbb{R}^{d\times d}$ and $c\in\mathbb{R}^{d}$.
+For any probability measure $P\in\mathcal{P}_{2}(\mathbb{R}^{d})$ with mean $\mu_{P}$ and covariance $\Sigma_{P}$,
+$\|T\|_{L^{2}(P)}^{2}:=\int_{\mathbb{R}^{d}}\|T(x)\|^{2}\,dP(x)=\text{Tr}(M^{T}M\Sigma_{P})+\|M\mu_{P}+c\|^{2}.$ (57)
+In particular, $\|T\|_{L^{2}(P)}^{2}$ depends on $P$ only through its first two moments.
+Proof.
+Expanding the squared norm:
+$\|T\|_{L^{2}(P)}^{2}$ $=\int\|Mx+c\|^{2}\,dP(x)$
+$=\int\|M(x-\mu_{P})+(M\mu_{P}+c)\|^{2}\,dP(x)$
+$=\int\|M(x-\mu_{P})\|^{2}\,dP(x)+2\langle M\mu_{P}+c,M\underbrace{\int(x-\mu_{P})\,dP(x)}_{=0}\rangle+\|M\mu_{P}+c\|^{2}$
+$=\int(x-\mu_{P})^{T}M^{T}M(x-\mu_{P})\,dP(x)+\|M\mu_{P}+c\|^{2}$
+$=\text{Tr}(M^{T}M\Sigma_{P})+\|M\mu_{P}+c\|^{2},$
+where the last step uses $\mathbb{E}_{x\sim P}[(x-\mu_{P})(x-\mu_{P})^{\top}]=\Sigma_{P}$. ∎
+Lemma D.2 (Gradient Norm Equality).
+For any $P\in\mathcal{L}(Q_{0})$ and its corresponding Gaussian $G_{\mu,\Sigma}$:
+$\|\nabla_{W}\mathbb{B}(P|\mathcal{Q})\|_{L^{2}(P)}^{2}=\|\nabla_{W}\mathbb{B}(G|\mathcal{G})\|_{L^{2}(G_{\mu,\Sigma})}^{2}.$ (58)
+Proof.
+First, define $T_{\text{LS}}=\nabla_{W}\mathbb{B}(P|\mathcal{Q})$ and $T_{\text{Gaussian}}=\nabla_{W}\mathbb{B}(G_{m,\Sigma}|\mathcal{G})$.
+By construction, both of these mappings are affine, and by Equation 56 $T_{\text{LS}}=T_{\text{Gaussian}}$.
+From Lemma D.2, since $P$ and $G_{\mu,\Sigma}$ share the first and second order moments by construction, they are equal, which means,
+$\lVert\nabla_{W}\mathbb{B}(P|\mathcal{Q})\rVert_{L_{2}(P)}^{2}=T_{\text{LS}}=T_{\text{Gaussian}}=\lVert\nabla_{W}\mathbb{B}(G|\mathcal{G})\rVert_{L_{2}(G_{\mu,\Sigma})}^{2},$
+which is the desired result. ∎
+We can now state and prove the following result.
+Proposition D.2 (PL Inequality Transfer).
+Let $\mathcal{L}(Q_{0})$ be a location-scatter family and $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}\subset\mathcal{L}(Q_{0})$.
+Let $\mathcal{G}=\{G_{\mu_{k},\Sigma_{k}}\}_{k=1}^{K}$ be the corresponding Gaussian with matching moments.
+If the Bures-Wasserstein barycenter functional $\mathbb{B}(\cdot|\mathcal{G})$ satisfies the PL inequality with constant $C_{\text{PL}}>0$, i.e.,
+$\|\nabla_{W}\mathbb{B}(G_{\mu,\Sigma}|\mathcal{G})\|_{L^{2}(G_{\mu,\Sigma})}^{2}\geq C_{\text{PL}}\left(\mathbb{B}(G_{\mu,\Sigma}|\mathcal{G})-\mathbb{B}(G_{\mu^{\star},\Sigma^{\star}}|\mathcal{G})\right),$ (59)
+then $\mathbb{B}(\cdot|\mathcal{Q})$ satisfies the PL inequality on $\mathcal{L}(Q_{0})$ with the same constant,
+$\|\nabla_{W}\mathbb{B}(P|\mathcal{Q})\|_{L^{2}(P_{m,\Sigma})}^{2}\geq C_{\mathrm{PL}}\left(\mathbb{B}(P|\mathcal{Q})-\mathbb{B}(P^{\star}|\mathcal{Q})\right).$ (60)
+Proof.
+Starting from the PL inequality for Gaussians (Equation 59):
+$\|\nabla_{W}\mathbb{B}(G_{\mu,\Sigma}|\mathcal{G})\|_{L^{2}(G_{\mu,\Sigma})}^{2}\geq C_{\text{PL}}\left(\mathbb{B}(G_{\mu,\Sigma}|\mathcal{G})-\mathbb{B}(G_{\mu^{\star},\Sigma^{\star}}|\mathcal{G})\right).$
+Applying Lemma D.2 to the left-hand side:
+$\|\nabla_{W}\mathbb{B}(P|\mathcal{Q})\|_{L^{2}(P)}^{2}\geq C_{\text{PL}}\left(\mathbb{B}(G_{\mu,\Sigma}|\mathcal{G})-\mathbb{B}(G_{\mu^{\star},\Sigma^{\star}}|\mathcal{G})\right).$
+Applying Equation (53) and barycenter correspondence to the right-hand side:
+$\|\nabla_{W}\mathbb{B}(P|\mathcal{Q})\|_{L^{2}(P)}^{2}\geq C_{\mathrm{PL}}\left(\mathbb{B}(P|\mathcal{Q})-\mathbb{B}(P^{\star}|\mathcal{Q})\right).\qed$
+Explicit PL Constant.
+By chewi2020gradient, the PL inequality holds for Gaussians under $\zeta$-regularity: if $\|\Sigma_{k}\|_{\mathrm{op}}\leq 1$ and $\det\Sigma_{k}\geq\zeta$ for all $k$, then $C_{\mathrm{PL}}=\zeta^{2}/4$.
+Then, Proposition D.2 extends this PL constant to location-scatter families satisfying the same spectral bounds.
+Scope and Limitations.
+Proposition D.2 applies to location-scatter families, which are generated by affine transformations of a base measure.
+This includes elliptical distributions (e.g., multivariate $t$-distributions, uniform distributions on ellipsoids) and, importantly, the Swiss roll measures used in our experiments.
+This result does not directly apply to arbitrary measures.
+Appendix E Domain Adaptation
+E.1 Introduction
+Statistical learning theory vapnik1999overview deals with the problem of learning a function $h:\mathcal{X}\rightarrow\mathcal{Y}$, such that it minimizes the risk of making a mistake, that is,
 $h^{\star}=\underset{h\in\mathcal{H}}{\text{arg inf}}\,\biggr\{\mathcal{R}_{Q}(h)=\int_{\mathcal{X}\times\mathcal{Y}}\mathcal{L}(y,h(x))dQ(x,y)\biggr\},$ (61)
@@
 Indeed, both $\mathcal{R}_{P}$ and $\hat{\mathcal{R}}_{P}$ evaluate the risk under a fixed probability measure $P$.
-However, machine learning models are oftentimes confronted with data following different, but related probability measures [quinonero2022dataset], a process known as distribution shift.
+However, machine learning models are oftentimes confronted with data following different, but related probability measures quinonero2022dataset, a process known as distribution shift.
 The main tool for tackling distribution shift is domain adaptation.
@@
 A natural extension to this problem is multi-source domain adaptation, where multiple source domains are available, i.e., $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}$, each with its own associated dataset $\{(x_{i}^{(Q_{k})},y_{i}^{(Q_{k})})\}_{i=1}^{n_{k}}$.
-Source-Only Training.
+E.2 Source-Only Training
 A straightforward baseline in domain adaptation is the source-only training.
@@
 This training produces an embedding space $\mathcal{U}$ that is discriminative of the classes in the domain adaptation problem.
-Following previous work [montesuma2021wasserstein, montesuma2023multi, montesuma2024distillation], we perform adaptation in this space.
-Wasserstein Barycenters and Domain Adaptation.
-The motivation for using Wasserstein barycenters in multi-source domain adaptation comes from a theoretical result of [redko2019optimal].
-Indeed, by [redko2019optimal, Theorem 2],
+Following previous work montesuma2021wasserstein, montesuma2023multi, montesuma2024distillation, we perform adaptation in this space.
+E.3 Wasserstein Barycenter Transport
+The motivation for using Wasserstein barycenters in multi-source domain adaptation comes from a theoretical result of redko2019optimal. Indeed, by [redko2019optimal, Theorem 2],
 $\mathcal{R}_{Q_{T}}(h)\leq\mathcal{R}_{Q_{S}}(h)+\mathbb{W}_{1}(\hat{Q}_{S},\hat{Q}_{T})+\mathcal{O}(n_{S}^{-1/2}+n_{T}^{-1/2})+\mathbb{R}(Q_{S},Q_{T}),$ (63)
@@
 $\mathbb{W}_{p}(\hat{Q}_{k},P)\leq\frac{1}{p}\mathbb{W}_{p}(\hat{Q}_{k},P)^{p}+\frac{p-1}{p}\text{, and }\mathbb{W}_{p}(P,\hat{Q}_{T})\leq\frac{1}{p}\mathbb{W}_{p}(P,\hat{Q}_{T})^{p}+\frac{p-1}{p}$
-Plugging this back into inequality 42,
+Plugging this back into inequality 64,
 $\mathcal{R}_{Q_{T}}(h)\leq\sum_{k=1}^{K}\lambda_{k}\mathcal{R}_{Q_{k}}(h)+$ $\dfrac{1}{p}\biggr(\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{p}(\hat{Q}_{k},P)^{p}+\mathbb{W}_{p}(P,\hat{Q}_{T})^{p}\biggr)+\dfrac{2(p-1)}{p}+\mathcal{O}(n^{-1/2})+\sum_{k=1}^{K}\lambda_{k}\mathbb{R}(Q_{k},Q_{T}).$
@@
 3:for $\tau\leftarrow 1$ to $n_{\text{iter}}$ do
-4:  Compute $P=\text{Bar}(\lambda,\mathcal{Q})$
-5:  Compute $\gamma^{\star}$ between barycenter and target using eq. 43
-6:  Transport $(x_{i}^{(P)},y_{i}^{(P)})$ Using eq. 44
+4:  Compute $P^{\star}=\text{Bar}(\lambda,\mathcal{Q})$
+5:  Compute $\gamma^{\star}$ between $P^{\star}$ and $Q_{T}$ (eq. 65)
+6:  Transport $(x_{i}^{(P)},y_{i}^{(P)})$ (eq. 66)
 7:end for
-Wasserstein Barycenter Transport $(p=2)$. [montesuma2021wasserstein] uses a two-step procedure for finding $P$. First, one solves,
+Wasserstein Barycenter Transport $(p=2)$. montesuma2021wasserstein uses a two-step procedure for finding $P$. First, one solves,
 $P^{\star}=\underset{P\in\mathcal{P}_{1,\text{ac}}(\Omega)}{\text{arg min}}\,\sum_{k=1}^{K}\lambda_{k}\mathbb{W}_{2}(Q_{k},P)^{2},$
@@
 $\gamma^{\star}$ $=\underset{\gamma\in\Gamma(P,\hat{Q}_{T})}{\text{arg min}}\,\sum_{i=1}^{n}\sum_{j=1}^{n_{T}}\gamma_{ij}\lVert x_{i}^{(P)}-x_{j}^{(Q_{T})}\rVert_{2}^{2},$ (65)
+and,
 $T_{P\rightarrow Q_{T}}(x_{i}^{(P)})$ $=n_{T}\sum_{j=1}^{n_{T}}\gamma_{ij}^{\star}x_{j}^{(Q_{T})},$ (66)
@@
 Note that since this procedure generates a labeled dataset on data following the target domain measure, we can train a classifier with the pairs $\{T_{P\rightarrow Q_{T}}(x_{i}^{(P)}),y_{i}^{(P)}\}$.
-Appendix D Missing Proofs and Derivations
-D.1 Connecting equations 12 and 16
-In this section, we the empirical flow in equation 16 from the continuity equation. As a reminder, the continuity equation is,
-$\partial_{t}P_{t}=\text{div}(P_{t}\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t})).$
-Here we want to plug $\hat{P}_{t}=n^{-1}\sum_{i=1}^{n}\delta_{z_{t,i}}$ into this equation. On the one hand, plugging the formula for the empirical measure into the r.h.s. of this equation gives us,
-$\text{div}(\hat{P}_{t}\nabla\mkern-10.0mu\nabla\mathbb{F}(\hat{P}_{t}))=\text{div}\biggr(\dfrac{1}{n}\sum_{i=1}^{n}\delta_{z_{t,i}}\nabla(\delta\mathbb{F}(\hat{P}_{t}))(z_{t,i})\biggr).$ (45)
-Here, we used the definition for the Wasserstein gradient (c.f., equation 36), i.e., $\nabla\mkern-10.0mu\nabla\mathbb{F}(P)=\nabla(\delta\mathbb{F}(P))(z_{t,i})$.
-On the other hand, due to the linearity of derivatives and the chain rule,
-$\partial_{t}\hat{P}_{t}=\dfrac{1}{n}\sum_{i=1}^{n}\partial_{t}\delta_{z_{t,i}^{(P)}}=-\dfrac{1}{n}\sum_{i=1}^{n}\text{div}(\partial_{t}z_{t,i}\delta_{z_{t,i}})=\text{div}\biggr(-\dfrac{1}{n}\sum_{i=1}^{n}\delta_{z_{t,i}}\partial_{t}z_{t,i}\biggr).$ (46)
-Equating the r.h.s.
-of equations 45 and 46 gives us the continuous-time update $\partial_{t}z_{t,i}=-\nabla\mkern-10.0mu\nabla\mathbb{F}(\hat{P}_{t})$.
-We can discretize this differential equation in time, using the Forward Euler scheme,
-$z_{\tau+1,i}=z_{\tau,i}-\alpha\nabla\mkern-10.0mu\nabla\mathbb{F}(\hat{P}_{\tau}),$
-where $\alpha>0$ is the time-step and $\tau$ is the discrete time index.
-D.2 Proposition 3.1
-As a recap, in this section we provide a formal proof that when,
-$P=\sum_{i}\pi_{i}^{(P)}(\mathcal{N}(\cdot|\mu_{i}^{(P)},\Sigma_{i}^{(P)})\otimes\delta_{\nu_{i}^{(P)}}),$
-and under the ground cost,
-$c(z,z^{\prime})=\lVert x-x^{\prime}\rVert_{2}^{2}+\rho(y,y^{\prime}),$
-the $\mathbb{MW}_{2}$ distance becomes,
-$\mathbb{MW}_{2}(P,Q)^{2}=\sum_{i=1}^{n}\sum_{j=1}^{m}\omega^{\star}_{ij}(\mathbb{W}_{2}(P_{i},Q_{j})^{2}+\rho(\nu_{i}^{(P)},\nu_{j}^{(Q)})^{2}).$
-As a reminder, we have the following proposition,
-Proposition D.1.
-Let $P=\sum\pi_{i}^{(P)}(\mathcal{N}(\mu_{i}^{(P)},\Sigma_{i}^{(P)})\otimes\delta_{\nu_{i}^{(P)}})$ and $Q=\sum\pi_{j}^{(Q)}(\mathcal{N}(\mu_{j}^{(Q)},\Sigma_{j}^{(Q)}\otimes\delta_{\nu_{j}^{(Q)}})$ be 2 GMMs over $\Omega=\mathcal{X}\times\mathcal{Y}$.
-Let the ground cost $c$ be,
-$c(z,z^{\prime})=\lVert x-x^{\prime}\rVert_{2}^{2}+\rho(y,y^{\prime})^{2},$
-where $\rho(y,y^{\prime})$ is a metric over $\mathcal{Y}$. Then,
-$\mathbb{MW}_{2}(P,Q)^{2}=\underset{\omega\in\Gamma(\pi^{(P)},\pi^{(Q)})}{\text{min}}\,\sum_{i=1}^{n}\sum_{j=1}^{m}\omega_{ij}C_{ij},$
-where $C_{ij}=\mathbb{W}_{2}(P_{i,x},Q_{j,x})^{2}+\rho(\nu_{i}^{(P)},\nu_{j}^{(P)})^{2}$.
-Proof.
-Our first step is decomposing the Wasserstein distance $\mathbb{W}_{2}(P_{i},Q_{j})^{2}$ into $\mathbb{W}_{2}(P_{i,x},Q_{j,x})^{2}+\rho(\nu_{i}^{(P)},\nu_{j}^{(Q)})^{2}$. Indeed,
-$\mathbb{W}_{2}(P_{i},Q_{j})^{2}=\int\int c(z,z^{\prime})d\gamma(z,z^{\prime}),$
-where $z=(x,y)$ is the feature-label pair.
-Note that since $P_{i}=P_{i,x}\otimes\delta_{\nu_{i}^{(P)}}$, $\gamma$ is supported on the set $\{(x,y,x,y^{\prime}):y=\nu_{i}^{(P)},y^{\prime}=\nu_{j}^{(Q)}\}$.
-Thus, the OT plan has the structure:
-$\gamma=\gamma_{x}\otimes\delta_{(\nu_{i}^{(P)},\nu_{j}^{(Q)})}$
-where $\gamma_{x}$ is the OT plan between Gaussians $P_{i,x}$ and $Q_{j,x}$. This means that the previous integral can be decomposed into,
-$\mathbb{W}_{2}(P_{i},Q_{j})^{2}=\int\int(\lVert x-x^{\prime}\rVert_{2}^{2}+\rho(y,y^{\prime})^{2})d\gamma(z,z^{\prime}),$
-which we can decompose two terms. The first,
-$\int\int\lVert x-x^{\prime}\rVert_{2}^{2}d\gamma(z,z^{\prime})$ $=\int\int\lVert x-x^{\prime}\rVert_{2}^{2}d\gamma_{x}(x,x^{\prime})=\mathbb{W}_{2}(P_{i,x},Q_{j,x})^{2},$
-and,
-$\int\int\rho(y,y^{\prime})^{2}d\gamma(z,z^{\prime})$ $=\int\int\rho(y,y^{\prime})^{2}\delta_{(\nu_{i}^{(P)},\nu_{j}^{(Q)})}dydy^{\prime}=\rho(\nu_{i}^{(P)},\nu_{j}^{(Q)})^{2},$
-which gives the desired decomposition. ∎
-The result in the previous proposition stems from the special structure we defined for the GMMs.
-Since the result is agnostic with respect the label metric, it holds for $\rho(\nu_{i}^{(P)},\nu_{j}^{(Q)})=\lVert\nu_{i}^{(P)}-\nu_{j}^{(Q)}\rVert_{2}$
-D.3 Theorem 3.1
-In this section, we present the proof for theorem 3.1. We mainly use two tools to obtain this result,
-  1. 1.
-    The exponential convergence given by the PL inequality,
-    $\lVert\nabla\mkern-10.0mu\nabla\mathbb{B}(P)\rVert_{L_{2}(P)}\geq C_{\text{PL}}(\mathbb{B}(P)-\mathbb{B}^{\star}),$ (47)
-  2. 2.
-    An uniform bound on $P\mapsto\mathbb{B}_{\mathcal{Q}}(P)$ controlling the error of approximating $Q_{k}$ with $\hat{Q}_{k}$.
-These results are available, respectively, in [chewi2024statistical] and [portales2025sample].
-For completeness, we re-state and demonstrate them.
-We start with a modified version of [chewi2024statistical, Corollary 5.17],
-Proposition D.2.
-(Measure theoretic Gronwall’s Lemma [chewi2024statistical]) Let $\mathbb{F}:\mathcal{P}_{2,\text{ac}}(\Omega)\rightarrow\mathbb{R}$ be a functional that satisfies the PL inequality (c.f.
-47) with constant $C_{\text{PL}}>0$, then,
-$\mathbb{F}(P_{\tau})-\mathbb{F}^{\star}\leq e^{-2C_{\text{PL}}\tau}(\mathbb{F}(P_{0})-\mathbb{F}^{\star}).$ (48)
-Proof.
-Define $\phi(t)=\mathbb{F}(P_{t})-\mathbb{F}^{\star}$. Then,
-$\partial_{t}\phi(t)$ $=\partial_{t}\mathbb{F}(P_{t})=-\lVert\nabla\mkern-10.0mu\nabla\mathbb{F}(P_{t})\rVert_{L_{2}(P_{t})}^{2},$
-$\geq-C_{\text{PL}}(\mathbb{F}(P_{0})-\mathbb{F}^{\star})=-C_{\text{PL}}\phi(t).$
-From Grönwall’s Lemma [chewi2024statistical, Lemma 5.16],
-$\underbrace{\mathbb{F}(P_{t})-\mathbb{F}^{\star}}_{\partial_{t}\phi(t)}\leq e^{-C_{\text{PL}}}\underbrace{(\mathbb{F}(P_{0})-\mathbb{F}(P)^{\star})}_{\phi(0)}$
-We now present the uniform approximation bound of [portales2025sample] for empirical barycenters,
-Theorem D.1.
-(Barycenter Empirical Approximation error [portales2025sample]) Let $\mathcal{Q}=\{Q_{1},\cdots,Q_{K}\}$ be a finite family of measures on $\mathcal{P}_{p,\text{ac}}(\mathcal{B}(0,R))$ for $R>0$.
-Let $\{z_{i}^{(Q_{k})}\}_{i=1}^{n_{k}}$, $i=1,\cdots,n_{k}$ be a set of $K$ i.i.d.
-samples from each $Q_{k}$.
-Let $\hat{P}=n^{-1}\sum_{i=1}^{n}\delta_{z_{i}^{(P)}}$.
-Let $p\in[1,+\infty)$.
-Then,
-$\mathbb{E}\biggr[\underset{z_{1}^{(P)},\cdots,z_{n}^{(P)}\in\Omega}{\text{sup}}\,|\mathbb{B}_{\mathcal{Q}}(\hat{P})-\widehat{\mathbb{B}}_{\mathcal{Q}}(\hat{P})|\biggr]\leq C_{p,R}\sqrt{\dfrac{C_{d,m}}{n}},$ (49)
-where,
-$C_{p,R}$ $=8\sqrt{2}\int_{0}^{3(2R)^{p}}\sqrt{\log\biggr(2+\dfrac{32p(2R)^{p}}{s}\biggr)}ds,$
-$C_{d,m}$ $=m(d+1).$
-With these bounds established, we can then prove our Theorem 3.1,
-Proof.
-Let $\tau>0$ be the current optimization step. We have the following,
-[…]
-[…]
-From equation 50, we can bound the expected empirical optimization gap,
-$\mathbb{E}[|\hat{\mathbb{B}}_{\mathcal{Q}}(\hat{P}_{\tau})-\hat{\mathbb{B}}^{\star}_{\mathcal{Q}}|]\leq$ $\mathbb{E}[|\hat{\mathbb{B}}_{\mathcal{Q}}(\hat{P}_{\tau})-\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})|]+\mathbb{E}[|\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})-\mathbb{B}_{\mathcal{Q}}^{\star}|]+\mathbb{E}[|\mathbb{B}_{\mathcal{Q}}^{\star}-\hat{\mathbb{B}}^{\star}_{\mathcal{Q}}|].$ (51)
-Now, we can bound the first and third terms using the approximation error of Portales et al. [portales2025sample] (i.e., equation 49),
-$\mathbb{E}[|\hat{\mathbb{B}}_{\mathcal{Q}}(\hat{P}_{\tau})-\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})|]$ $\leq C_{p,R}\sqrt{\dfrac{C_{d,m}}{n}},$ (52)
-$\mathbb{E}[|\hat{\mathbb{B}}_{\mathcal{Q}}^{\star}-\mathbb{B}_{\mathcal{Q}}^{\star}|]$ $\leq C_{p,R}\sqrt{\dfrac{C_{d,m}}{n}},$
-here, we recall that $\mathbb{B}_{\mathcal{Q}}^{\star}=\mathbb{B}_{\mathcal{Q}}(P^{\star})=\text{inf}_{P}\mathbb{B}_{\mathcal{Q}}(P)$.
-Coming back to the second term, note that the true optimization gap $\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})-\mathbb{B}^{\star}_{\mathcal{Q}}$ is deterministic, as it does not involve sampling from the measures in $\mathcal{Q}$.
-This means that,
-$\mathbb{E}[|\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})-\mathbb{B}_{\mathcal{Q}}^{\star}|]=|\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})-\mathbb{B}_{\mathcal{Q}}^{\star}|\leq e^{-C_{\text{PL}}\tau}(\hat{\mathbb{B}}_{\mathcal{Q}}(\hat{P}_{0})-\hat{\mathbb{B}}_{\mathcal{Q}}^{\star}).$ (53)
-Combining equations 5352 and 51, we get,
-$\mathbb{E}[|\hat{\mathbb{B}}_{\mathcal{Q}}(\hat{P}_{\tau})-\hat{\mathbb{B}}^{\star}_{\mathcal{Q}}|]\leq e^{-C_{\text{PL}}\tau}(\mathbb{B}_{\mathcal{Q}}(\hat{P}_{0})-\mathbb{B}^{\star}_{\mathcal{Q}})+C_{p,R}\sqrt{\dfrac{C_{d,m}}{n}},$
-which is the desired inequality. ∎
-D.4 Theorem 3.2
-The goal of this section is to stablish the proof of theorem 3.2.
-The proof roughly follows the same idea of the empirical case.
-However, since all the involved measures are GMMs by hypothesis, we do not have empirical approximation gaps per se.
-In this regard, we do consider the stochasticity of fitting GMMs, provided data from each underlying $Q_{k}$.
-As a result, we need an approximation tool, in the spirit of theorem D.1 of [portales2025sample].
-We do so through the following result of [boite2025differentiable],
-Proposition D.3.
-(GMM approximation bound [boite2025differentiable]) Let $P$ and $Q$ be GMMs with parameters $\{(\pi^{(P)}_{i},\mu_{i}^{(P)},\Sigma_{i}^{(P)})\}_{i=1}^{n}$ (resp.
-$Q$), and let $\hat{P}$ and $\hat{Q}$ be the corresponding GMMs empirically estimated by the expectation-maximization algorithm.
-Assume the following,
-  • Hypothesis 1 (bounded parameters): There are constants $R_{\mu}>0$ and $R_{\Sigma}>0$ such that,
-    $\lVert\mu_{i}^{(P)}\rVert_{2}\leq R_{\mu},\sqrt{\text{Tr}(\Sigma_{i}^{(P)})}\leq R_{\Sigma},\lVert\hat{\mu}_{i}^{(P)}\rVert_{2}\leq R_{\mu},\sqrt{\text{Tr}(\hat{\Sigma}_{i}^{(P)})}\leq R_{\Sigma}\text{, resp.
-}Q,$
-  • Hypothesis 2 (empirical convergence): Assume that exist $\rho_{\pi}>0$, $\rho_{\mu}$, and $\rho_{\Sigma}>0$ such that,
-    $\mathbb{E}[\lVert\pi^{(P)}-\hat{\pi}^{(P)}\rVert_{1}]\leq\rho_{\pi}\quad\mathbb{E}[\lVert\mu_{i}^{(P)}-\hat{\mu}^{(P)}_{i}\rVert_{2}]$ $\leq\rho_{\mu}\quad\mathbb{E}[d_{\text{bures}}(\Sigma_{i}^{(P)},\hat{\Sigma}_{i}^{(P)})]$ $\leq\rho_{\Sigma}$
-    $i=1,\cdots,n$ (resp. $Q$), where,
-    $d_{\text{bures}}(\Sigma,\Sigma^{\prime})=\sqrt{\text{Tr}(\Sigma+\Sigma^{\prime}-2(\Sigma^{1/2}\Sigma^{\prime}\Sigma^{1/2})^{1/2})}.$
-In these conditions,
-$\mathbb{E}[|\mathbb{MW}_{2}(\hat{P},\hat{Q})^{2}-\mathbb{MW}_{2}(P,Q)^{2}|]\leq\underbrace{8R_{\mu}\rho_{\mu}+8R_{\Sigma}\rho_{\Sigma}+8(R_{\mu}^{2}+R_{\Sigma}^{2})\rho_{\pi}}_{C_{\pi,\mu,\Sigma}},$ (54)
-Note that the same result applies for labeled GMM.
-Indeed, since we fit a GMM per class, the labels are deterministic.
-As a result, $\nu_{i}^{(P)}=\hat{\nu}_{i}^{(P)}$.
-These previous theorems allow us to bound the barycenter functional,
-Corollary D.1.
-Let $\{Q_{k}\}_{k=1}^{K}$ be $K>0$ GMMs over $\Omega$, and let $\{\hat{Q}_{k}\}_{k=1}^{K}$ be their corresponding empirical approximations.
-Assume Hypothesis 1 and 2 of Proposition 4.3.
-Let $P$ be a GMM, and $\lambda\in\Delta_{K}$, then,
-$\mathbb{E}[|\mathbb{B}(P)-\hat{\mathbb{B}}(P)|]\leq\sum_{k=1}^{K}\lambda_{k}C_{k}=C_{\lambda,\mathcal{Q}},$
-where $C_{k}=8R_{\mu_{k}}+8R_{\Sigma_{k}}\rho_{\Sigma_{k}}+8(R_{\mu_{k}}^{2}+R_{\Sigma_{k}}^{2})\rho_{\pi_{k}}$.
-Proof.
-The proof relies on the triangle inequality:
-$|\mathbb{B}(P)-\hat{\mathbb{B}}(P)|$ $\leq\biggr|\sum_{k=1}^{K}\lambda_{k}\mathbb{MW}_{2}(P,Q_{k})^{2}-\sum_{k}\lambda_{k}\mathbb{MW}_{2}(P,\hat{Q}_{k})^{2}\biggr|$
-$\leq\sum_{k=1}^{K}\lambda_{k}|\mathbb{MW}_{2}(P,Q_{k})^{2}-\mathbb{MW}_{2}(P,\hat{Q}_{k})^{2}|,$
-then, taking the expectation on both sides,
-$\mathbb{E}[|\mathbb{B}(P)-\hat{\mathbb{B}}(P)|]\leq\sum_{k=1}^{K}\lambda_{k}\mathbb{E}[|\mathbb{MW}_{2}(P,Q_{k})^{2}-\mathbb{MW}_{2}(P,\hat{Q}_{k})^{2}|]\leq\sum_{k=1}^{K}\lambda_{k}C_{k}=C_{\lambda,\mathcal{Q}}.$
-Furthermore, note that the constant can be written explicitly as,
-$C_{\lambda,\mathcal{Q}}=\sum_{k=1}^{K}\lambda_{k}(8R_{\mu_{k}}+8R_{\Sigma_{k}}\rho_{\Sigma_{k}}+8(R_{\mu_{k}}^{2}+R_{\Sigma_{k}}^{2})\rho_{\pi_{k}})$ (55)
-With these results established, we now prove Theorem 3.2,
-Proof.
-Our arguments largely follow those in section D.3. For instance, the steps taken in equations 50 and 51 directly apply to the GMM setting. Now, in place of equation 52, we have,
-$\mathbb{E}[|\hat{\mathbb{B}}_{\mathcal{Q}}(\hat{P}_{\tau})-\mathbb{B}_{\mathcal{Q}}(\hat{P}_{\tau})|]\leq C_{\lambda,\mathcal{Q}},\quad\mathbb{E}[|\hat{\mathbb{B}}_{\mathcal{Q}}^{\star}-\mathbb{B}_{\mathcal{Q}}^{\star}|]\leq C_{\lambda,\mathcal{Q}}.$ (56)
-Furthermore, equation 53 also holds in the GMM case. Indeed, the optimization gap is deterministic. Therefore, combining equations 56 with 51 and 53, we get,
-$\mathbb{E}[|\hat{\mathbb{B}}_{\mathcal{Q}}(\hat{P}_{\tau})-\hat{\mathbb{B}}^{\star}_{\mathcal{Q}}|]\leq e^{-C_{\text{PL}}\tau}(\mathbb{B}_{\mathcal{Q}}(\hat{P}_{0})-\mathbb{B}^{\star}_{\mathcal{Q}})+C_{\lambda,\mathcal{Q}},$
-which is the desired inequality.
-Appendix E Additional Toy Experiments
-E.1 Two Moons
-Figure 5: Two moons samples alongside the fitted GMMs. Stars and ellipses denote the means and covariance matrices.
-In this experiment, similarly to the Swiss-roll example we explored in the main paper, we deal with Location-Scatter families of measures.
-The idea is that we are given $Q_{0}\in\mathcal{P}_{2,\text{ac}}$, and we define a finite family of measures,
-$\mathcal{Q}=\{T_{k,\sharp}Q_{0}:T_{k}(x)=A_{k}x+b_{k},A_{k}\in\text{PD}(d)\}_{k=1}^{K}$ (57)
-where $\text{PD}(d)$ denotes the set of $d\times d$ positive-definite matrices. The barycenter of $Q_{1},\cdots,Q_{K}$ is itself a push-forward of $Q_{0}$, noted $T$. Furthermore, $T$ has the form,
-$T(x)=\sum_{k=1}^{K}\lambda_{k}T_{k}(x)=\biggr(\sum_{k=1}^{K}\lambda_{k}A_{k}\biggr)x+\sum_{k=1}^{K}\lambda_{k}b_{k}$
-For this experiment, we use $Q_{0}$ as the two moons dataset [^2] More specifically, we use https://scikit-learn.org/stable/modules/generated/sklearn.datasets.make_moons.html to generate data from the two moons measure, and https://scikit-learn.org/stable/modules/generated/sklearn.datasets.make_swiss_roll.html to generate data from the Swiss roll measure..
-For the transformations, we use a family of rotation-shift affine functions, i.e.,
-$A_{k}=\begin{bmatrix}\cos\theta_{k}&-\sin\theta_{k}\\ \sin\theta_{k}&\cos\theta_{k}\end{bmatrix},\quad b_{k}\in\mathbb{R}^{2}.$
-where $\theta_{k}\in\{-30,-15,15,30\}$ degrees, and,
-$b_{k}\in\{\begin{bmatrix}0,0\end{bmatrix},\begin{bmatrix}10,0\end{bmatrix},\begin{bmatrix}0,10\end{bmatrix},\begin{bmatrix}10,10\end{bmatrix}\}$.
-The generated family of measures is shown in Figure 5.
-Figure 6: From left to right, comparison of the empirical barycenter with and without an interaction energy term, and the Gaussian mixture barycenter.
-Overall, due its clustering effect, the GMM barycenter better respects the underlying geometry of the input measures.
-We show a comparison of the empirical and GMM barycenters in Figure 6.
-Note that, without an interaction term, the empirical barycenter mixes a few samples from opposing classes.
-This artifact comes from the rotations in the input measures.
-Adding the interaction energy term $U_{R}$ solves this problem, at the cost of some deformation of the class clusters.
-Finally, the GMM flow is able to compute an accurate barycenter due its clustering effect on the Gaussian components.
-We refer readers to [montesuma2024lighter] and [montesuma2025optimal] for more details on this effect.
-E.2 Handling Class Imbalance
-A challenge classic Wasserstein barycenter algorithms face is handling class imbalance.
-For instance, since the algorithm in [montesuma2023multi] relies on label propagation, the soft-labels in the Wasserstein barycenter might end up mixing classes.
-This phenomenon leads to Wasserstein barycenters with mixed classes and fuzzy labels.
-We tackle these problems with the functionals described in Section 3.3 in the main paper.
-We show, in Figure 7, an example of the imbalance problem.
-Here, $\pi=[\pi_{0},\pi_{1}]$ denotes the class proportions.
-Figure 7: Imbalanced Gaussian blobs for the class imbalance example. We draw $\pi_{0}\sim\mathcal{U}[0,1]$ and let $\pi_{1}=1-\pi_{0}$.
-We start by comparing the joint Wasserstein barycenter algorithm of [montesuma2023multi] (a) with our mini-batch strategy (b).
-Note that using mini-batches already helps with the fuzzy label problem, but many labels are still uncertain in the final barycenter.
-Adding the label entropy penalty $\mathbb{V}$, with underlying functional $V$ defined in equation 20 in the main paper, leads to sharp labels, as we show in (c).
-Finally, by adding the interaction energy term we can separate the classes, as show in Figure 8 (d).
-(a) [montesuma2023multi].
-(b) $\mathbb{B}_{\mathcal{Q}}$ (ours).
-(c) $\mathbb{B}_{\mathcal{Q}}+\mathbb{V}$ (ours).
-(d) $\mathbb{B}_{\mathcal{Q}}+\mathbb{V}+\mathbb{U}$ (ours).
-Figure 8: Discrete (a), and mini-batch Wasserstein barycenters (b and c).
-Overall, the discrete algorithm of [montesuma2023multi] struggles with class imbalances, since the transport plan matches points from different classes.
-This phenomenon leads to fuzzy samples near the class boundary.
-This issue is partly solved through mini-batching (b), but a definitive solution comes with adding the entropy functional defined in equation 20 in the main paper.
-Appendix F Additional details on domain adaptation
-F.1 Evaluation Protocol, Benchmarks, Fine-Grained Results
-Benchmarks.
+E.4 Adapting Neural Network Barycenter Solvers to DA
+Here, we cover how we adapted previous neural network-based barycenters to perform domain adaptation.
+There is one main challenge with using unlabeled barycenters for domain adaptation: we cannot directly move the barycenter to the target domain, as we would end up with more unlabeled data.
+A straightforward solution to this problem is moving the available labeled data that we have, that is, the source domain data.
+For this scenario, the barycenter domain works as a pivot domain, as we transport the source to the target through the composition,
+$T_{Q_{k}\rightarrow Q_{T}}(x_{i}^{(Q_{k})})=T_{P\rightarrow Q_{T}}(T_{Q_{k}\rightarrow P}(x_{i}^{(Q_{k})})).$ (67)
+In comparison, by computing a labeled barycenter, we synthesize data in $P^{\star}$.
+As we demonstrated in Tables 3 through 7, this is generally beneficial for domain adaptation.
+In this way, we can acquire a labeled dataset in the target domain through $\{\{T_{P\rightarrow Q_{T}}(T_{Q_{k}\rightarrow P}(x_{i}^{(Q_{k})})),y_{i}^{(Q_{k})}\}_{i=1}^{n_{k}}\}_{k=1}^{K}$.
+To isolate the effect of the quality of the computed barycenters, we estimate the mappings as we did in equations 65 and 66, that is,
+$T_{Q_{k}\rightarrow P}(x_{i}^{(Q_{k})})=n\sum_{j=1}^{n}\gamma_{k,i,j}^{\star}x_{j}^{(P)}\quad T_{P\rightarrow Q_{k}}(x_{i}^{(P)})=n_{T}\sum_{j=1}^{n_{T}}\gamma_{ij}^{\star}x_{j}^{(Q_{T})},$ (68)
+where $\gamma_{k}^{\star}$ is the OT plan from $Q_{k}$ to $P$, and $\gamma^{\star}$ is the OT plan from $P$ to $Q_{T}$.
+Appendix F Additional Experiments
 In total, we compare five domain adaptation benchmarks, divided into computer vision, neuroscience and chemical engineering.
@@
 Here, we give further details on each of the benchmarks.
-For computer vision, we use the Office 31 [saenko2010adapting] and Office Home [venkateswara2017deep] benchmarks.
-For neuroscience, we use the BCI-CIV-2a [brunner2008bci] and ISRUC [khalighi2013automatic] benchmarks.
-For chemical engineering, we use the TEP benchmark [montesuma2024benchmarking].
-Evaluation Protocol. As we describe in the following, all benchmarks deal with multi-class classification. We measure performance using the standard classification accuracy,
+For computer vision, we use the Office 31 saenko2010adapting and Office Home venkateswara2017deep benchmarks.
+For neuroscience, we use the BCI-CIV-2a brunner2008bci and ISRUC khalighi2013automatic benchmarks.
+For chemical engineering, we use the TEP benchmark montesuma2024benchmarking.
+F.1 Evaluation Protocol and Backbones
+As we describe in the following, all benchmarks deal with multi-class classification. We measure performance using the standard classification accuracy,
 $\text{accuracy}(y,\hat{y})=\dfrac{1}{n}\sum_{i=1}^{n}\delta(y_{i}-\hat{y}_{i})\times 100\%.$
 For the Office 31, Office Home, Tennessee Eastman Process, and BCI-CIV-2a, we adopt the so-called leave-one-domain-out evaluation.
-For instance, given $3$ domains $\{A,B,C\}$, we perform $3$ experiments, leaving one domain as the target domain at each time.
-For instance, one of these experiments would be $\{A,B\}\rightarrow C$, where $\{A,B\}$ are the sources, and $C$ is the target.
-Backbones.
+For example, given $3$ domains $\{A,B,C\}$, we perform $3$ experiments, leaving one domain as the target domain at each time.
+Then, one of these experiments would be $\{A,B\}\rightarrow C$, where $\{A,B\}$ are the sources, and $C$ is the target.
 As we discuss in the main paper, we perform adaptation over embeddings.
-To that end, we use the same experimental setting as [montesuma2024lighter] for the Office 31, Office Home, and TEP benchmarks.
+To that end, we use the same experimental setting as montesuma2024lighter for the Office 31, Office Home, and TEP benchmarks.
 More specifically, we train a ResNet 50 and ResNet 101 for the Office 31 and Office Home benchmarks, and a CNN for the TEP benchmark.
-For this latter backbone, we refer readers to [montesuma2024benchmarking] for further details.
-For neuroscience benchmarks, we follow the experimental protocol of [wang2024cbramod], where we fine-tune CBraMod, a transformer-based foundation model, on available source domain data.
+For this latter backbone, we refer readers to montesuma2024benchmarking for further details.
+For neuroscience benchmarks, we follow the experimental protocol of wang2024cbramod, where we fine-tune CBraMod, a transformer-based foundation model, on available source domain data.
 We use the exact same training settings as these authors.
+F.2 Fine-Grained Results
 Algorithm $\mathcal{X}\times\mathcal{Y}$ A D W Avg. $\uparrow$
 ResNet50 - 67.50 95.00 96.83 86.40
-WJDOT [turrisi2022multi] - 67.77 97.32 95.32 86.80
-DaDiL-E [montesuma2023multi] - 70.55 100.00 98.83 89.79
-DaDiL-R [montesuma2023multi] - 70.90 100.00 98.83 89.91
-GMM-DaDiL [montesuma2024lighter] - 72.47 100.00 99.41 90.63
-Discrete [cuturi2014fast] 64.45 88.39 92.98 81.94
-NormFlow [visentin2025computing] 65.67 97.32 94.73 85.91
-CW2B [korotin2021continuous] 69.16 94.64 95.32 86.37
-NOT [kolesov2024estimating] 69.33 92.85 96.49 86.22
-U-NOT [gazdieva2024robust] 68.29 97.32 95.32 86.97
-Discrete [montesuma2023multi] 67.94 98.21 97.66 87.93
-GMM [montesuma2024lighter] 70.13 99.11 96.49 88.54
-WGF (ours) 71.25 98.21 97.07 88.84
-WGF-GMM (ours) 71.77 99.11 97.07 89.34
-Table 4: Office 31.
-Office 31 [saenko2010adapting] contains, in total, 4652 images among 31 classes of objects.
+WJDOT turrisi2022multi - 67.77 97.32 95.32 86.80
+DaDiL-E montesuma2023multi - 70.55 100.00 98.83 89.79
+DaDiL-R montesuma2023multi - 70.90 100.00 98.83 89.91
+GMM-DaDiL montesuma2024lighter - 72.47 100.00 99.41 90.63
+Discrete cuturi2014fast 70.55 96.43 96.49 87.82
+NormFlow visentin2025computing 65.67 97.32 94.73 85.91
+CW2B korotin2021continuous 69.16 94.64 95.32 86.37
+NOT kolesov2024estimating 69.33 92.85 96.49 86.22
+U-NOT gazdieva2024robust 68.29 97.32 95.32 86.97
+WGF (ours) 69.16 98.21 96.49 87.96
+Discrete montesuma2023multi 67.94 98.21 97.66 87.93
+GMM montesuma2024lighter 70.13 99.11 96.49 88.54
+WGF (ours) 70.03 99.11 99.41 89.52
+Table 3: Classification accuracy per domain for the Office 31 benchmark.
+Office 31 saenko2010adapting contains, in total, 4652 images among 31 classes of objects.
 These images are divided into 3 different domains: Amazon, dSLR, and Webcam.
@@
 There 795 images in this domain.
-As analyzed in [ringwald2021adaptiope], this benchmark is one of the most commonly used in the literature.
+As analyzed in ringwald2021adaptiope, this benchmark is one of the most commonly used in the literature.
 Despite its moderate size, it presents challenges of its own.
@@
 We show our results in the Office 31 benchmark in Table 3.
-Overall, dictionary learning methods such as DaDiL [montesuma2023multi] or GMM-DaDiL [montesuma2024lighter] remain state-of-the-art for this benchmark.
+Overall, dictionary learning methods such as DaDiL montesuma2023multi or GMM-DaDiL montesuma2024lighter remain state-of-the-art for this benchmark.
 However, our gradient flow algorithm is able to improve previous methods, closing the gap between these two kinds of algorithms.
@@
 ResNet101 - 72.90 62.20 83.70 85.00 75.95
-WJDOT [turrisi2022multi] - 74.28 63.80 83.78 84.52 76.59
-DaDiL-E [montesuma2023multi] - 77.16 64.95 85.47 84.97 78.14
-DaDiL-R [montesuma2023multi] - 75.92 64.83 85.36 85.32 77.86
-GMM-DaDiL [montesuma2024lighter] - 77.16 66.21 86.15 85.32 78.81
-Discrete [cuturi2014fast] 68.11 59.79 75.56 78.21 70.42
-NormFlow [visentin2025computing] 77.16 62.54 83.10 84.63 76.86
-CW2B [korotin2021continuous] 74.69 63.46 82.54 85.09 76.44
-NOT [kolesov2024estimating] 72.22 63.57 82.77 82.91 75.36
-U-NOT [gazdieva2024robust] 75.30 64.94 82.88 84.28 76.85
-Discrete [montesuma2023multi] 75.72 63.80 84.23 84.63 77.09
-GMM [montesuma2024lighter] 75.31 64.26 86.71 85.21 77.87
-WGF (ours) 74.48 65.17 85.92 85.77 77.84
-WGF-GMM (ours) 74.89 65.97 87.16 86.81 78.71
-Table 5: Office-Home.
-Office-Home [venkateswara2017deep] contains, 15,500 images among 65 classes of objects.
+WJDOT turrisi2022multi - 74.28 63.80 83.78 84.52 76.59
+DaDiL-E montesuma2023multi - 77.16 64.95 85.47 84.97 78.14
+DaDiL-R montesuma2023multi - 75.92 64.83 85.36 85.32 77.86
+GMM-DaDiL montesuma2024lighter - 77.16 66.21 86.15 85.32 78.81
+Discrete cuturi2014fast 76.13 63.91 83.11 85.55 77.17
+NormFlow visentin2025computing 77.16 62.54 83.10 84.63 76.86
+CW2B korotin2021continuous 74.69 63.46 82.54 85.09 76.44
+NOT kolesov2024estimating 72.22 63.57 82.77 82.91 75.36
+U-NOT gazdieva2024robust 75.30 64.94 82.88 84.28 76.85
+WGF (ours) 74.89 64.60 83.89 85.21 77.56
+Discrete montesuma2023multi 76.54 63.91 83.78 84.74 77.25
+GMM montesuma2024lighter 75.31 64.26 86.71 85.21 77.87
+WGF (ours) 76.13 65.52 85.81 86.24 78.42
+Table 4: Classification accuracy per domain for the Office-Home benchmark.
+Office-Home venkateswara2017deep contains, 15,500 images among 65 classes of objects.
 There are four domains: Art, Clipart, Product, and Real-World.
@@
 In this sense, the Product domain is similar to the Amazon domain in the Office 31 benchmark, and the Real-World is similar to dSLR or Webcam.
-In comparison with the Amazon benchmark, the Office-Home benchmark adds another factor of domain variation: style (e.g., Art vs.
+In comparison with the Office31 benchmark, the Office-Home benchmark adds another factor of domain variation: style (e.g., Art vs.
 Real-World).
-Furthermore, due the larger number of classes, one might expect a harder classification problem.
 We show our results in Table 4.
 In comparison with the Office 31 benchmark, our methods get even closer to the state-of-the-art.
-For instance, WGF-GMM has only a $0.10\%$ gap in performance with GMM-DaDiL, while showing state-of-the-art performance on domains Product and Real-World.
-Again, we verify that using labels in the ground cost ($\mathcal{X}\times\mathcal{Y}$ = ✓) offers a clear advantage.
+Again, using labels in the ground cost ($\mathcal{X}\times\mathcal{Y}$ = ✓) offers a clear advantage.
 Remark F.1.
-We follow previous works on visual domain adaptation, especially [saenko2010adapting] and [venkateswara2017deep], and report the classification accuracy on fixed partitions of the target domains.
-For that reason, Tables 4 and 5 do not have confidence intervals.
-While it would be possible to run these experiments with random partitions (e.g., $k$-fold cross validation), this would mean that our results would not be comparable to previous established results in [montesuma2024lighter], for instance.
-Tennessee-Eastman Process (TEP) [montesuma2024benchmarking] is a benchmark in chemical engineering.
+We follow previous works on visual domain adaptation, especially saenko2010adapting and venkateswara2017deep, and report the classification accuracy on fixed partitions of the target domains.
+For that reason, Tables 3 and 4 do not have confidence intervals.
+While it would be possible to run these experiments with random partitions (e.g., $k$-fold cross validation), this would mean that our results would not be comparable to previous established results in montesuma2024lighter, for instance.
+Tennessee-Eastman Process (TEP) [reinartz2021extended, montesuma2024benchmarking] is a benchmark in chemical engineering.
 This benchmark comprises a set of simulations of a chemical reaction that occurs inside a reactor.
@@
 As a result, each domain corresponds to one of these modes of operation.
-We refer readers to [montesuma2024benchmarking] and [reinartz2021extended] for more details.
+We refer readers to montesuma2024benchmarking and reinartz2021extended for more details.
 Algorithm $\mathcal{X}\times\mathcal{Y}$ Mode 1 Mode 2 Mode 3 Mode 4 Mode 5 Mode 6 Avg. $\uparrow$
-CNN 80.82 $\pm$ 0.96 63.69 $\pm$ 1.71 87.47 $\pm$ 0.99 79.96 $\pm$ 1.07 74.44 $\pm$ 1.52 84.53 $\pm$ 1.12 78.48
-WJDOT [turrisi2022multi] - 89.06 $\pm$ 1.34 75.60 $\pm$ 1.84 89.99 $\pm$ 0.86 89.38 $\pm$ 0.77 85.32 $\pm$ 1.29 87.43 $\pm$ 1.23 86.13
-DaDiL-E [montesuma2023multi] - 90.45 $\pm$ 1.02 77.08 $\pm$ 1.21 86.79 $\pm$ 2.14 89.01 $\pm$ 1.35 84.04 $\pm$ 3.16 87.85 $\pm$ 1.06 85.87
-DaDiL-R [montesuma2023multi] - 91.97 $\pm$ 1.22 77.15 $\pm$ 1.32 85.41 $\pm$ 1.69 89.39 $\pm$ 1.03 84.49 $\pm$ 1.95 88.44 $\pm$ 1.29 86.14
-GMM-DaDiL [montesuma2024lighter] - 91.72 $\pm$ 1.41 76.41 $\pm$ 1.89 89.68 $\pm$ 1.49 89.18 $\pm$ 1.17 86.05 $\pm$ 1.46 88.02 $\pm$ 1.12 86.85
-Discrete [cuturi2014fast] 90.38 $\pm$ 1.29 71.17 $\pm$ 1.26 85.27 $\pm$ 1.09 87.29 $\pm$ 0.99 83.83 $\pm$ 1.74 84.94 $\pm$ 1.87 83.81
-NormFlow [visentin2025computing] 87.06 $\pm$ 0.57 66.25 $\pm$ 1.11 88.58 $\pm$ 1.00 89.35 $\pm$ 0.34 80.99 $\pm$ 2.21 85.08 $\pm$ 1.45 82.89
-CW2B [korotin2021continuous] 91.72 $\pm$ 0.79 73.71 $\pm$ 1.42 88.37 $\pm$ 1.31 89.32 $\pm$ 1.14 84.66 $\pm$ 1.75 87.19 $\pm$ 1.24 85.83
-NOT [kolesov2024estimating] 90.31 $\pm$ 1.08 72.37 $\pm$ 0.94 87.99 $\pm$ 1.31 89.28 $\pm$ 0.81 84.35 $\pm$ 1.68 85.29 $\pm$ 2.02 84.93
-U-NOT [gazdieva2024robust] 90.82 $\pm$ 0.65 72.68 $\pm$ 1.58 88.65 $\pm$ 0.95 89.00 $\pm$ 1.28 85.22 $\pm$ 1.07 86.19 $\pm$ 1.70 85.43
-Discrete [montesuma2023multi] 92.38 $\pm$ 0.66 73.74 $\pm$ 1.07 88.89 $\pm$ 0.85 89.38 $\pm$ 1.26 85.53 $\pm$ 1.35 86.60 $\pm$ 1.63 86.09
-GMM [montesuma2024lighter] 92.23 $\pm$ 0.70 71.81 $\pm$ 1.78 84.72 $\pm$ 1.92 89.28 $\pm$ 1.55 87.51 $\pm$ 1.73 82.49 $\pm$ 1.81 84.67
-WGF (ours) 92.41 $\pm$ 0.90 74.16 $\pm$ 1.29 88.78 $\pm$ 0.91 89.84 $\pm$ 0.62 84.66 $\pm$ 2.37 87.39 $\pm$ 2.11 86.21
-WGF-GMM (ours) 92.55 $\pm$ 1.12 75.95 $\pm$ 1.51 89.54 $\pm$ 1.11 89.21 $\pm$ 1.10 85.63 $\pm$ 1.17 87.74 $\pm$ 1.48 86.77
-Table 6: TEP.
+CNN - 80.82 $\pm$ 0.96 63.69 $\pm$ 1.71 87.47 $\pm$ 0.99 79.96 $\pm$ 1.07 74.44 $\pm$ 1.52 84.53 $\pm$ 1.12 78.48
+WJDOT turrisi2022multi - 89.06 $\pm$ 1.34 75.60 $\pm$ 1.84 89.99 $\pm$ 0.86 89.38 $\pm$ 0.77 85.32 $\pm$ 1.29 87.43 $\pm$ 1.23 86.13
+DaDiL-E montesuma2023multi - 90.45 $\pm$ 1.02 77.08 $\pm$ 1.21 86.79 $\pm$ 2.14 89.01 $\pm$ 1.35 84.04 $\pm$ 3.16 87.85 $\pm$ 1.06 85.87
+DaDiL-R montesuma2023multi - 91.97 $\pm$ 1.22 77.15 $\pm$ 1.32 85.41 $\pm$ 1.69 89.39 $\pm$ 1.03 84.49 $\pm$ 1.95 88.44 $\pm$ 1.29 86.14
+GMM-DaDiL montesuma2024lighter - 91.72 $\pm$ 1.41 76.41 $\pm$ 1.89 89.68 $\pm$ 1.49 89.18 $\pm$ 1.17 86.05 $\pm$ 1.46 88.02 $\pm$ 1.12 86.85
+Discrete cuturi2014fast 90.38 $\pm$ 1.29 71.17 $\pm$ 1.26 85.27 $\pm$ 1.09 87.29 $\pm$ 0.99 83.83 $\pm$ 1.74 84.94 $\pm$ 1.87 83.81
+NormFlow visentin2025computing 87.06 $\pm$ 0.57 66.25 $\pm$ 1.11 88.58 $\pm$ 1.00 89.35 $\pm$ 0.34 80.99 $\pm$ 2.21 85.08 $\pm$ 1.45 82.89
+CW2B korotin2021continuous 91.72 $\pm$ 0.79 73.71 $\pm$ 1.42 88.37 $\pm$ 1.31 89.32 $\pm$ 1.14 84.66 $\pm$ 1.75 87.19 $\pm$ 1.24 85.83
+NOT kolesov2024estimating 90.31 $\pm$ 1.08 72.37 $\pm$ 0.94 87.99 $\pm$ 1.31 89.28 $\pm$ 0.81 84.35 $\pm$ 1.68 85.29 $\pm$ 2.02 84.93
+U-NOT gazdieva2024robust 90.82 $\pm$ 0.65 72.68 $\pm$ 1.58 88.65 $\pm$ 0.95 89.00 $\pm$ 1.28 85.22 $\pm$ 1.07 86.19 $\pm$ 1.70 85.43
+WGF (ours) 91.51 $\pm$ 1.32 73.77 $\pm$ 1.24 87.44 $\pm$ 1.16 89.38 $\pm$ 0.62 84.87 $\pm$ 1.75 86.08 $\pm$ 2.23 85.51
+Discrete montesuma2023multi 92.38 $\pm$ 0.66 73.74 $\pm$ 1.07 88.89 $\pm$ 0.85 89.38 $\pm$ 1.26 85.53 $\pm$ 1.35 86.60 $\pm$ 1.63 86.09
+GMM montesuma2024lighter 92.23 $\pm$ 0.70 71.81 $\pm$ 1.78 84.72 $\pm$ 1.92 89.28 $\pm$ 1.55 87.51 $\pm$ 1.73 82.49 $\pm$ 1.81 84.67
+WGF (ours) 92.34 $\pm$ 1.05 76.10 $\pm$ 1.56 89.78 $\pm$ 0.90 89.32 $\pm$ 1.03 85.92 $\pm$ 1.35 87.74 $\pm$ 1.52 86.87
+Table 5: Classification accuracy per domain on the Tennessee Eastman Process benchmark.
 We report our results in Table 5.
 For this benchmark, the gap between unlabeled ($\mathcal{X}\times\mathcal{Y}$ = ✗) and labeled ($\mathcal{X}\times\mathcal{Y}$ = ✓) barycenter methods is closer.
-For instance, CW2B [korotin2021continuous] achieves 85.83 for the average target domain performance versus 86.09 for the discrete barycenter.
-Furthermore, it surpasses the GMM method of [montesuma2024lighter].
-In comparison, our methods establish a new state-of-the-art for barycenter-based domain adaptation in this benchmark.
-Note that our WGF-GMM is able to get the second best performance overall, close to GMM-DaDiL.
+For instance, CW2B korotin2021continuous achieves 85.83 for the average target domain performance versus 86.09 for the labeled discrete barycenter.
+Furthermore, it surpasses the GMM method of montesuma2024lighter.
+In comparison, our methods establish a new state-of-the-art for barycenter-based domain adaptation in this benchmark, having similar performance to GMM-DaDiL montesuma2024lighter.
 Algorithm $\mathcal{X}\times\mathcal{Y}$ 1 2 3 4 5 6 7 8 9 Avg. $\uparrow$
 CBraMod - 52.22 40.45 64.23 40.41 47.74 48.34 53.64 50.52 55.03 50.30
-DaDiL-R [montesuma2023multi] - 58.68 40.27 66.14 41.45 59.54 51.38 53.12 57.46 55.21 53.69
-GMM-DaDiL [montesuma2024lighter] - 60.59 42.36 70.31 48.12 61.11 50.34 62.15 58.85 60.00 57.10
-Discrete [cuturi2014fast] 60.93 43.40 69.96 50.00 61.28 50.69 61.63 61.45 57.11 57.38
-NormFlow [visentin2025computing] 57.11 43.22 67.53 47.71 60.24 51.04 56.25 60.76 55.20 55.45
-CW2B [korotin2021continuous] 60.00 43.57 70.13 48.95 60.93 51.21 61.28 61.63 57.46 57.25
-NOT [kolesov2024estimating] 61.28 43.22 70.31 47.91 60.59 51.21 61.46 61.80 57.29 57.23
-U-NOT [gazdieva2024robust] 60.59 44.09 69.44 47.70 61.11 50.34 61.97 61.11 58.15 57.17
-Discrete [montesuma2023multi] 60.76 42.53 70.66 48.54 61.28 50.69 61.63 61.28 59.54 57.43
-GMM [montesuma2024lighter] 60.59 42.53 69.79 47.08 61.45 50.34 62.32 59.54 59.72 57.04
-WGF (ours) 61.63 43.40 71.00 47.91 61.45 51.38 61.97 61.63 57.12 57.50
-WGF-GMM (ours) 60.59 43.05 70.48 47.29 61.80 51.56 62.15 61.80 57.46 57.35
-Table 7: BCI-CIV-2a.
-BCI-CIV-2a [brunner2008bci] Contains electroencephalogram (EEG) data from nine subjects.
+DaDiL-R montesuma2023multi - 58.68 40.27 66.14 41.45 59.54 51.38 53.12 57.46 55.21 53.69
+GMM-DaDiL montesuma2024lighter - 60.59 42.36 70.31 48.12 61.11 50.34 62.15 58.85 60.00 57.10
+Discrete cuturi2014fast 60.93 43.40 69.96 50.00 61.28 50.69 61.63 61.45 57.11 57.38
+NormFlow visentin2025computing 57.11 43.22 67.53 47.71 60.24 51.04 56.25 60.76 55.20 55.45
+CW2B korotin2021continuous 60.00 43.57 70.13 48.95 60.93 51.21 61.28 61.63 57.46 57.25
+NOT kolesov2024estimating 61.28 43.22 70.31 47.91 60.59 51.21 61.46 61.80 57.29 57.23
+U-NOT gazdieva2024robust 60.59 44.09 69.44 47.70 61.11 50.34 61.97 61.11 58.15 57.17
+WGF (ours) 60.93 43.75 69.96 49.16 60.76 51.38 61.97 63.54 57.64 57.68
+Discrete montesuma2023multi 60.76 42.53 70.66 48.54 61.28 50.69 61.63 61.28 59.54 57.43
+GMM montesuma2024lighter 60.59 42.53 69.79 47.08 61.45 50.34 62.32 59.54 59.72 57.04
+WGF (ours) 60.93 46.18 68.92 50.21 62.32 50.69 61.63 61.11 57.11 57.68
+Table 6: Classification accuracy per domain on the BCI-CIV-2a benchmark.
+BCI-CIV-2a brunner2008bci Contains electroencephalogram (EEG) data from nine subjects.
 In brief, an EEG measures the electrical activity of the brain over time.
@@
 We show our results in Table 6.
-For cross-subject studies, we highlight that, besides improving individual domain performance, it is desirable to not degrade the source-only performance.
+For cross-subject studies, we highlight that, besides improving individual domain performance, it is desirable to not degrade over the source-only performance.
 This is indeed the case for almost all methods.
-For instance, DaDiL [montesuma2023multi] degrades performance of subjects 2 and 7 by a small margin.
+For instance, DaDiL montesuma2023multi degrades performance of subjects 2 and 7 by a small margin.
 For this benchmark, the difference between unlabeled ($\mathcal{X}\times\mathcal{Y}$ = ✗) and labeled ($\mathcal{X}\times\mathcal{Y}$ = ✓) is not as marked as before.
 We hypothesize that this benchmark falls into the covariate shift hypothesis, namely, $Q_{S}(X)\neq Q_{T}(X)$, rather than the more general joint shift $Q_{S}(X,Y)\neq Q_{T}(X,Y)$.
-Algorithm 1 2 3 4 5 6 7 8 9 10 Avg. $\uparrow$
-CBraMod 78.97 80.35 70.46 85.59 80.69 72.85 77.00 73.29 83.41 68.95 76.63
-WJDOT [turrisi2022multi] 80.41 78.25 65.69 82.85 81.74 77.74 76.33 68.75 86.95 70.81 76.95
-DaDiL-E [montesuma2023multi] 76.12 77.90 65.93 83.21 81.28 68.45 75.66 72.61 79.75 65.93 75.14
-DaDiL-R [montesuma2023multi] 78.06 78.13 66.97 84.64 81.51 72.02 76.33 72.84 80.85 67.55 75.89
-GMM-DaDiL [montesuma2024lighter] 76.63 78.95 66.51 82.73 77.44 73.81 77.00 71.70 83.05 66.86 75.47
-Discrete [cuturi2014fast] 78.77 80.58 69.76 85.83 79.06 73.45 78.33 72.50 84.14 68.48 77.09
-NormFlow [visentin2025computing] 81.12 82.90 75.58 84.28 80.34 78.69 75.88 70.34 80.48 70.81 78.05
-CW2B [korotin2021continuous] 79.08 78.25 69.07 86.54 78.48 70.35 76.88 70.90 82.80 66.04 75.84
-NOT [kolesov2024estimating] 78.57 79.88 67.32 85.59 78.72 73.21 77.77 72.50 83.41 67.21 76.44
-U-NOT [gazdieva2024robust] 80.20 83.48 73.72 87.50 76.51 72.62 76.66 70.22 78.17 71.51 77.06
-Discrete [montesuma2023multi] 78.87 80.69 72.09 86.07 78.60 78.45 78.00 72.95 85.48 70.81 78.20
-GMM [montesuma2024lighter] 75.20 75.46 69.88 80.71 78.60 71.90 76.33 68.41 78.17 68.14 74.28
-WGF (ours) 80.20 81.62 76.86 86.42 79.30 80.00 75.88 76.13 86.95 74.42 79.78
-WGF-GMM (ours) 79.89 81.74 72.32 86.42 80.69 79.04 77.55 72.72 85.73 71.04 78.72
-Table 8: ISRUC.
-ISRUC [khalighi2013automatic] is a benchmark for EEG data for sleep staging.
+Algorithm $\mathcal{X}\times\mathcal{Y}$ 1 2 3 4 5 6 7 8 9 10 Avg. $\uparrow$
+CBraMod - 78.97 80.35 70.46 85.59 80.69 72.85 77.00 73.29 83.41 68.95 76.63
+WJDOT turrisi2022multi - 80.41 78.25 65.69 82.85 81.74 77.74 76.33 68.75 86.95 70.81 76.95
+DaDiL-E montesuma2023multi - 76.12 77.90 65.93 83.21 81.28 68.45 75.66 72.61 79.75 65.93 75.14
+DaDiL-R montesuma2023multi - 78.06 78.13 66.97 84.64 81.51 72.02 76.33 72.84 80.85 67.55 75.89
+GMM-DaDiL montesuma2024lighter - 76.63 78.95 66.51 82.73 77.44 73.81 77.00 71.70 83.05 66.86 75.47
+Discrete cuturi2014fast 78.77 80.58 69.76 85.83 79.06 73.45 78.33 72.50 84.14 68.48 77.09
+NormFlow visentin2025computing 81.12 82.90 75.58 84.28 80.34 78.69 75.88 70.34 80.48 70.81 78.05
+CW2B korotin2021continuous 79.08 78.25 69.07 86.54 78.48 70.35 76.88 70.90 82.80 66.04 75.84
+NOT kolesov2024estimating 78.57 79.88 67.32 85.59 78.72 73.21 77.77 72.50 83.41 67.21 76.44
+U-NOT gazdieva2024robust 80.20 83.48 73.72 87.50 76.51 72.62 76.66 70.22 78.17 71.51 77.06
+WGF (ours) 79.69 81.86 70.58 86.66 78.14 77.74 78.77 72.16 85.97 70.00 78.16
+Discrete montesuma2023multi 78.87 80.69 72.09 86.07 78.60 78.45 78.00 72.95 85.48 70.81 78.20
+GMM montesuma2024lighter 75.20 75.46 69.88 80.71 78.60 71.90 76.33 68.41 78.17 68.14 74.28
+WGF (ours) 80.10 80.81 76.62 85.95 76.86 80.35 76.55 77.72 84.87 73.95 80.02
+Table 7: Classification accuracy per domain on the ISRUC benchmark.
+ISRUC khalighi2013automatic is a benchmark for EEG data for sleep staging.
 In brief, sleep staging is the classification of physiological data, such as EEG, recorded during sleep into different sleep stages.
@@
 This means that we compute Wasserstein barycenters of $K=90$ input measures.
-In this benchmark, using our gradient flow strategy yields the best result overall (e.g., $79.78$ average domain performance for our empirical gradient flow).
-Furthermore, we see an important gap between the GMM barycenter algorithm of [montesuma2024lighter] and ours, which is similar to the gap found in the TEP benchmark (c.f.
-Table 6).
-There are mainly two reasons for this.
-First, we can freely choose the learning rate for updating the GMM components, which leads to an overall more stable algorithm.
-Second, we can regularize the barycenter problem, with additional functionals.
-F.2 Adapting Neural Network Barycenter Methods to Domain Adaptation
-In this section, we cover how we adapted previous neural network-based barycenters to perform domain adaptation.
-There is one main challenge with using unlabeled barycenters for domain adaptation: we cannot directly move the barycenter to the target domain, as we would end up with more unlabeled data.
-A straightforward solution to this problem is moving the available labeled data that we have, that is, the source domain data.
-For this scenario, the barycenter domain works as a pivot domain, as we transport the source to the target through the composition,
-$T_{Q_{k}\rightarrow Q_{T}}(x_{i}^{(Q_{k})})=T_{P\rightarrow Q_{T}}(T_{Q_{k}\rightarrow P}(x_{i}^{(Q_{k})})).$
-In comparison, by computing a labeled barycenter, we are synthesizing data in $P^{\star}$.
-As we demonstrated in Tables 4 through 8, this is generally beneficial for domain adaptation.
-In this way, we can acquire a labeled dataset in the target domain through $\{\{T_{P\rightarrow Q_{T}}(T_{Q_{k}\rightarrow P}(x_{i}^{(Q_{k})})),y_{i}^{(Q_{k})}\}_{i=1}^{n_{k}}\}_{k=1}^{K}$.
-To isolate the effect of the quality of the computed barycenters, we estimate the mappings as we did in equations 43 and $4$, that is,
-$T_{Q_{k}\rightarrow P}(x_{i}^{(Q_{k})})=n\sum_{j=1}^{n}\gamma_{k,i,j}^{\star}x_{j}^{(P)}\quad T_{P\rightarrow Q_{k}}(x_{i}^{(P)})=n_{T}\sum_{j=1}^{n_{T}}\gamma_{ij}^{\star}x_{j}^{(Q_{T})},$
-where $\gamma_{k}^{\star}$ is the OT plan from $Q_{k}$ to $P$, and $\gamma^{\star}$ is the OT plan from $P$ to $Q_{T}$.
-F.3 Complexity Analysis and Running Time
-Complexity Analysis. We start with analyzing the complexity of evaluating the functionals used in our work. We recall that,
+Overall, our WGF method has clear advantage over all other methods, with unsupervised WGF (78.16% average target accuracy) being the best over unlabeled barycenters, and labeled barycenters as well (80.02% average test accuracy).
+F.3 Complexity Analysis and Additional Running Time
+We start with analyzing the complexity of evaluating the functionals used in our work. We recall that,
 $\mathbb{V}(P)=\dfrac{1}{n}\sum_{i=1}^{n}V(z_{i}^{(P)}),\quad\mathbb{U}(P)=\dfrac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}U(z_{i}^{(P)},z_{j}^{(P)}),$
 where, assuming that the complexity of evaluating $V$ and $U$ is $\mathcal{O}(1)$, the complexity of evaluating $\mathbb{V}$ is $\mathcal{O}(n)$ and for $\mathbb{U}$ is $\mathcal{O}(n^{2})$.
-For $\mathbb{B}_{\mathcal{Q}}$, the complexity depends on the number of elements used in OT, and the number of input measures.
-In other words, we have $\mathcal{O}(Kn_{b}^{3}\log n_{b})$ for the empirical case, and $\mathcal{O}(Kn_{\text{components}}\log n_{\text{components}})$ for WGF-GMM.
-We assume this cost dominates over the evaluation of other functionals.
-Method Complexity per iteration Parameter Complexity Running time per iteration (s) # Iterations Running Time (s)
-Discrete [cuturi2014fast, montesuma2023multi] $Kn^{3}\log n$ $n$ 4.28 50 214.37
-NormFlow [visentin2025computing] N/A $|f|$ 0.17 5000 878.32
-CW2B [korotin2021continuous] N/A $K(|f|+|g|)$ 2.18 1000 2181.00
-NOT [kolesov2024estimating] N/A $K(|f|+|g|)$ 41.63 25 1040.75
-U-NOT [gazdieva2024robust] N/A $K(|f|+|g|)$ 13.29 100 1329.77
-WGF (ours) $Kn_{b}^{3}\log n_{b}$ $n$ 0.58 200 117.24
-WGF-GMM (ours) $Kn_{\text{components}}^{3}\log n_{\text{components}}$ $n_{\text{components}}$ 0.11 150 17.14
-Table 9: Complexity and running time (in seconds) of different barycenter calculation strategies on the ISRUC benchmark.
-$n_{b}$ denotes the batch size, $n$ number of samples, $K$ number of input measures, $d$ number of dimensions, $n_{\text{components}}$ number of components in GMMs.
-For NormFlow, $f$ denotes the neural net implementing the normalizing flow.
-For CW2b, NOT and U-NOT, $f$ and $g$ denote the neural networks approximating the Kantorovich potentials.
-$|f|$ and $|g|$ denote the number of their parameters.
-Running Time.
-We perform a running time analysis of the tested methods on the ISRUC benchmark, which contains the most number of samples and domains.
+The complexity of $\mathbb{B}$ is usually the bottleneck, as it involves computing $K$ OT problems.
+For exact OT, this means we need to solve $K$ OT problems of size $m\times n$, where $m$ is the batch size and $n$ is the barycenter support size.
+These OT problems need to be solved sequentially, as parallelizing $K$ independent linear programs is non-trivial.
+In practice, we use the network simplex solver of bonneel2011displacement, wrapped in Python by flamary2021pot.
+This algorithm has worst-case complexity of $\mathcal{O}(n^{3})$ but in practice, runs in $\mathcal{O}(n^{2})$.
+Assuming $n\gg m$, the complexity becomes,
+$\text{Exact barycenter functional complexity}=\mathcal{O}(K\times m\times n),$
+which scales linearly with the barycenter support size $n$.
+Though this might appear contradictory with the usual $\mathcal{O}(n^{3})$ complexity of OT, we highlight that we are keeping $m$ constant as we scale the barycenter support size $n$.
+The same analysis can be done for entropic OT,
+$\text{Entropic barycenter functional complexity}=\mathcal{O}(K\times L\times m\times n),$
+where, instead of a factor $m^{2}$, we have $L\times m$, where $L>0$ is the number of Sinkhorn iterations.
+Both of these approaches scale linearly with the number of samples in the barycenter support.
+However, there are 2 important advantages of using entropic OT.
+First, we can use GPU acceleration for computing the OT variables.
+Second, as we shown in Section 3.1, we can vectorize the $K$ OT problems.
+While this does not alleviate the $K$ term in the complexity, it does lead to further speedups.
+These factors explain the difference, in running time, between exact and entropic OT.
+(a) Scaling with $n$ ($\epsilon=0$)
+(b) Scaling with $n$ ($\epsilon>0$)
+(c) Scaling with $\epsilon$ ($m=256$)
+Figure 8: Running time analysis in the Swiss roll example.
+In (a) and (b), we show that our WGF approach improves the scaling of the discrete solver of cuturi2014fast.
+In (c), we isolate the speedups of GPU parallelism and Sinkhorn vectorization.
+Next, we perform a running time analysis of the tested methods on the ISRUC benchmark, which contains the most number of samples and domains.
 For all neural net methods, we fix a batch size of $256$ samples.
 Overall, our proposed methods run significantly faster than previous methods.
-For instance, in comparison with the discrete solver of [cuturi2014fast], our algorithm is approximately seven times faster per iteration.
+For instance, in comparison with the discrete solver of cuturi2014fast, our algorithm is approximately seven times faster per iteration.
 It is noteworthy that some neural net methods are relatively faster per iteration (e.g.
@@
 5000 iterations).
-F.4 Hyperparameter Setting
+Method Complexity per iteration Parameter Complexity Running time per iteration (s) # Iterations Running Time (s)
+Discrete cuturi2014fast, montesuma2023multi $Kn^{3}\log n$ $n$ 4.28 50 214.37
+NormFlow visentin2025computing N/A $|f|$ 0.17 5000 878.32
+CW2B korotin2021continuous N/A $K(|f|+|g|)$ 2.18 1000 2181.00
+NOT kolesov2024estimating N/A $K(|f|+|g|)$ 41.63 25 1040.75
+U-NOT gazdieva2024robust N/A $K(|f|+|g|)$ 13.29 100 1329.77
+WGF (ours, $\epsilon=0$) $Km^{2}n$ $n$ 0.05 200 10.43
+WGF (ours, $\epsilon=10^{-3}$) KLmnKLmn $n$ 0.03 200 7.00
+Table 8: Complexity and running time (in seconds) of different barycenter calculation strategies on the ISRUC benchmark.
+$n_{b}$ denotes the batch size, $n$ number of samples, $K$ number of input measures, $d$ number of dimensions.
+For NormFlow, $f$ denotes the neural net implementing the normalizing flow.
+For CW2B, NOT and U-NOT, $f$ and $g$ denote the neural networks approximating the Kantorovich potentials.
+$|f|$ and $|g|$ denote the number of their parameters.
+F.4 Effect of Different Functionals
+(a) $(10^{-3},10^{-3})$
+(b) $(10^{-4},10^{-3})$
+(c) $(10^{-5},10^{-3})$
+(d) $(10^{-3},0.0)$
+(e) $(10^{-4},0.0)$
+(f) $(10^{-5},0.0)$
+Figure 9: Wasserstein barycenter as a function of $(\eta_{G},\epsilon)$, i.e., diffusion coefficient $\xi$, and entropy regularization $\epsilon$.
+On the effect of $\mathbb{G}$.
+The internal energy functional [chizat2025doubly] controls the smoothness of $\mathbb{G}$, and acts, in the case of the entropy functional $G(s)=s\log s$, as a diffusion term.
+In Figure 9, we show the effect of varying $\xi$ (c.f.
+equation 16) in the final barycenter.
+Overall, an aggressive diffusion coefficient leads to fuzzy barycenters (c.f.
+Figure 9 (a)), while a coefficient too small leads to low variance measures (c.f.
+Figure 9 (c)).
+This is, in fact, an artifact of the entropic OT.
+While a large diffusion coefficient does lead to fuzzy barycenters with exact OT (c.f.
+Figure 9 (d)), using a small coefficient does not degrade the quality of the computed barycenter.
+(a) Unsupervised
+(b) Unsupervised
+(c) Supervised
+(d) Supervised
+Figure 10: t-SNE embeddings of unsupervised (a, b) and supervised (c, d) barycenters.
+While it is possible to group some samples according their pseudo-labels based on $\mathbb{U}$, the task is in general difficult.
+For the supervised case, the interaction energy is able to make classes more separable (d, in comparison to c).
+On the effect of $\mathbb{U}$.
+We now visualize the effect of adding the repulsion interaction energy functional in the Office-Home benchmark.
+For computer vision experiments we use ResNets [he2016deep] as the backbone, which produce $2048$-dimensional embeddings.
+For that reason, we use the cosine distance $d(x,x^{\prime})=1-\text{cossim}(x,x^{\prime})$ in equation 22 and we fix $\text{margin}=1$, which encourages class clusters to live in orthogonal sub-spaces of the latent space.
+Overall, adding $\mathbb{U}$ leads to well separated classes.
+Figure 11: Distribution of label entropy values as a function of potential energy scaling $\eta$.
+In Figure 10, we show a comparison between unsupervised and supervised regularized barycenters.
+In (a), unsupervised, unregularized barycenters do not have any label structure, since there is no label supervision.
+In (b), we show that by adding the interaction energy functional, we can cluster samples with the same assigned label – even without input measure label supervision.
+In (c) and (d), we show the supervised case, in which labels in the input measures are available.
+In (c), classes are well clustered, but some of them overlap in the embedding space.
+Adding the interaction energy functional $\mathbb{U}$ effectively pushes these apart, as shown in (d).
+Overall, we conclude that enforcing well separated class clusters is not straightforward without the inductive bias coming from the labels.
+On the effect of $\mathbb{V}$. In Figure 11, we show the distribution of label entropy values (c.f. equation 22) as a function of the potential scaling $\eta$,
+$\mathbb{V}(P)=\frac{\eta}{n}\sum_{i=1}^{n}V(y_{i}^{(P)}).$
+We see that increasing the weight of this functional leads to increasingly sharp labels.
+F.5 Hyperparameter Setting
 In this section, we detail the hyperparameters of our methods. For both flows, we have mainly three parameters,
-  • Number of samples in the barycenter support, $n$. This hyperparameter is replaced by the number of components in the GMM barycenter, $K$.
+  • number of samples in the barycenter support, $n$.
   • learning rate $\alpha$ for the gradient flow,
   • number of iterations $n_{\text{iter}}$,
-additionally, for the empirical flow, we have the batch size $n_{b}$ for the input measures.
+additionally, for the empirical flow, we have the batch size $m$ for the input measures.
 We show the complete list of hyperparameters in Table 9.
 Additionally to these parameters, we also have the choice of regularizing functionals, $\mathbb{V}$ and $\mathbb{U}$.
-Remark F.2.
-In all of our domain adaptation experiments, we use a diagonal GMM. We empirically verified that optimizing with respect to complete covariances is numerically unstable and time consuming.
-WGF WGF-GMM
-Benchmark $n$ $n_{b}$ $\alpha$ $n_{\text{iter}}$ $\mathbb{V}$ $\mathbb{U}$ $K$ $\alpha$ $n_{\text{iter}}$ $\mathbb{V}$ $\mathbb{U}$
-Office31 1550 1550 1.0 200 $\mathbb{W}_{2}(P,Q_{T})$ Repulsion 124 1.0 150 - -
-Office-Home 1300 650 1.0 200 $\mathbb{W}_{2}(P,Q_{T})$ Repulsion 975 1.0 200 $\mathbb{W}_{2}(P,Q_{T})$ Repulsion
-TEP 870 290 1.0 150 $\mathbb{W}_{2}(P,Q_{T})$ - 145 0.1 200 - -
-BCI-CIV-2a 800 200 1.0 - - Repulsion 400 0.01 200 Entropy Repulsion
-ISRUC 1000 1000 0.1 200 - - 40 0.1 150 Entropy Repulsion
-Table 10: Hyperparameters used in each benchmark. $n$ denotes number of samples, $\alpha$ denotes learning rate, $n_{\text{iter}}$ denotes number of iterations.
+Benchmark $n$ $m$ $\alpha$ $n_{\text{iter}}$ $\epsilon$ $\mathbb{V}$ $\eta_{V}$ $\mathbb{U}$ $\eta_{U}$ margin $\eta_{G}$
+Supervised $(\mathcal{X}\times\mathcal{Y})$
+Office31 620 620 1.0 100 $10^{-3}$ Entropy $10^{2}$ Repulsion $^{3} 1.01.0$ $10^{-2}$
+[…]
+TEP 870 145 1.0 50 $0.0$ Entropy $10^{2}$ Repulsion $10^{3}$ 1.0 0.0
+BCI-CIV-2a 40 80 1.0 50 $10^{-2}$ Entropy $10^{2}$ Repulsion $10^{5}$ 1.0 0.0
+ISRUC 200 100 0.1 200 0.0 - - - - - -
+Unsupervised $(\mathcal{X})$
+[…]
+[…]
+TEP 580 145 1.0 50 0.0 - - - - - -
+BCI-CIV-2a 80 40 1.0 50 $10^{-2}$ - - - - - -
+ISRUC 250 50 0.1 200 $10^{-3}$ - - - - - -
+Table 9: Hyperparameters used in each benchmark. $n$ denotes number of samples, $m$ denotes the batch size, $\alpha$ denotes learning rate, $n_{\text{iter}}$ denotes number of iterations.
 We offer some additional reasoning on how to choose these parameters.
-The number of samples and components controls the complexity of the calculated barycenter.
+The number of samples controls the complexity of the calculated barycenter.
 We express these values in terms of the number of classes.
 For instance, $n=1550$ in the Office 31 benchmark translates to $n=31\times 50$ samples, that is, $50$ samples per class.
-The same reasoning applies for the batch size.
+The same reasoning applies to the batch size $m$.
 Furthermore, our method is fairly robust to the choice of learning rate.
@@
 This generally strikes a balance on how fast the algorithm converges.
+F.6 Ablation
+Table 10 ablates the effect of the functionals $\mathbb{G},\mathbb{V},\mathbb{U}$ in the Office-Home benchmark.
+First, our WGF consistently outperforms the discrete baseline (77.17 vs.
+77.53 and 77.24 vs.
+78.42 in the unlabeled and labeled cases).
+Second, $\mathbb{V}$ acts as a stabilizer with high entropic regularization ($\epsilon=10^{-2}$).
+While the entropic regularizer blurs the barycentric measure, $\mathbb{V}$ has a sharpness effect.
+Third, our WGF is more sensitive to the entropic regularization, $\epsilon$, particularly when used as a loss function (our case) rather than for OT plan calculations, as in cuturi2014fast.
+This latter point highlights the importance of using a sharpness functional.
+We explore the effects of $\mathbb{G}$, $\mathbb{V}$, and $\mathbb{U}$ in further detail in Appendix F.4.
+Method $\epsilon$ $\mathbb{B}$ $\mathbb{B}{+}\mathbb{U}$ $\mathbb{B}{+}\mathbb{V}$ $\mathbb{B}{+}\mathbb{V}{+}\mathbb{U}$ $\mathbb{B}{+}\mathbb{G}$ $\mathbb{B}{+}\mathbb{G}{+}\mathbb{V}$ $\mathbb{B}{+}\mathbb{G}{+}\mathbb{U}$ $\mathbb{B}{+}\mathbb{G}{+}\mathbb{V}{+}\mathbb{U}$
+Unlabeled setting ($\mathcal{X}$ only)
+Discrete 0.0 77.17
+Discrete 0.01 76.61
+Discrete 0.001 76.56
+WGF (ours) 0.0 77.41 76.83 77.41 77.13 76.88 76.88 77.53 77.53
+WGF (ours) 0.01 75.54 74.42 75.26 75.26 74.55 74.55 75.00 75.00
+WGF (ours) 0.001 77.15 76.88 76.94 76.94 77.28 77.28 77.26 77.26
+Labeled setting ($\mathcal{X}\times\mathcal{Y}$)
+Discrete 0.0 77.24
+Discrete 0.01 76.73
+Discrete 0.001 76.64
+WGF (ours) 0.0 77.97 78.12 78.07 78.25 78.11 78.11 78.19 78.19
+WGF (ours) 0.01 77.59 77.87 77.83 77.83 77.83 77.88 77.89 77.94
+WGF (ours) 0.001 78.16 78.16 78.15 78.16 77.99 78.07 78.42 78.37
+Table 10: Ablation on the contribution of each energy functional on domain adaptation performance.